### v3

# arXiv:2402.18042v3[math.AG] 01 Mar 2026

Abstract
Let $n\geq 0$ and $r>0$ be integers. Let $\mathcal{O}_{X,x}^{h}$ be the henselization of the local ring $\mathcal{O}_{X,x}$ of a scheme $X$ at a point $x\in X$. For a normal crossing variety $Y$ over the spectrum of a field $k$ of positive characteristic $p>0$, K.Sato defined an étale logarithmic Hodge-Witt sheaf $\lambda^{n}_{Y,r}$ (cf. [26, p.726, Definition 3.1.1 (1)]) on the étale site $Y_{\mathrm{\acute{e}t}}$ which agrees with $W_{r}\Omega^{n}_{Y,\log}$ (cf. [12]) in the case where $Y$ is smooth over $\operatorname{Spec}(k)$. In this paper, we prove the Gersten-type conjecture for étale sheaves which satisfy some properties over $\mathcal{O}_{Y,y}^{h}$. For example, $\lambda_{Y,r}^{n}$ and $\mu_{l}^{\otimes n}$ satisfy these properties where $\mu_{l}$ is the étale sheaf of $l$-th roots of unity for an integer $l$ which is prime to the characteristic of $Y$. Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $\mathfrak{X}$ a semistable family over $\operatorname{Spec}(B)$. Suppose that $B$ contains $p$-th roots of unity. As an application of the Gersten-type conjecture for $\lambda^{n}_{r}$, we prove the relative version of the Gersten-type conjecture for the $p$-adic étale Tate twist $\mathfrak{T}_{1}(n)$ (cf. [27, pp.537–538, Definition 4.2.4]) over $\mathcal{O}_{\mathfrak{X},x}^{h}$. Moreover, we prove a generalization of Artin’s theorem (cf. [11, p.98, Théorème (3.1)]) about the Brauer groups.
Indian Institute of Technology Gandhinagar
E-mail address: sakagaito43@gmail.com
1 Introduction
Let $X$ be an equidimensional scheme and noetherian of dimension $d$, $X^{(s)}$ the set of points $x\in X$ of codimension $s$ and $\mathcal{F}^{\bullet}$ a bounded below complex of sheaves of abelian groups on the étale site $X_{\mathrm{\acute{e}t}}$. Let us denote
$\operatorname{H}^{s+t}_{x}(X_{\mathrm{\acute{e}t}},\mathcal{F}^{\bullet}):=\displaystyle\lim_{\begin{subarray}{c}\to\\ x\in U\end{subarray}}\operatorname{H}^{s+t}_{\bar{\{x\}}\cap U}(U_{\mathrm{\acute{e}t}},\mathcal{F}^{\bullet})$
where $U$ runs through open neighborhoods of $x$ in $X$. Then we have a coniveau spectral sequence
$E^{s,t}_{1}=\displaystyle\bigoplus_{x\in X^{(s)}}\operatorname{H}^{s+t}_{x}(X_{\mathrm{\acute{e}t}},\mathcal{F}^{\bullet})\Rightarrow E^{s+t}=\operatorname{H}^{s+t}_{\mathrm{\acute{e}t}}(X,\mathcal{F}^{\bullet})$ (1)
and it induces the Cousin complex
$0\to\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(X,\mathcal{F}^{\bullet})\to\displaystyle\bigoplus_{x\in X^{(0)}}\operatorname{H}^{n}_{x}(X_{\mathrm{\acute{e}t}},\mathcal{F}^{\bullet})\to\displaystyle\bigoplus_{x\in X^{(1)}}\operatorname{H}^{n+1}_{x}(X_{\mathrm{\acute{e}t}},\mathcal{F}^{\bullet})\to\cdots$ (2)
(cf. [4, Part 1, §1], see also [25, Theorem 2.1]). We call the exactness of the Cousin complex (2) the Gersten-type conjecture for étale (hyper)cohomology with values in $\mathcal{F}^{\bullet}$.
Let $Y$ be a normal crossing variety over the spectrum of a field $k$ of positive characteristic $p>0$, that is, $Y$ is everywhere étale locally isomorphic to
$\operatorname{Spec}(k[T_{0},T_{1},\cdots,T_{N}]/(T_{0}T_{1}\cdots T_{a}))$
for some integer $a$ with $0\leq a\leq N=\operatorname{dim}(Y)$ (cf. [28, pp.180–181, Definition 2.1]). Then K.Sato defined the logarithmic Hodge-Witt sheaves
$\lambda^{n}_{Y,r}:=\operatorname{Im}\left(d\log:(\mathbb{G}_{m,Y})^{\otimes n}\to\displaystyle\bigoplus_{x\in Y^{(0)}}i_{y*}W_{r}\Omega^{n}_{y,\log}\right)$
(cf. [26, p.726, Definition 3.1.1 (1)]) and
$\nu^{n}_{Y,r}:=\operatorname{Ker}\left(\delta:\displaystyle\bigoplus_{y\in Y^{(0)}}i_{y*}W_{r}\Omega^{n}_{y,\log}\to\displaystyle\bigoplus_{y\in Y^{(1)}}i_{y*}W_{r}\Omega^{n-1}_{y,\log}\right)$
(cf. [26, p.715, Definition 2.1.1]) on $Y_{\mathrm{\acute{e}t}}$ for any integers $n\geq 0$ and $r>0$ which agree with the logarithmic Hodge-Witt sheaves $W_{r}\Omega^{n}_{Y,\log}$ (cf. [12]) in the case where $Y$ is a smooth scheme over $\operatorname{Spec}(k)$. In general, $\lambda^{n}_{Y,r}\neq\nu^{n}_{Y,r}$ (cf. [26, p.737, Remark 4.2.3]).
In the case where $X$ is the spectrum of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$ and $\mathcal{F}=\nu^{n}_{X,r}$, the Cousin complex (2) is exact by [24, Theorem 1.1]. This is an extension of ([10, Théorème 1.4], [29, p.600, Theorem 4.1]).
In §2, we prove the Gersten-type conjecture for $\lambda_{X,r}^{n}$. Precisely, we prove the following:
Theorem 1.1.
(cf. Theorem 2.26) Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ at a point $y\in Y$. Then the sequence
$0\to\operatorname{H}^{u}_{\mathrm{\acute{e}t}}(A,\lambda^{n}_{r})\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{u}_{x}(A_{\mathrm{\acute{e}t}},\lambda_{r}^{n})\to\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{u+1}_{x}(A_{\mathrm{\acute{e}t}},\lambda_{r}^{n})\to\cdots$
is exact for any integers $n\geq 0$, $u$ and $r>0$.
The outline of the proof is as follows. If $u<0$, then we have
$\operatorname{H}^{u+s}_{x}(A_{\mathrm{\acute{e}t}},\lambda_{r}^{n})=0$
for $x\in\operatorname{Spec}(A)^{(s)}$ by Proposition 2.25. So Theorem 1.1 holds for $u<0$. In the case where $u\geq 1$, Theorem 1.1 follows from the following theorem:
Theorem 1.2.
(cf. Proposition 2.4, Theorem 2.22) Let $Y$ be a normal crossing variety over the spectrum of a field $k$. Let $Y_{1},\cdots,Y_{a}$ be the irreducible components of $Y$.
Let $(\mathcal{F}_{X},N)$ be a pair of an étale sheaf of abelian groups on an object $X$ in $\mathcal{V}$ and a non-negative integer which satisfies Condition 2.11. Here $\mathcal{V}$ is the category whose objects are normal crossing varieties over the spectrum of a field $k$ and whose morphisms are morphisms of schemes. Suppose that $\operatorname{cd}_{q}(k)<\infty$ for a prime number $q$. For an abelian group $M$, $M_{T_{q}}$ denotes the $q$-power torsion subgroup of $M$. Put
$T:=\left\{\begin{array}[]{ll}0&\textrm{if $\mathcal{F}_{Y}$ is a sheaf of torsion groups}\\[5.0pt] 1&\textrm{otherwise}\end{array}\right$.
and $N_{1}=\operatorname{max}\{N,T\}$.
Then $\operatorname{H}_{y}^{u}(Y_{\mathrm{\acute{e}t}},\mathcal{F}_{Y})$ is a torsion group for $y\in Y^{(s)}$ and $u\geq s+T$. Moreover, we have the followings:
  • Property $\operatorname{P}_{1}(a)$: Let $s\geq 0$ be an integer and $i:Z\to Y$ a closed immersion with
    $Z=\displaystyle\bigcup^{a-1}_{m=1}Y_{m}.$
    Then the homomorphism
    $\operatorname{H}^{N^{\prime}+s}_{y}(Z_{\mathrm{\acute{e}t}},i^{*}\mathcal{F}_{Y})_{T_{q}}\xrightarrow{\sim}\operatorname{H}^{N^{\prime}+s}_{y}(Z_{\mathrm{\acute{e}t}},\mathcal{F}_{Z})_{T_{q}}$
    is an isomorphism for $N^{\prime}\geq N_{1}$ and $y\in Y^{(s)}\cap Z$.
  • Property $\operatorname{P}_{2}(a)$: Let the notations be the same as above and $j:U\hookrightarrow Y$ the open complement $Y\setminus Z$. Then a distinguished triangle
    $\cdots\to j_{!}\mathcal{F}_{U}\to\mathcal{F}_{X}\to i_{*}i^{*}\mathcal{F}_{X}\to\cdots$ (3)
    induces an exact sequence
    $0\to\operatorname{H}^{N^{\prime}+s}_{y}(Y_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\operatorname{H}^{N^{\prime}+s}_{y}(Y_{\mathrm{\acute{e}t}},\mathcal{F}_{Y})_{T_{q}}\to\operatorname{H}^{N^{\prime}+s}_{y}(Z_{\mathrm{\acute{e}t}},\mathcal{F}_{Z})_{T_{q}}\to 0$ (4)
    for $y\in Y^{(s)}\cap Z$ and $N^{\prime}\geq N_{1}$.
  • Property $\operatorname{P}_{3}(a)$: Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point of $y\in Y$. Then the sequence
    $0\to\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}$
    $\to\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\to\cdots$
    is exact for $N^{\prime}\geq N_{1}$.
By Remark 2.12, $(\mathcal{F}_{X},N)=(\lambda^{n}_{X,r},1)$ satisfies Condition 2.11 and so Theorem 1.1 holds for $u\geq 1$ by Property $\operatorname{P}_{3}(a)$ of Theorem 1.2. Then we are able to prove Theorem 1.1 for $u=0$ by using the coniveau spectral sequence (1). This concludes the outline of the proof of Theorem 1.1.
Let $l$ be an integer which is prime to $\operatorname{char}(k)$ and $\mu_{l}$ the étale sheaf of $l$-th roots of unity. Then $(\mathcal{F}_{X},N)=(\mu_{l}^{\otimes n},0)$ satisfies Condition 2.11 (cf. Remark 2.12). As an application of Theorem 1.2, we also have the following:
Corollary 1.3.
(cf. Corollary 2.23) Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ over the spectrum of a field $k$ at a point $y\in Y$ and $l$ an integer which is prime to $\operatorname{char}(k)$. Then the sequence
$0\to\operatorname{H}^{u}_{\mathrm{\acute{e}t}}(A,\mu^{\otimes n}_{l})\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{u}_{x}(A_{\mathrm{\acute{e}t}},\mu_{l}^{\otimes n})\to\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{u+1}_{x}(A_{\mathrm{\acute{e}t}},\mu_{l}^{\otimes n})\to\cdots$
is exact for any integers $n\geq 0$ and $u$.
This is an extension of the Bloch-Gabber-Ogus theorem (cf. [4, Corollary 2.2.2]).
In the proof of Theorem 1.2, we use Lemma 2.21. The proof of Theorem 1.2 is similar to that of Lemma 2.21 and we prove them by downward induction on $N^{\prime}$ and induction on the dimension of a scheme. We use Propositions 2.7 and 2.9 to apply downward induction on $N^{\prime}$ and we are able to prove them by using Lemma 2.2. We use Lemma 2.19 to apply induction on the dimension of a scheme. See §2.1 below for more detailed outline of the proof.
As another application of Theorem 1.2, we obtain the relative version of the Gersten-type conjecture in mixed characteristic cases.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $K$ the quotient field of $B$. Let $\mathfrak{X}$ be a semistable family over $\operatorname{Spec}(B)$, that is, a regular scheme of pure dimension which is flat of finite type over $\operatorname{Spec}(B)$, the generic fiber $\mathfrak{X}\otimes_{B}K$ is smooth over $\operatorname{Spec}(K)$ and the special fiber $Y$ of $\mathfrak{X}$ is a reduced divisor with normal crossings on $\mathfrak{X}$. Let $i:Y\to\mathfrak{X}$ be the inclusion of the closed fiber of $\mathfrak{X}$ and $j:\mathfrak{X}\otimes_{B}K\to\mathfrak{X}$ the inclusion of the generic fiber of $\mathfrak{X}$. Then $p$-adic étale Tate twist is defined as follows:
Definition 1.4.
($p$-adic étale Tate twist) (cf. [27, pp.537–538, Definition 4.2.4], [28, p.187, Remark 3.7]) Let the notations be the same as above and $r$ any positive integer. For $n=0$,
$\mathfrak{T}_{r}(0):=\mathbb{Z}/p^{r}.$
For $n\geq 1$, $\mathfrak{T}_{r}(n)$ is defined as a complex which fits into the distinguished triangle
$i_{*}\nu_{Y,r}^{n-1}[-n-1]\to\mathfrak{T}_{r}(n)\to\tau_{\leq n}Rj_{*}\mu_{p^{r}}^{\otimes n}\xrightarrow{\sigma^{n}_{\mathfrak{X},r}}i_{*}\nu_{Y,r}^{n-1}[-n]$
where $\mu_{p^{r}}$ is the sheaf of $p^{r}$-th roots of unity and $\sigma^{n}_{\mathfrak{X},r}$ is induced by the homomorphism $\sigma:i^{*}R^{n}j_{*}\mu_{p^{r}}^{\otimes n}\to\nu^{n-1}_{Y,r}$ which is defined in [28, p.186, Theorem 3.4].
Let $\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$ and $\mathbb{Z}(n)$ be Bloch’s cycle complex for the étale and Zariski topology, respectively (cf. [7, p.779]). For a positive integer $m$, $\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}$ (resp. $\mathbb{Z}/m(n)$) denotes $\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\otimes\mathbb{Z}/m\mathbb{Z}$ (resp. $\mathbb{Z}(n)\otimes\mathbb{Z}/m\mathbb{Z}$). By [28, pp.209–210, Remark 7.2] and [24, Proposition 1.5], we have an isomorphism
$\tau_{\leq n+1}\left(\mathbb{Z}/p^{r}(n)_{\mathrm{\acute{e}t}}\right)\simeq\mathfrak{T}_{r}(n)$
in $D^{b}(\mathfrak{X},\mathbb{Z}/p^{r}\mathbb{Z})$ which is the derived category of bounded complexes of étale $\mathbb{Z}/p^{r}\mathbb{Z}$-sheaves on $\mathfrak{X}_{\mathrm{\acute{e}t}}$. If $\mathfrak{X}$ is smooth over $\operatorname{Spec}(B)$, then trancation is unnecessary by [7, p.786, Corollary 4.4]. By an application of Theorem 1.1, we have the relative version of the Gersten-type conjecture for $\mathfrak{T}_{r}(n)$ as follows:
Theorem 1.5.
(cf. Theorem 2.28) Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let $\mathfrak{X}$ be a semistable family over $\operatorname{Spec}(B)$ and $R$ the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of $\mathfrak{X}$ at a point $x$ of the closed fiber of $\mathfrak{X}$. Put $Z=\operatorname{Spec}(R/(\pi))$. Suppose that $B$ contains $p^{r}$-th roots of unity. Then the sequence
$0\to\operatorname{H}_{Z}^{q}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\to\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}_{z}^{q}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\to\displaystyle\bigoplus_{z\in Z^{(1)}}\operatorname{H}_{z}^{q+1}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\to\cdots$
is exact for $q\geq n+2$.
In §3, we observe a relation between the Gersten-type conjecture and motivic cohomology groups. As an application of the Beilinson-Lichtenbaum conjecture (cf. [7, p.774, Theorem 1.2.2] and [30]), we prove the following:
Proposition 1.6.
(cf. Proposition 3.1) Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $R$ a local ring (resp. a henselian local ring) of a regular scheme of finite type over $\operatorname{Spec}(B)$. Let $U$ be the generic fiber of $\operatorname{Spec}(R)$ and $n$ a non-negative integer. Suppose that $B$ contains $p$-th roots of unity.
Then the sequence
$0\to$ $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(U,\mu_{p}^{\otimes n})\to\displaystyle\bigoplus_{x\in U^{(0)}}\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{p}^{\otimes n})\to\displaystyle\bigoplus_{x\in U^{(1)}}\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{p}^{\otimes(n-1)})$
$\to$ $\cdots$ (5)
is exact if and only if we have isomorphisms
$\operatorname{H}^{n+1}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n))\simeq 0$
and
$\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n))\simeq\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n+1))$
for any integer $s\geq 1$.
By Proposition 1.6, the sequence (1.6) is exact for any integer $n\geq 0$ if and only if
$\operatorname{H}_{\mathrm{Zar}}^{u}(U,\mathbb{Z}/p(n))=0$ (6)
for any integers $n\geq 0$ and $u>n$. In the case where $R$ is (the henselization of) the local ring $\mathcal{O}_{\mathfrak{X},x}$ of a semistable family $\mathfrak{X}$ over $\operatorname{Spec}(B)$, the equation (6) holds for any integers $n\geq 0$ and $u>n+1$ if and only if the equation
$\operatorname{H}^{u}_{\mathrm{Zar}}(R,\mathbb{Z}/p(n))=0$
holds for any integers $n\geq 0$ and $u>n+1$ by Proposition 1.6 and [24, Proposition 2.1] (cf. Remark 3.2). By an application of Theorem 1.5 and Proposition 1.6, we have the following:
Corollary 1.7.
(cf. Corollary 3.4) Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $R$ the henselization of a semistable family over $\operatorname{Spec}(B)$. Suppose that $\operatorname{dim}(R)=2$ and $B$ contains $p$-th roots of unity. Then the sequence
$0\to$ $\operatorname{H}^{s}_{\mathrm{\acute{e}t}}(R,\mathfrak{T}_{1}(n))\to\bigoplus_{x\in\operatorname{Spec}(R)^{(0)}}\operatorname{H}^{s}_{x}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to$
$\bigoplus_{x\in\operatorname{Spec}(R)^{(1)}}\operatorname{H}^{s+1}_{x}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to\bigoplus_{x\in\operatorname{Spec}(R)^{(2)}}\operatorname{H}^{s+2}_{x}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to 0$
is exact for any integer $s\geq 0$.
In §4, we compute motivic cohomology groups in global cases. We prove the following:
Proposition 1.8.
(cf. Proposition 4.1, Proposition 4.5 and Proposition 4.10) Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let
$C=B[T_{0},T_{1},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi)$
for $0\leq a\leq N$. Then we have
$\operatorname{H}^{q}_{\mathrm{Zar}}(C,\mathbb{Z}/m(n))=0$ (7)
for integers $m>0$ and $q\geq n+1$. Moreover, we have
$\operatorname{H}^{q}_{\mathrm{Zar}}(C/(\pi),\mathbb{Z}(n))=0$
for $q\geq n+1$.
By using the localization theorem ([7, p.779, Theorem 3.2.a)]), we are able to prove Proposition 1.8 by induction on $a$. The outline of the proof of the isomorphism (7) is as follows: We have a homomorphism of polynomial rings over $B$
$B[T^{\prime}_{0},\cdots,T^{\prime}_{N}]\to B[T_{0},\cdots,T_{N}]$
which sends $T^{\prime}_{a-1}$ to $T_{a-1}T_{a}$ and sends $T^{\prime}_{r}$ to $T_{r}$ for $r\neq a-1$. Then this homomorphism induces an isomorphism
$\left(B[T^{\prime}_{0},\cdots,T^{\prime}_{N}]/(T^{\prime}_{0}\cdots T^{\prime}_{a-1}-\pi)\right)_{T^{\prime}_{a}}$
$\simeq$ $\left(B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi)\right)_{T_{a}}$
for $a>0$. So, if the homomorphism
$\operatorname{H}^{n}_{\mathrm{Zar}}(C_{T_{a}},\mathbb{Z}/m(n))\to\operatorname{H}^{n-1}_{\mathrm{Zar}}(C/(T_{a}),\mathbb{Z}/m(n-1))$ (8)
is surjective, the isomorphism (7) follows from the localization theorem and an inductive argument on $a$. Moreover, the surjectivity of the homomorphism (8) is proved by using the following:
Proposition 1.9.
(cf. Proposition 4.6 and Proposition 4.7) Let $\mathfrak{X}$ be a regular scheme which is essentially of finite type over the spectrum of a discrete valuation ring $B$ in mixed characteristic $(0,p)$.
  1. 1.
    Let $i:Z\to\mathfrak{X}$ be a closed immersion of codimension $1$ and $Z$ an irreducible closed scheme of $\mathfrak{X}$ with $\operatorname{char}(Z)=p$. Then the canonical map induces a quasi-isomorphism
    $\tau_{\leq n+2}\left(\mathbb{Z}(n-1)_{\mathrm{\acute{e}t}}^{Z}[-2]\right)\xrightarrow{\simeq}\tau_{\leq n+2}Ri^{!}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}.$
  2. 2.
    The canonical map induces a quasi-isomorphism
    $\tau_{\leq n+1}\left(\mathbb{Z}(n)^{\mathfrak{X}}\right)\xrightarrow{\simeq}\tau_{\leq n+1}\left(R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\right)$
    where $\epsilon:\mathfrak{X}_{\mathrm{\acute{e}t}}\to\mathfrak{X}_{\mathrm{Zar}}$ is the canonical map of sites and $\epsilon_{*}$ is the forgetful functor.
In the case where $\mathfrak{X}$ is smooth over $\operatorname{Spec}(B)$, $\mathcal{H}^{q}(\mathbb{Z}(n))=0$ for $q\geq n+1$ by [7, p.786, Corollary 4.4]. So Proposition 1.9 is an extension of [7, p.774, Theorem 1.2.1 and Theorem 1.2.2]. As an application of Proposition 1.8, we obtain the following:
Proposition 1.10.
(cf. Corollary 4.18 and Remark 4.19) Let $B$ be a henselian discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let
$\mathfrak{X}=\operatorname{Proj}\left(B[T_{0},\cdots,T_{N+1}]/(T_{0}\cdots T_{a}-\pi T_{N+1}^{a+1})\right)$
for $0\leq a\leq N$ and $i:Y\to\mathfrak{X}$ the inclusion of the closed fiber of $\mathfrak{X}$. Then the homomorphism
$\Gamma\left(\mathfrak{X},R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)\to\Gamma\left(Y,i^{*}R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$
is injective for any integers $n\geq 0$ and $l>0$ with $(l,p)=1$. Here $\alpha:\mathfrak{X}_{\mathrm{\acute{e}t}}\to\mathfrak{X}_{\mathrm{Nis}}$ is the canonical map of sites.
In §5, we prove the following as an application of the proper base change theorem ([17, p.224, VI, Corollary 2.7]) and the local-global principle ([24, Theorem 1.7]):
Theorem 1.11.
(cf. Theorem 5.5) Let $B$ be a henselian excellent discrete valuation ring of mixed characteristic $(0,p)$ and $\mathfrak{X}$ a semistable family and proper over $\operatorname{Spec}(B)$. Let $i:Y\to\mathfrak{X}$ be the inclusion of the closed fiber of $\mathfrak{X}$. Suppose that $\operatorname{dim}(\mathfrak{X})=2$ and $B$ contains $p$-th roots of unity. Then the homomorphism
$\operatorname{H}^{s}_{\mathrm{Nis}}(\mathfrak{X},R^{t}\alpha_{*}\mathfrak{T}_{1}(n))\xrightarrow{\simeq}\operatorname{H}^{s}_{\mathrm{Nis}}(Y,i^{*}R^{t}\alpha_{*}\mathfrak{T}_{1}(n))$
is an isomorphism for integers $s\geq 0$ and $t\geq 2$. Moreover, we have an isomorphism
$\operatorname{H}^{s}_{\mathrm{Nis}}(\mathfrak{X},R^{n+1}\alpha_{*}\mathfrak{T}_{1}(n))\xrightarrow{\simeq}\operatorname{H}^{s}_{\mathrm{Nis}}(Y,R^{1}\alpha_{*}\lambda_{1}^{n})$
for integers $s\geq 0$ and $n\geq 1$. Thus, the sequence
$0\to\displaystyle\bigoplus_{x\in\mathfrak{X}^{(0)}}\operatorname{H}_{x}^{n+r}(\mathfrak{X}_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to\displaystyle\bigoplus_{x\in\mathfrak{X}^{(1)}}\operatorname{H}_{x}^{n+r+1}(\mathfrak{X}_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))$
$\to\displaystyle\bigoplus_{x\in\mathfrak{X}^{(2)}}\operatorname{H}_{x}^{n+r+2}(\mathfrak{X}_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to 0$
is exact for integers $n\geq 1$ and $r\geq 2$.
Let the notation be the same as above. Then Theorem 1.11 relates to a generalization of Artin’s theorem that the Brauer group of $\mathfrak{X}$ is isomorphic to that of $Y$ (cf. Remark 5.6).
At the end of the paper, we raise questions (see Question 5.10 and Question 5.18) which relates to Kato conjecture (cf. [13], [14, p.125, Conjecture 0.2 and Conjecture 0.3]).
Notations
For a scheme $X$, $X^{(s)}$ denotes the set of points of codimension $s$ and $k(X)$ denotes the ring of rational functions on $X$. For a ring $A$, $k(A)$ denotes $k(\operatorname{Spec}(A))$. For a point $x\in X$, $\kappa(x)$ denotes the residue field of $x$, $\mathcal{O}_{X,x}$ denotes the local ring of $X$ at $x\in X$ and $\mathcal{O}^{h}_{X,x}$ denotes the henselization of $\mathcal{O}_{X,x}$.
$X_{\mathrm{\acute{e}t}}$, $X_{\mathrm{Nis}}$ and $X_{\mathrm{Zar}}$ denote the category of étale schemes over a scheme $X$ equipped with the étale, Nisnevich and Zariski topology, respectively. The category of sheaves on $X_{\mathrm{\acute{e}t}}$ is denoted by $\mathbb{S}_{X_{\mathrm{\acute{e}t}}}$. Let $\epsilon:X_{\mathrm{\acute{e}t}}\to X_{\mathrm{Zar}}$ (resp. $\alpha:X_{\mathrm{\acute{e}t}}\to X_{\mathrm{Nis}}$) be the canonical map of sites. Then $\epsilon_{*}$ (resp. $\alpha_{*}$) denotes the forgetful functor.
$\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$ (resp. $\mathbb{Z}(n)$, $\mathbb{Z}(n)_{\mathrm{Nis}}$) denotes Bloch’s cycle complex for étale (resp. Zariski, Nisnevich) topology and $\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}=\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\otimes\mathbb{Z}/m\mathbb{Z}$ (resp. $\mathbb{Z}/m(n)=\mathbb{Z}(n)\otimes\mathbb{Z}/m\mathbb{Z}$, $\mathbb{Z}/m(n)_{\mathrm{Nis}}=\mathbb{Z}(n)_{\mathrm{Nis}}\otimes\mathbb{Z}/m\mathbb{Z}$) for a positive integer $m$. For an integer $m>0$, $\mu_{m}$ denotes the sheaf of $m$-th roots of unity. For a smooth scheme $Y$ over the spectrum of a field of positive characteristic $p>0$, $\nu^{n}_{Y,r}=W_{r}\Omega_{Y,\log}^{n}$ denotes the logarithmic de Rham-Witt sheaf (cf. [12], [29, p.575, Definition 2.6]).
2 Gersten-type conjecture
Let $Y$ be a normal crossing variety over the spectrum of a field $k$ of positive characteristic $p>0$, that is, $Y$ is everywhere étale locally isomorphic to
$\operatorname{Spec}(k[T_{0},T_{1},\cdots,T_{M}]/(T_{0}T_{1}\cdots T_{a}))$
for some integer $a$ with $0\leq a\leq M=\operatorname{dim}(Y)$ (cf. [28, pp.180–181, Definition 2.1]). In this section, we prove the Gersten-type conjecture for a pair of a sheaf $\mathcal{F}$ on $Y_{\mathrm{\acute{e}t}}$ and a non-negative integer $N$ which satisfies Condition 2.11 over the henselization ring of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$ ( Theorem 1.2, Theorem 2.22). Let
$\lambda^{n}_{Y,r}:=\operatorname{Im}\left(d\log:(\mathbb{G}_{m,Y})^{\otimes n}\to\displaystyle\bigoplus_{y\in Y^{(0)}}i_{x*}W_{r}\Omega^{n}_{y,\log}\right)$
be the logarithmic Hodge-Witt sheaves (cf. [26, p.726, Definition 3.1.1 (1)]). For example, a pair $(\mathcal{F},N)=(\lambda^{n}_{Y,r},1)$ satisfies Condition 2.11 (see Remark 2.12.2. (ii)).
Let $\mathfrak{X}$ be a semistable family over the spectrum of a discrete valuation ring $B$ of mixed characteristic $(0,p)$. Suppose that $B$ contains $p$-th roots of unity. As an application of Theorem 1.2, we prove the relative version of Gersten-type conjecture for $\mathfrak{T}_{1}(n)$ over the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ at a point $x\in\mathfrak{X}$ ( Theorem 2.26).
2.1 The outline of the proof of Theorem 1.2
The outline of the proof of Theorem 1.2 is as follows: By [17, p.93, III, Corollary 1.28], it suffices to prove $\operatorname{P}_{1}(a)$ and $\operatorname{P}_{2}(a)$ in the case where $Y$ is replaced by the spectrum of the henselization $A=\mathcal{O}^{h}_{Y,y}$ of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$ and a point $y\in Y$ is replaced by the maximal ideal of $A$. We put $\operatorname{Spec}(C)=\operatorname{Spec}(A)\times_{Y}Z$ and prove Theorem 1.2 by induction on $a=\#(\operatorname{Spec}(A)^{(0)})$.
In the case where $a=1$, then $\operatorname{Spec}(C)=\emptyset$ and so $\operatorname{P}_{1}(1)$ and $\operatorname{P}_{2}(1)$ hold. Since $Y$ is a smooth scheme over the spectrum of a field, $\operatorname{P}_{3}(1)$ holds by Condition 2.11.
Assume that $\operatorname{P}_{1}(u)$, $\operatorname{P}_{2}(u)$ and $\operatorname{P}_{3}(u)$ hold for $u\leq a$, then we prove $\operatorname{P}_{1}(a+1)$, $\operatorname{P}_{2}(a+1)$ and $\operatorname{P}_{3}(a+1)$.
First, we prove $\operatorname{P}_{1}(a+1)$. Since $(\operatorname{Spec}(C)^{(0)})=a$, $\operatorname{P}_{1}(a+1)$ follows from $\operatorname{P}_{3}(a)$ and Lemma 2.18.
Next we prove $\operatorname{P}_{2}(a+1)$ and $\operatorname{P}_{3}(a+1)$ by downward induction on $N^{\prime}(\geq N_{1})$. Let $\mathcal{G}$ be a sheaf of abelian groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Then $\operatorname{H}^{N^{\prime}+s}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{G})_{T_{q}}$ vanishes for $x\in\operatorname{Spec}(A)^{(s)}$ and sufficiently large $N^{\prime}$ by Proposition 2.7 and Remark 2.8. Since $\operatorname{cd}_{q}(k)$ and $\operatorname{tr.deg}_{k}\kappa(y)$ are finite, $\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}$ vanishes for sufficiently large $N^{\prime}$ by the isomorphism (14). So $\operatorname{P}_{2}(a+1)$ and $\operatorname{P}_{3}(a+1)$ hold for sufficiently large $N^{\prime}$.
Assume that $\operatorname{P}_{2}(a+1)$ and $\operatorname{P}_{3}(a+1)$ hold for $N^{\prime}>w(\geq N_{1})$. By the definition, $\operatorname{Spec}(A)\times_{Y}Y_{a+1}$ is the spectrum of the henselization $A_{a+1}$ of the local ring $\mathcal{O}_{Y_{a+1},y}$ of a smooth scheme $Y_{a+1}$ at a point $y\in Y_{a+1}$. Moreover, $\operatorname{Spec}(C)\times_{Y}Y_{a+1}$ is the spectrum of the henselization $C_{a+1}$ of the local ring $\mathcal{O}_{Z\times_{Y}Y_{a+1}},y$ of a normal crossing variety $Z\times_{Y}Y_{a+1}$ at a point $y\in Z\times_{Y}Y_{a+1}$. Since we have
$U^{\prime}=\operatorname{Spec}(A)\setminus\operatorname{Spec}(C)=\operatorname{Spec}(A_{a+1})\setminus\operatorname{Spec}(C_{a+1}),$
there is an open immersion $j^{\prime}:U^{\prime}\to\operatorname{Spec}(A_{a+1})$. Let $i_{a+1}:\operatorname{Spec}(A_{a+1})\to\operatorname{Spec}(A)$ be a closed immersion. Then $j=i_{a+1}\circ j^{\prime}$ and so we have an isomorphism
$\operatorname{H}^{u}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})=\left\{\begin{array}[]{ll}\operatorname{H}^{u}_{x}((A_{a+1})_{\mathrm{\acute{e}t}},(j^{\prime})_{!}\mathcal{F}_{U^{\prime}}),&x\in\operatorname{Spec}(A)\cap\operatorname{Spec}(A_{a+1})\\[5.0pt] 0&x\in\operatorname{Spec}(A)\setminus\operatorname{Spec}(A_{a+1})\end{array}\right$.
for any $u\geq 0$ by Lemma 2.20 and [25, Lemma 3.7]. Moreover, the sequence
$0$ $\to\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A_{a+1},(j^{\prime})_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A_{a+1})^{(0)}}\operatorname{H}^{N^{\prime}}_{x}((A_{a+1})_{\mathrm{\acute{e}t}},(j^{\prime})_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}$
$\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A_{a+1})^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}((A_{a+1})_{\mathrm{\acute{e}t}},(j^{\prime})_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\to\cdots$ (9)
is exact by Lemma 2.21 and $\operatorname{P}_{3}(a)$. Since we have the exact sequence (2.1) and $\operatorname{P}_{3}(a)$ holds by the assumption, we are able to prove that $\operatorname{P}_{2}(a+1)$ holds for $N^{\prime}=w$ by Proposition 2.9, Lemma 2.19 and induction on $s=\operatorname{dim}(A)$. Moreover, $\operatorname{P}_{3}(a+1)$ follows from the exact sequence (2.1), $\operatorname{P}_{3}(a)$ and $\operatorname{P}_{2}(a+1)$. This concludes the outline of the proof.
2.2 Equi-characteristic cases
Lemma 2.1.
Let $X$ be an equidimensional scheme and noetherian of dimension $d\geq 1$. Let $\mathcal{F}^{\bullet}$ be a complex of sheaves of abelian groups on $X_{\mathrm{\acute{e}t}}$. Consider the coniveau spectral sequence
$E^{u,v}_{1}=\bigoplus_{x\in X^{(u)}}\operatorname{H}_{x}^{u+v}(X_{\mathrm{\acute{e}t}},\mathcal{F}^{\bullet})\Rightarrow E^{u+v}=\operatorname{H}^{u+v}_{\mathrm{\acute{e}t}}(X,\mathcal{F}^{\bullet})$ (10)
(cf. [4, Part 1, §1], [25, Theorem 2.1]). Let $n$ be an integer. Let $e$ be a positive integer such that d1d-1 or $d$. Moreover, we assume the following two conditions:
  • (a)
    The homomorphism
    $E^{n+e}\to E_{2}^{0,n+e}$
    is injective.
  • (b)
    If $e\neq 1$, we have
    $E^{u,v}_{2}=\begin{cases}E^{u+v}&(\textrm{if}~~u=0)\\ \\ 0&(\textrm{if}~~e\geq 3~~\textrm{and}~~0<u\leq e-2)\end{cases}$
    for u+v=n+e1u+v=n+e-1.
Then we have
$E_{2}^{e,n}=0.$
Proof.
We remark that there exists the spectral sequence (10) because the proof of [25, Theorem 2.1] is valid without assuming that $\mathcal{F}^{\bullet}$ is a bounded below complex.
By the definition of the spectral sequence, there is a filtration
$E^{m}=E^{m}_{0}\supset E^{m}_{1}\supset\cdots\supset E^{m}_{m}\supset 0$
such that
$E^{m}_{u}/E^{m}_{u+1}=E^{u,m-u}_{\infty}$
for $0\leq u\leq d$. By the assumption (a), the homomorphism
$E^{n+e}\to E_{\infty}^{0,n+e}$
is an isomorphism and so we have
$E_{\infty}^{e,n}=0,$
that is, we have
$E_{r}^{e,n}=0$
for sufficiently large $r$. So it suffices to prove that
$E_{r+1}^{e,n}=\operatorname{Coker}\left(E_{r}^{e-r,n+r-1}\to E_{r}^{e,n}\right)=E_{r}^{e,n}$ (11)
for $r\geq 2$. If $r>e$, then er<0e-r<0 and so the isomorphism (11) holds for $r>e$. By the assumption (b), we have isomorphisms
$E^{n+e-1}\xrightarrow{\sim}E^{0,n+e-1}_{\infty}\xrightarrow{\sim}E^{0,n+e-1}_{r}\xrightarrow{\sim}E^{0,n+e+1}_{2}$
for $r\geq 2$. So we have
$\operatorname{Ker}\left(E^{0,n+e-1}_{e}\to E^{e,n}_{e}\right)=E^{0,n+e-1}_{e+1}=E^{0,n+e-1}_{e}$
and the isomorphism (11) holds for $r=e$. If $2\leq r\leq e-1$, then
$E_{r}^{e-r,n+r-1}=0$
by the assumption (b). So the isomorphism (11) holds for $2\leq r\leq e-1$. Hence, the isomorphism (11) holds for $r\geq 2$. This completes the proof. ∎
Lemma 2.2.
Let the notations be the same as in Lemma 2.1. Let $q$ be a prime number. For an abelian group $M$, $M_{T_{q}}$ denotes the $q$-power torsion subgroup of $M$. We assume the following three conditions:
  • (a)
    The homomorphism
    $(E^{n+e})_{T_{q}}\to(E_{2}^{0,n+e})_{T_{q}}$
    is injective.
  • (b)
    If u+v=n+e1u+v=n+e-1 and $u\leq e-2$, then $E_{2}^{u,v}$ is a torsion group. Moreover, $E^{n+e}$ is also a torsion group.
  • (c)
    If $e\neq 1$, we have
    $(E^{u,v}_{2})_{T_{q}}=\begin{cases}(E^{u+v})_{T_{q}}&(\textrm{if}~~u=0)\\ \\ 0&(\textrm{if}~~e\geq 3~~\textrm{and}~~0<u\leq e-2)\end{cases}$
    for u+v=n+e1u+v=n+e-1.
Then we have
$(E_{2}^{e,n})_{T_{q}}=0.$
Proof.
Let the notations be the same as in the proof of Lemma 2.1. By the assumption (b), $E^{n+e}_{r}$ is a torsion group for any $r$ and so $E^{n+e}_{r}/E^{n+e}_{r+1}=E_{\infty}^{r,n+e-r}$ is a torsion group for any $r$. Especially, $E_{\infty}^{e,n}$ is a torsion group. Moreover, we have
$(E_{\infty}^{e,n})_{T_{q}}=0$
by the assumption (a). By the assumption (b), $E_{r}^{e-r,n+r-1}$ is a torsion group for $r\geq 2$ and so we have
$(E_{r+1}^{e,n})_{T_{q}}=\operatorname{Coker}\left((E_{r}^{e-r,n+r-1})_{T_{q}}\to(E_{r}^{e,n})_{T_{q}}\right)$
for $r\geq 2$ by (11). In the same way as in the proof of Lemma 2.1, we are able to prove that
$(E_{r+1}^{e,n})_{T_{q}}=(E_{r}^{e,n})_{T_{q}}$
for $r\geq 2$ by the assumption (c). Hence the statement follows. ∎
In this section, we frequently assume that a ring $A$ is a formally equidimensional noetherian local ring (cf. [16, p.251, Definition]). For example, an equidimensional local ring which is essentially of finite type over a regular local ring is formally equidimensional by [16, p.251, Theorem 31.6 (iii)].
Remark 2.3.
Let $A$ be a formally equidimensional noetherian local ring. By [16, p.251, Theorem 31.6 (iv)] and [16, p.250, Theorem 31.5], AA is equidimensional. By [16, p.251, Theorem 31.6 (i)], $A_{\mathfrak{p}}$ is also formally equidimensional for any $\mathfrak{p}\in\operatorname{Spec}(A)$. Moreover, the henselization $A_{\mathfrak{p}}^{h}$ of $A_{\mathfrak{p}}$ is formally equidimensional by the definition.
Proposition 2.4.
Let $A$ be a formally equidimensional noetherian local ring, $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$ and $x\in\operatorname{Spec}(A)^{(s)}$. Put
$T:=\left\{\begin{array}[]{ll}0&\textrm{if $\mathcal{F}$ is a sheaf of torsion groups}\\[5.0pt] 1&\textrm{otherwise}.\end{array}\right$.
Then $\operatorname{H}^{t}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})$ is a torsion group for $t\geq s+T$.
Proof.
By Remark 2.3 and [17, p.93. III, Corollary 1.28], it suffices to prove the statement in the case where $A$ is a formally equidimensional henselian local ring and $x$ is the maximal ideal of $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
In the case where $\operatorname{dim}(A)=0$, the statement follows from [31, p.180, Theorem 6.5.8].
Suppose that the statement holds in the case where $\operatorname{dim}(A)\leq s$. Suppose that $\operatorname{dim}(A)=s+1$. Consider the spectral sequence
$E^{u,v}_{1}=\bigoplus_{x\in(\operatorname{Spec}(A))^{(u)}}\operatorname{H}_{x}^{u+v}(A_{\mathrm{\acute{e}t}},\mathcal{F})\Rightarrow E^{u+v}=\operatorname{H}_{\mathrm{\acute{e}t}}^{u+v}(A,\mathcal{F}).$ (12)
By [31, p.180, Theorem 6.5.8], $E^{u+v}$ is a torsion group for $u+v\geq T$ and so $E^{s+1,v}_{\infty}$ is also a torsion group for $v\geq 1$. By the assumption, $E^{u,v}_{r}$ is a torsion group for $u\leq s$, $v\geq T$ and $r\geq 1$. By the definition of the spectral sequence, the sequence
$E_{r}^{s-r+1,v+r-1}\to E^{s+1,v}_{r}\to E^{s+1,v}_{r+1}\to 0$ (13)
is exact for $v\geq T$ and $r\geq 1$. Hence $E^{s+1,v}_{1}$ is a torsion group for $v\geq T$. This completes the proof. ∎
Corollary 2.5.
Let the notations $A$, $\mathcal{F}$ and $T$ be the same as in Proposition 2.4. Suppose that $\operatorname{dim}(A)=s$. Then $\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,\mathcal{F})$ is a torsion group for $t\geq s+T$.
Proof.
By the spectral sequence (12), the statement follows from Proposition 2.4 and [17, p.93, III, Corollary 1.28]. ∎
Remark 2.6.
Let $A$ be a henselian local ring, $I$ an ideal of $A$ and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Then we have an isomorphism
$\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,\mathcal{F})\xrightarrow{\simeq}\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A/I,i^{*}\mathcal{F})$ (14)
for $t\geq 0$ where $i:\operatorname{Spec}(A/I)\to\operatorname{Spec}(A)$ is the closed immersion by [7, p.777, The proof of Proposition 2.2.b)]. So $\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,\mathcal{F})$ is a torsion group for $t>0$. Moreover, we have an isomorphism
$\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,j_{!}\mathcal{F}_{U})\simeq 0$
for any $t\geq 0$ by (14) where $j:U\to\operatorname{Spec}(A)$ is the immersion of the open complement of $\operatorname{Spec}(A/I)$ in $\operatorname{Spec}(A)$.
Proposition 2.7.
Let $A$ be a formally equidimentional noetherian local ring over a field $k$, $l$ a prime number which is prime to $\operatorname{char}(k)$ and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Let the notation $T$ be the same as in Proposition 2.4. Suppose that $\operatorname{cd}_{l}(k)<\infty$ and $\operatorname{tr.deg}_{k}A_{x}<\infty$ for any $x\in\operatorname{Spec}(A)^{(0)}$. Put
$N:=\operatorname{cd}_{l}(k)+\max\{\operatorname{tr.deg}_{k}A_{x}|x\in\operatorname{Spec}(A)^{(0)}\}.$
For a group $M$, $M_{T_{l}}$ denotes the $l$-power torsion subgroup of $M$. Then we have
$\operatorname{H}^{m}_{x}(A,\mathcal{F})_{T_{l}}=0$
for $x\in\operatorname{Spec}(A)^{(s)}$ and $m\geq s+N+T+1$.
Proof.
It suffices to prove the statement in the case where $A$ is a formally equidimensional henselian local ring and $x$ is the maximal ideal of $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
In the case where $\operatorname{dim}(A)=0$, the statement follows from [18, (3.3.3) Proposition] and [18, (6.5.14) Theorem].
Suppose that the statement holds in the case where $\operatorname{dim}(A)\leq s$. Suppose that $\operatorname{dim}(A)=s+1$. Consider the spectral sequence (12). By the assumption, it suffices to prove that the homomorphism
$(E_{1}^{s,v})_{T_{l}}\to(E_{1}^{s+1,v})_{T_{l}}$
is surjective for $v\geq N+T+1$, that is,
$(E_{2}^{s+1,v})_{T_{l}}=0$ (15)
for $v\geq N+T+1$. In order to prove the statement, we use Lemma 2.2. Since $A$ is henselian, we have an isomorphism
$\operatorname{H}^{m}_{\mathrm{\acute{e}t}}(A,\mathcal{F})\simeq\operatorname{H}^{m}_{\mathrm{\acute{e}t}}(\kappa(x),i^{*}\mathcal{F})$ (16)
for $m\geq 0$ where $i:\operatorname{Spec}(\kappa(x))\to\operatorname{Spec}(A)$ is the closed immersion. Since $A$ is universally catenary by [16, p.251, Theorem 31.6 (iv)], we have
$\operatorname{tr.deg}_{k}\kappa(x)=(N-\operatorname{cd}_{l}(k))-(s+1)$
by [16, p.119, Theorem 15.6]. So we have
$(E^{m})_{T_{l}}=\operatorname{H}^{m}_{\mathrm{\acute{e}t}}(A,\mathcal{F})_{T_{l}}=0$ (17)
for $m\geq N+T-s$ by [18, (6.5.14) Theorem] and [18, (3.3.3) Proposition]. Especially, we have
$(E^{m})_{T_{l}}=0$
for $m\geq s+N+T+2$ and so the assumption (a) of Lemma 2.2 is satisfied. If $u+v=s+N+T+1$ and $u\leq s-1$, then $v\geq N+T+2$ and so $E^{u,v}_{2}$ is a torsion group by Proposition 2.4. Moreover, $E^{m}$ is a torsion group for $m\geq s+N+T+2$ by (16) and [31, p.180, Theorem 6.5.8]. So the assumption (b) of Lemma 2.2 is satisfied. By the assumption, we have
$(E^{u+v})_{T_{q}}=(E^{u,v}_{2})_{T_{q}}=0$
for $u+v\geq s+N+T+1$ and $u\leq s-1$. So the assumption (c) of Lemma 2.2 is satisfied. Hence, the isomorphism (15) holds. This completes the proof.
Remark 2.8.
Let $A$ be a formally equidimensional local ring of positive characteristic $p>0$ and $\mathcal{F}$ a sheaf of groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Since $\operatorname{cd}_{p}(A)=1$, we have
$\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,\mathcal{F})_{T_{p}}=0$
for $t\geq T+2$ by a similar argument as in the proof of [18, Chapter III, §3, (3.3.3) Proposition]. Moreover, we have
$\operatorname{H}^{t}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{p}}=0$
for $x\in\operatorname{Spec}(A)^{(s)}$ and $t\geq T+s+2$ by using a similar argument as in the proof of Proposition 2.7 (or [18, Chapter III, §3, (3.3.3) Proposition]).
Proposition 2.9.
Let $A$ be a formally equidimensional local ring and $\mathcal{F}$ a sheaf on the étale site $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$. Let $N$ be a non-negative integer and $l$ a prime number. Suppose that $\operatorname{dim}(A)=s>1$ and $N\geq T$ where the notation $T$ is the same as in Proposition 2.4.
If $\operatorname{H}^{N+s-1}_{\mathrm{\acute{e}t}}(A,\mathcal{F})$ is a torsion group and the sequence
$0\to$ $\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F})_{T_{l}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{l}}\to\cdots\to\bigoplus_{x\in\operatorname{Spec}(A)^{(s-2)}}\operatorname{H}^{N^{\prime}+s-2}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{l}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(A)^{(s-1)}}\operatorname{H}^{N^{\prime}+s-1}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{l}}$
is exact for $N^{\prime}\geq N+1$, then the sequence
$\bigoplus_{x\in\operatorname{Spec}(A)^{(s-2)}}\operatorname{H}^{N+s-2}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{l}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(s-1)}}\operatorname{H}^{N+s-1}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{l}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(A)^{(s)}}\operatorname{H}^{N+s}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{l}}\to 0$
is exact.
Proof.
Consider the spectral sequence
$E^{u,v}_{1}=\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(u)}}\operatorname{H}^{u+v}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})\Rightarrow E^{u+v}=\operatorname{H}^{u+v}_{\mathrm{\acute{e}t}}(A,\mathcal{F}).$
Then it suffices to prove that
$(E^{e,N}_{2})_{T_{l}}=0$ (18)
for $s-1\leq e\leq s$. We prove the isomorphism (18) by using Lemma 2.2.
By the assumption, the homomorphism
$(E^{e+N})_{T_{l}}\to(E^{0,e+N}_{2})_{T_{l}}$
is injective. If u+v=N+e1u+v=N+e-1 and $u\leq e-2$, then $v\geq N+1$. If $e\neq 1$, then $N+e-1\geq N+1$. So we have
$(E^{u,v}_{2})_{T_{q}}=\begin{cases}(E^{u+v})_{T_{q}}&(\textrm{if}~~u=0)\\ \\ 0&(\textrm{if}~~e\geq 3~~\textrm{and}~~0<u\leq e-2)\end{cases}$
for $e\neq 1$ and u+v=N+e1u+v=N+e-1 by the assumption. By Proposition 2.4, $E^{u,v}_{2}$ is a torsion group in the case where u+v=N+e1u+v=N+e-1 and $u\geq e-2$. Moreover, $E^{N+e}$ is a torsion group by Corollary 2.5 and the assumption. Hence, the assumptions of Lemma 2.2 are satisfied and the isomorphism (18) holds. This completes the proof. ∎
Corollary 2.10.
Let the notations AA and $\mathcal{F}$ be the same as in Proposition 2.9. Let the notation $T$ be the same as in Proposition 2.4. Suppose that $\operatorname{char}(A)=p>0$. Then the sequence
$\bigoplus_{x\in\operatorname{Spec}(A)^{(s-2)}}\operatorname{H}^{T+s-1}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{p}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(s-1)}}\operatorname{H}^{T+s}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{p}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(A)^{(s)}}\operatorname{H}^{T+s+1}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{p}}\to 0$
is exact.
Proof.
By Corollary 2.5, $\operatorname{H}_{\mathrm{\acute{e}t}}^{m}(A,\mathcal{F})$ is a torsion group for $m\geq 1+T+(s-1)=s+T$. So the statement follows from Proposition 2.9 and Remark 2.8. ∎
Condition 2.11.
Let $\mathcal{V}$ be the category whose objects are normal crossing varieties over the spectrum of a field $k$ and whose morphisms are morphisms of schemes. Let $X$ be an object in $\mathcal{V}$, $\mathcal{F}_{X}$ an étale sheaf of abelian groups on $X$ and $f:X\to Y$ a morphism in $\mathcal{V}$. Let $N$ be a non-negative integer. Then $(\mathcal{F}_{X},N)\in\mathbb{S}_{X_{\mathrm{\acute{e}t}}}\times\mathbb{N}_{\geq 0}$ satisfies the followings:
  1. 1.
    We have a pull-back map
    $f^{*}:\mathcal{F}_{Y}\to f_{*}\mathcal{F}_{X}$
    which satisfies the following properties:
    1. (i)
      For two morphisms $f:X\to Y$ and $g:Y\to Z$ in $\mathcal{V}$, the composite map
      $\mathcal{F}_{Z}\xrightarrow[g^{*}]{}g_{*}\mathcal{F}_{Y}\xrightarrow[g_{*}(f^{*})]{}(g\circ f)_{*}\mathcal{F}_{X}$
      agrees with the pull-back map $(g\circ f)^{*}$.
    2. (ii)
      Let $X$ be an object in $\mathcal{V}$ and $j:U\to X$ an open immersion. Then the homomorphism
      $j^{*}\mathcal{F}_{X}\to\mathcal{F}_{U}$
      is an isomorphism where the homomorphism is induced by the pull-back map $j^{*}$.
    3. (iii)
      Let $R=\mathcal{O}_{X,x}^{h}$ be the henselization of the local ring of a normal crossing variety $X$ at a point $x\in X$ and a closed immersion $i:Z\to\operatorname{Spec}(R)$ be a morphism in $\mathcal{V}$. Then the homomorphism
      $\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(Z,i^{*}\mathcal{F}_{\operatorname{Spec}(R)})\to\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(Z,\mathcal{F}_{Z})$
      is an isomorphism for $N^{\prime}\geq N$ where the homomorphism is induced by the pull-back map $i^{*}$.
  2. 2.
    Let $X$ be a smooth scheme over $\operatorname{Spec}(k)$ and $R$ the henselization of the local ring $\mathcal{O}_{X,x}$ at a point $x\in X$. Then the sequence
    $0\to\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(R,\mathcal{F}_{\operatorname{Spec}(R)})\to\bigoplus_{x\in\operatorname{Spec}(R)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(R_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(R)})$
    $\to\bigoplus_{x\in\operatorname{Spec}(R)^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}(R_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(R)})\to\cdots$
    is exact for $N^{\prime}\geq N$.
Remark 2.12.
Let $\mathcal{V}$ be the same as in the above and $X$ an object in $\mathcal{V}$.
  1. 1.
    Let $R$ be the henselization of the local ring of $X$ at a point $x\in X$ and $i:Z\to\operatorname{Spec}(R)$ a closed immersion which is a morphism in $\mathcal{V}$. If $(\mathcal{F}_{X},N)\in\mathbb{S}_{X_{\mathrm{\acute{e}t}}}\times\mathbb{N}_{\geq 0}$ satisfies Condition 2.11.1, then we have an isomorphism
    $\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(R,\mathcal{F}_{\operatorname{Spec}(R)})\xrightarrow{\sim}\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(Z,\mathcal{F}_{Z})$
    for $N^{\prime}\geq N$ by the isomorphism (14).
  2. 2.
    In the following cases, Condition 2.11 is satisfied:
    1. (i)
      Let $l$ be an integer which is prime to $\operatorname{char}(k)$ and $n$ a non-negative integer. If $\mathcal{F}_{X}=\mu_{l}^{\otimes n}$ and $N=0$, then $(\mathcal{F}_{X},N)\in\mathbb{S}_{X_{\mathrm{\acute{e}t}}}\times\mathbb{N}_{\geq 0}$ satisfies Condition 2.11 by [4, Corollary 2.2.2].
    2. (ii)
      Suppose that $\operatorname{char}(k)=p>0$. If $A^{\prime}$ is a commutative ring over $k$, then $\operatorname{cd}_{p}(A^{\prime})=1$.
      Let $r$, $n$ be non-negative integers. If $\mathcal{F}_{X}=\lambda_{X,r}^{n}$ and $N=1$, then $(\mathcal{F}_{X},N)\in\mathbb{S}_{X_{\mathrm{\acute{e}t}}}\times\mathbb{N}_{\geq 0}$ satisfies Condition 2.11.1 by [26, p.732, Theorem 3.5.1], [26, p.734, Corollary 3.5.3] and [24, p.30, Theorem 2.8]. Moreover, $(\mathcal{F}_{X},N)\in\mathbb{S}_{X_{\mathrm{\acute{e}t}}}\times\mathbb{N}_{\geq 0}$ satisfies Condition 2.11.2 by [29, p.600, Theorem 4.1].
In order to give another example which satisfies Condition 2.11.2, we recall the followings:
Lemma 2.13.
Let $X$ be essentially of finite type over the spectrum of a regular ring of dimension at most $1$, $n$ an integer and $\mathbb{Z}(n)$ Bloch’s cycle complex for the Zariski topology. Let $i:Z\to X$ be a closed immersion of codimension $c$ with open complement $j:U\to X$, $i_{1}:W\to X$ a closed immersion of codimension $c^{\prime}$ and
${Z_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i^{\prime}}$$\scriptstyle{i_{2}}$${W\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{1}}$${U_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j^{\prime}}$$\scriptstyle{i_{3}}$${Z\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${X}$${U\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j}$
Cartesian products.
Then the natural inclusion map
$i_{*}:i_{*}\mathbb{Z}(n-c)^{Z}[-2c]\to\mathbb{Z}(n)^{X}$
induces a quasi-isomorphism
$\operatorname{Gys}_{i}:\mathbb{Z}(n-c)^{Z}[-2c]\to Ri^{!}\mathbb{Z}(n)^{X}.$
Moreover,
$\left((i^{\prime})_{*}(\operatorname{Gys}_{i\circ i_{2}}),\operatorname{Gys}_{i_{1}},R(j^{\prime})_{*}(\operatorname{Gys}_{i_{3}})\right)$
is a morphism of triangles, that is, we have a commutative diagram
$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${(i^{\prime})_{*}\mathbb{Z}(n-c^{\prime\prime})^{Z_{1}}[-2c^{\prime\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(i^{\prime})_{*}}$$\scriptstyle{(i^{\prime})_{*}(\operatorname{Gys}_{i\circ i_{2}})}$${\mathbb{Z}(n-c^{\prime})^{W}[-2c^{\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\operatorname{Gys}_{i_{1}}}$${R(j^{\prime})_{*}\mathbb{Z}(n-c^{\prime})^{U_{1}}[-2c^{\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{R(j^{\prime})_{*}(\operatorname{Gys}_{i_{3}})}\textstyle{\cdots}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(i^{\prime})_{*}R(i\circ i_{2})^{!}\mathbb{Z}(n)^{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R(i_{1})^{!}\mathbb{Z}(n)^{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R(j^{\prime})_{*}R(i_{3})^{!}\mathbb{Z}(n)^{U}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots}$
where $c^{\prime\prime}=\operatorname{codim}(Z_{1},X)$.
Proof.
By [7, p.780, Corollary 3.3.a)] and $i^{!}j_{*}=0$ (cf.[17, p.76, II, Proposition 3.14 (c)]), $\operatorname{Gys}_{i}$ is an isomorphism. Since we have a commutative diagram
${(i\circ i_{2})_{*}\mathbb{Z}(n-c^{\prime\prime})^{Z_{1}}[-2c^{\prime\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(i_{1})_{*}((i^{\prime})_{*})}$$\scriptstyle{i_{*}((i_{2})_{*})}$${(i_{1})_{*}\mathbb{Z}(n-c^{\prime})^{W}[-2c^{\prime}]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(i_{1})_{*}}$${i_{*}\mathbb{Z}(n-c)^{Z}[-2c]\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{*}}$${\mathbb{Z}(n)^{X}}$
by the functoriality of the natural inclusion map, there exists a morphism
$h:Rj_{*}(i_{3})_{*}\mathbb{Z}(n-c^{\prime})^{U_{1}}\to Rj_{*}\mathbb{Z}(n)^{U}$
so that $\left(i_{*}((i_{2})_{*}),(i_{1})_{*},h\right)$ is a morphism of triangles by [7, p.780, Corollary 3.3.a)]. Moreover, we have
$j^{*}(h)=j^{*}((i_{1})_{*})$ and $h=Rj_{*}(i_{3})_{*}$
by $j^{*}i_{*}=0$ (cf.[17, p.76, II, Proposition 3.14 (c)]). So $\left(i_{*}((i_{2})_{*}),(i_{1})_{*},Rj_{*}(i_{3})_{*}\right)$ is a morphism of triangles. Moreover, we have
$R(i_{1})^{!}i_{*}=(i^{\prime})_{*}R(i_{2})^{!}$ and $R(i_{1})^{!}Rj_{*}=R(j^{\prime})_{*}R(i_{3})^{!}.$
So $\left((i^{\prime})_{*}(\operatorname{Gys}_{i_{2}}),\operatorname{Gys}_{i_{1}},R(j^{\prime})_{*}(\operatorname{Gys}_{i_{3}})\right)$ is a morphism of triangles. Since $i_{*}$ is a left adjoint functor to $Ri^{!}$, we have a commutative diagram
${i_{*}\mathbb{Z}(n-c)^{Z}[-2c]\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i_{*}}$$\scriptstyle{i_{*}(\operatorname{Gys}_{i})}$${\mathbb{Z}(n)^{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${i_{*}Ri^{!}\mathbb{Z}(n)^{X}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}(n)^{X}}$
and so we are able to prove that $\left(i_{*}(\operatorname{Gys}_{i}),\operatorname{id}_{\mathbb{Z}(n)},\operatorname{id}_{Rj_{*}\mathbb{Z}(n)}\right)$ is a morphism of triangles as above. Moreover, we have
$\operatorname{Gys}_{i\circ i_{2}}=R(i_{2})^{!}(\operatorname{Gys}_{i})\circ\operatorname{Gys}_{i_{2}}$
by the functoriality of the natural inclusion map. Hence
$\left((i^{\prime})_{*}(\operatorname{Gys}_{i\circ i_{2}}),\operatorname{Gys}_{i_{1}},R(j^{\prime})_{*}(\operatorname{Gys}_{i_{3}})\right)$
is a morphism of triangles. This completes the proof. ∎
Proposition 2.14.
Let $X$ be a smooth scheme over the spectrum of a field $k$, AA the local ring $\mathcal{O}_{X,y}$ of $X$ at a point $y\in X$ and $x\in\operatorname{Spec}(A)^{(s)}$. Let $n$ be an integer. Let $\mathbb{Z}(n)$ (resp. $\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$) be Bloch’s cycle complex for the Zariski (resp. étale) topology and $\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}=\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\otimes_{\mathbb{Z}}\mathbb{Q}/\mathbb{Z}$. Then we have an isomorphism
$\operatorname{H}^{i}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})=\left\{\begin{array}[]{ll}\operatorname{H}^{i}_{x}(A_{\mathrm{Zar}},\mathbb{Z}(n))&\textrm{if $i\leq s+n+1$}\\[5.0pt] \operatorname{H}^{i-1}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}})&\textrm{if $i>s+n+1$}.\end{array}\right$. (21)
Moreover, the sequence
$0\to\operatorname{H}^{i}_{\mathrm{\acute{e}t}}(A,\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{i}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})$
$\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{i+1}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to\cdots$
is exact for any integer $i$.
Proof.
By [2, p.299, Theorem (10.1)], [7, p.779, Theorem 3.2.b)] and Lemma 2.13, the sequence
$0\to\operatorname{H}^{i}_{\mathrm{Zar}}(A,\mathbb{Z}(n))\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{i}_{x}(A_{\mathrm{Zar}},\mathbb{Z}(n))$
$\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{i+1}_{x}(A_{\mathrm{Zar}},\mathbb{Z}(n))\to\cdots$ (22)
is exact for any integer $i$. If $\operatorname{char}(k)=p\geq 0$, then we have a quasi-isomorphism
$\mathbb{Z}/m\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\simeq\left\{\begin{array}[]{ll}\mu_{m}^{\otimes n}&\textrm{if $(m,p)=1$}\\[5.0pt] \nu_{r}^{n}&\textrm{if $m=p^{r}$, $r>0$}\end{array}\right$.
by [7, p.774, Theorem 1.2.4] and [9, p.491, Theorem 8.3]. So the sequence
$0\to\operatorname{H}^{i}_{\mathrm{\acute{e}t}}(A,\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{i}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}})$
$\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{i+1}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to\cdots$ (23)
is exact for any integer $i$ by [4, Corollary 2.2.2] and [29, p.600, Theorem 4.1]. Since we have a distinguished triangle
$\cdots\to\mathbb{Z}(n)\xrightarrow{\times m}\mathbb{Z}(n)\to\mathbb{Z}/m(n)\to\cdots$
and an isomorphism
$\mathbb{Z}/m(n)\simeq\frac{1}{m}\mathbb{Z}/\mathbb{Z}(n)$
for any positive integer $m$, we have
$\operatorname{H}^{i}_{\mathrm{\acute{e}t}}(A,\mathbb{Q}(n)_{\mathrm{\acute{e}t}})=\operatorname{H}^{i}_{\mathrm{Zar}}(A,\mathbb{Q}(n))=\operatorname{H}^{i}_{\mathrm{Zar}}(A,\mathbb{Q}/\mathbb{Z}(n))=0$
for $i\geq n+1$ by [7, p.781, Proposition 3.6] and [7, p.786, Corollary 4.4]. Hence we have an isomorphism
$\operatorname{H}^{i}_{\mathrm{\acute{e}t}}(A,\mathbb{Z}(n)_{\mathrm{\acute{e}t}})=\left\{\begin{array}[]{ll}\operatorname{H}^{i}_{\mathrm{Zar}}(A,\mathbb{Z}(n))&\textrm{if $i\leq n+1$}\\[5.0pt] \operatorname{H}^{i-1}_{\mathrm{\acute{e}t}}(A,\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}})&\textrm{if $i>n+1$}\end{array}\right$. (26)
by [7, p.774, Theorem 1.2.2] and [30]. In order to prove the statement, it suffices to prove the isomorphism (21) in the case where $\operatorname{dim}(A)=s$. We prove the isomorphism (21) by induction on $\operatorname{dim}(A)=s$.
In the case where $\operatorname{dim}(A)=s=0$, the statement follows from (26).
Assume that $\operatorname{dim}(A)=s=1$. Since the statement holds in the case where $\operatorname{dim}(A)=s=0$, the homomorphism
$\operatorname{H}^{i}_{\mathrm{\acute{e}t}}(A,\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{i}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})$
is injective by (2.2), (2.2) and (26). So the homomorphism
$\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{i}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{i+1}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})$
is surjective and the sequence
$\operatorname{H}^{i}_{\mathrm{\acute{e}t}}(A,\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{i}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{i+1}_{x}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\to 0$
is exact. Hence the statement holds in the case where $\operatorname{dim}(A)=s=1$ by (2.2), (2.2) and (26).
Assume that the statement holds in the case where $\operatorname{dim}(A)=s\leq t$ and $t\geq 1$. Suppose that $\operatorname{dim}(A)=s=t+1$. Consider the spectral sequence
$E_{1}^{u,v}=\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(u)}}\operatorname{H}_{x}^{u+v}(A_{\mathrm{\acute{e}t}},\mathbb{Z}(n)_{\mathrm{\acute{e}t}})\Rightarrow E^{u+v}=\operatorname{H}^{u+v}_{\mathrm{\acute{e}t}}(A,\mathbb{Z}(n)_{\mathrm{\acute{e}t}}).$
By (2.2), (2.2) and the assumption, it suffices to prove that the sequence
$E_{1}^{t-1,i}\to E_{1}^{t,i}\to E_{1}^{t+1,i}\to 0$
is exact for any integer $i$, that is,
$E^{e,i}_{2}=0$ (27)
for any integer $i$ and $t\leq e\leq t+1$. By (2.2), (2.2), (26) and the assumption, we have
$E^{u,i}_{2}=\begin{cases}E^{i}&(\textrm{if}~~u=0)\\ \\ 0&(\textrm{if}~~e\geq 3~~\textrm{and}~~0<u\leq e-2)\end{cases}$
for any integer $i$. So the isomorphism (27) follows from Lemma 2.1. This completes the proof. ∎
Remark 2.15.
If $\mathcal{F}_{X}=\mathbb{G}_{m,X}$ and $N=1$, then $(\mathcal{F}_{X},N)\in\mathbb{S}_{X_{\mathrm{\acute{e}t}}}\times\mathbb{N}_{\geq 0}$ satisfies Condition 2.11 by [17, p.116, III, Remark 3.11 (a)] and Proposition 2.14.
Moreover, let us remark the following:
Lemma 2.16.
Let $f:X\to Y$ and $g:Y\to Z$ be morphisms of schemes. Let $\mathcal{H}\in\mathbb{S}(Z_{\mathrm{\acute{e}t}})$, $\mathcal{G}\in\mathbb{S}(Y_{\mathrm{\acute{e}t}})$ and $\mathcal{F}\in\mathbb{S}(X_{\mathrm{\acute{e}t}})$. Let
$\alpha\in\operatorname{Hom}_{\mathbb{S}(Y_{\mathrm{\acute{e}t}})}(g^{*}(\mathcal{H}),\mathcal{G}),$ $\beta\in\operatorname{Hom}_{\mathbb{S}(X_{\mathrm{\acute{e}t}})}(f^{*}(\mathcal{G}),\mathcal{F})$
and
$\gamma\in\operatorname{Hom}_{\mathbb{S}(X_{\mathrm{\acute{e}t}})}((g\circ f)^{*}(\mathcal{H}),\mathcal{F}).$
Let
$\alpha^{\prime}\in\operatorname{Hom}_{\mathbb{S}(Z_{\mathrm{\acute{e}t}})}(\mathcal{H},g_{*}(\mathcal{G})),$ $\beta^{\prime}\in\operatorname{Hom}_{\mathbb{S}(Y_{\mathrm{\acute{e}t}})}(\mathcal{G},f_{*}\mathcal{F})$
and
$\gamma^{\prime}\in\operatorname{Hom}_{\mathbb{S}(Z_{\mathrm{\acute{e}t}})}(\mathcal{H},(g\circ f)_{*}\mathcal{F})$
be the right adjoints of $\alpha$, $\beta$ and $\gamma$. Then the following are equivalent:
  1. (i)
    $\beta\circ f^{*}(\alpha)$ agrees with $\gamma$.
  2. (ii)
    $g_{*}(\beta^{\prime})\circ\alpha^{\prime}$ agrees with $\gamma^{\prime}$.
Proof.
Since the proof of the implication (ii) $\Rightarrow$ (i) is similar to that of the implication (i) $\Rightarrow$ (ii), it suffices to show that (i) implies (ii). We have the commutative diagram
$\mathcal{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${g_{*}g^{*}\mathcal{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${g_{*}\mathcal{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(g\circ f)_{*}(g\circ f)^{*}\mathcal{H}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${g_{*}(f_{*}f^{*})\mathcal{G}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(g\circ f)_{*}\mathcal{F}}$
and the composition of the lower sequence agrees with $(g\circ f)_{*}\gamma$ by (i). So (ii) holds and the statement follows. ∎
Lemma 2.17.
Let $Y$ be a normal crossing variety over the spectrum of a field $k$, AA the henselization of the local ring $\mathcal{O}_{Y,y}$ at a point $y\in Y$ and $\mathcal{F}_{X}$ an étale sheaf on an object $X$ in $\mathcal{V}$ which satisfies Condition 2.11. Then the homomorphism
$\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})\to\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(\kappa(x),\mathcal{F}_{\operatorname{Spec}(\kappa(x))})$
is injective for $x\in\operatorname{Spec}(A)^{(0)}$ and $N^{\prime}\geq N$.
Proof.
By Condition 2.11, we have a commutative diagram
${\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(\kappa(x),\mathcal{F}_{\operatorname{Spec}(\kappa(x))})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A/x,\mathcal{F}_{\operatorname{Spec}(A/x)})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(\kappa(x),\mathcal{F}_{\operatorname{Spec}(\kappa(x))}).}$
Then the left map is an isomorphism by Condition 2.11. Moreover, the lower map is injective by Condition 2.11 and Lemma 2.16. So the statement follows. ∎
Lemma 2.18.
Let $Y$ be a normal crossing variety over the spectrum of a field $k$, $i:Z\to Y$ a closed immersion with a normal crossing variety $Z$ over the spectrum of a field $k$ and $C$ the henselization of the local ring $\mathcal{O}_{Z,z}$ at a point $z\in Z$.
Let $(\mathcal{F}_{X},N)$ be a pair of an étale sheaf on an object in $\mathcal{V}$ and a non-negative integer which satisfies Condition 2.11. Let $q$ be a prime number. Put
$T:=\left\{\begin{array}[]{ll}0&\textrm{if $\mathcal{F}_{Y}$ is a sheaf of torsion groups}\\[5.0pt] 1&\textrm{otherwise}\end{array}\right$.
and $N_{1}=\operatorname{max}\{N,T\}$. Suppose that the sequence
$0\to\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(C)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}$
$\to\bigoplus_{x\in\operatorname{Spec}(C)^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\cdots$ (28)
is exact for any point $z\in Z$ and $N^{\prime}\geq N_{1}+1$. Then we have an isomorphism
$\operatorname{H}^{N^{\prime}+s}_{x}(Z_{\mathrm{\acute{e}t}},i^{*}\mathcal{F}_{Y})_{T_{q}}\xrightarrow{\simeq}\operatorname{H}^{N^{\prime}+s}_{x}(Z_{\mathrm{\acute{e}t}},\mathcal{F}_{Z})_{T_{q}}$ (29)
for $N^{\prime}\geq N_{1}$ and $x\in Y^{(s)}\cap Z$.
Proof.
Let $B$ be the henselization of the local ring $\mathcal{O}_{Y,z}$ at a point $z\in Y\cap Z$. By [17, p.93, III, Corollary 1.28], we may prove the statement in the case where $Z$ is replaced by $\operatorname{Spec}(C)$ with $\operatorname{dim}(C)=s$ and $Y$ is replaced by $\operatorname{Spec}(B)$. We prove the statement by induction on $s$.
In the case where $s=0$, we have isomorphisms
$\operatorname{H}^{N^{\prime}}_{x}(C_{\mathrm{\acute{e}t}},i^{*}\mathcal{F}_{\operatorname{Spec}(B)})\simeq\operatorname{H}^{N^{\prime}}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})\simeq\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(\kappa(x),\mathcal{F}_{\operatorname{Spec}(\kappa(x))})$
for $x\in\operatorname{Spec}(C)^{(0)}$ and $N^{\prime}\geq N$. So the statement holds in this case.
In the case where $s=1$, if $\mathcal{F}^{\prime}$ is $i^{*}\mathcal{F}_{\operatorname{Spec}(B)}$ or $\mathcal{F}_{\operatorname{Spec}(C)}$, then the sequence
$0\to\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(C,\mathcal{F}^{\prime})\to\bigoplus_{x\in\operatorname{Spec}(C)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}^{\prime})$
$\to\bigoplus_{x\in\operatorname{Spec}(C)^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}^{\prime})\to 0$
is exact for $N^{\prime}\geq N$ by Condition 2.11 and Lemma 2.17. Since the statement holds in the case where $s=0$, the statement follows from Condition 2.11.
Assume that the statement holds in the case where $s\leq s^{\prime}$ and $s^{\prime}\geq 1$. Suppose that $\operatorname{dim}(C)=s^{\prime}+1$. Let $\mathcal{F}^{\prime}$ be $i^{*}\mathcal{F}_{\operatorname{Spec}(B)}$ or $\mathcal{F}_{\operatorname{Spec}(C)}$. Let $N^{\prime}\geq N_{1}$. By the assumption, we have the isomorphism (29) for $x\in\operatorname{Spec}(C)^{(s)}$ and $s\leq s^{\prime}$. So it suffices to show that the sequence
$\bigoplus_{x\in\operatorname{Spec}(C)^{(s^{\prime}-1)}}\operatorname{H}^{N^{\prime}+s^{\prime}-1}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}^{\prime})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(C)^{(s^{\prime})}}\operatorname{H}^{N^{\prime}+s^{\prime}}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}^{\prime})_{T_{q}}$
$\to\bigoplus_{x\in\operatorname{Spec}(C)^{(s^{\prime}+1)}}\operatorname{H}^{N^{\prime}+s^{\prime}+1}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}^{\prime})_{T_{q}}\to 0$
is exact. Consider the spectral sequence
$E^{u,v}_{1}(\mathcal{G})=\bigoplus_{x\in\operatorname{Spec}(C)^{(u)}}\operatorname{H}^{u+v}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{G})\Rightarrow E^{u+v}(\mathcal{G})=\operatorname{H}^{u+v}_{\mathrm{\acute{e}t}}(C,\mathcal{G})$
for $\mathcal{G}\in\mathbb{S}(\operatorname{Spec}(C)_{\mathrm{\acute{e}t}})$. Then we have isomorphisms
$E^{u,N^{\prime}+1}_{2}(i^{*}\mathcal{F}_{\operatorname{Spec}(B)})_{T_{q}}\simeq E^{u,N^{\prime}+1}_{2}(\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\simeq\left\{\begin{array}[]{ll}\operatorname{H}^{N^{\prime}+1}_{\mathrm{\acute{e}t}}(C,\mathcal{F}^{\prime})_{T_{q}}&(u=0)\\[5.0pt] 0&(u\leq s^{\prime}-1)\end{array}\right$. (32)
by (2.18) and the assumption. Moreover, we have an isomorphism
$\operatorname{H}^{m}_{\mathrm{\acute{e}t}}(C,\mathcal{F}_{\operatorname{Spec}(C)})\simeq\operatorname{H}^{m}_{\mathrm{\acute{e}t}}(\kappa(z),i^{*}(\mathcal{F}_{\operatorname{Spec}(C)}))$
for $m\geq 0$. So $\operatorname{H}_{\mathrm{\acute{e}t}}^{N^{\prime}+s^{\prime}}(C,\mathcal{F}_{\operatorname{Spec}(C)})$ is a torsion group because $N^{\prime}+s^{\prime}\geq 1$. Hence the statement follows from Proposition 2.9. ∎
Moreover, let us remark the following facts:
Lemma 2.19.
Let $\mathcal{A}$ be a small abelian category. Then we have the followings:
  1. 1.
    Consider the following commutative diagram in $\mathcal{A}$:
    ${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${D_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f}$${0}$${A_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${D_{2}}$${A_{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${B_{3}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${C_{3}}$${0}$
    where all columns are exact and all rows are complex. Assume that the sequences
    $C_{1}\to D_{1}\to 0$
    $B_{2}\to C_{2}\to D_{2}$
    $A_{3}\to B_{3}\to C_{3}$
    are exact. Then the morphism $f:D_{1}\to D_{2}$ is injective.
  2. 2.
    Consider the following diagram in $\mathcal{A}$:
[…]
    where the sequences are exact. Then the morphism
    $\delta:\operatorname{Ker}(h)\to A_{1}$
    is defined by the commutative diagram. Moreover, assume that
    $\operatorname{Im}(\delta)\subset\operatorname{Ker}(f).$
    Then the morphism $g:A_{2}\to B_{2}$ is injective.
Proof.
By the Freyd-Mitchell Embedding Theorem, the statement follows by diagram chasing. ∎
Lemma 2.20.
Let
${Z\cap U\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i^{\prime}}$$\scriptstyle{j^{\prime}}$${U\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{j}$${Z\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{i}$${X}$
be Cartesian. If $j$ is an open immersion and $i$ is a closed immersion, then we have
$i_{*}(j^{\prime})_{!}$ $=j_{!}(i^{\prime})_{*}.$
In particular, we have
$i_{*}(j^{\prime})_{!}=j_{!}$
in the case where $U=Z\cap U\subset Z$.
Proof.
The statement follows from [17, p.71, II, Corollary 3.5] and the definition of the extension by zero (cf. [17, p.78]). ∎
Lemma 2.21.
Let $Y$ be a smooth scheme over the spectrum of a field $k$ and $Z$ a normal crossing variety over $\operatorname{Spec}(k)$ which is a closed scheme of $Y$ of codimension $1$.
Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y$ of $Y$, $C$ the henselization of the local ring $\mathcal{O}_{Z,y}$ and $i:\operatorname{Spec}(C)\to\operatorname{Spec}(A)$ the corresponding closed immersion.
Let $U=\operatorname{Spec}(A)\setminus\operatorname{Spec}(C)$ and $j:U\to\operatorname{Spec}(A)$ the corresponding open immersion.
Let $(\mathcal{F}_{X},N)$ be a pair of an étale sheaf on an object $X$ in $\mathcal{V}$ and a non-negative integer which satisfies Condition 2.11 and $q$ a prime number. Suppose that $N^{\prime}\geq\operatorname{max}\{N,T\}=N_{1}$ and $\operatorname{cd}_{q}(k)<\infty$.
If the sequence
$0\to\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(C)^{(0)}}\operatorname{H}^{t}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to$
$\bigoplus_{x\in\operatorname{Spec}(C)^{(1)}}\operatorname{H}^{t+1}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\cdots$ (33)
is exact for any point $y\in Y$ and $t\geq N^{\prime}$, then a distinguished triangle
$\cdots\to j_{!}\mathcal{F}_{U}\to\mathcal{F}_{\operatorname{Spec}(A)}\to i_{*}i^{*}\mathcal{F}_{\operatorname{Spec}(A)}\to\cdots$ (34)
induces an exact sequence
$0\to\operatorname{H}_{x}^{t+s-1}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\operatorname{H}_{x}^{t+s}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\operatorname{H}_{x}^{t+s}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\to 0$ (35)
for $x\in(\operatorname{Spec}(A))^{(s)}\cap\operatorname{Spec}(C)$ and the sequence
$0\to\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,j_{!}\mathcal{F}_{U})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{t}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to$
$\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{t+1}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\cdots$ (36)
is exact for $t\geq N^{\prime}$.
Proof.
We prove the exactness of the sequences (35) and (2.21) by downward induction on $N^{\prime}$. By [25, Lemma 3.7] and Lemma 2.18, we have an isomorphism
$\operatorname{H}^{u}_{x}(A_{\mathrm{\acute{e}t}},i_{*}i^{*}\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\simeq\left\{\begin{array}[]{ll}\operatorname{H}_{x}^{u}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}},&x\in(\operatorname{Spec}(A))^{(s)}\cap\operatorname{Spec}(C)\\[5.0pt] 0,&x\in(\operatorname{Spec}(A))^{(s)}\setminus\operatorname{Spec}(C)\end{array}\right$.
for $u\geq N^{\prime}+s$. So the distinguished triangle (34) induces an exact sequence
$\operatorname{H}^{N^{\prime}+s-1}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\operatorname{H}^{N^{\prime}+s}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\operatorname{H}^{N^{\prime}+s}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}$
$\to$ $\operatorname{H}^{N^{\prime}+s}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\cdots$
for $x\in(\operatorname{Spec}(A))^{(s)}$. Let $\mathcal{G}$ be a sheaf of abelian groups on $\operatorname{Spec}(A)_{\mathrm{\acute{e}t}}$ and $x\in\operatorname{Spec}(A)^{(s)}$. Then
$\operatorname{H}^{t+s}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{G})_{T_{q}}=0$
for sufficiently large $t$ by Proposition 2.7 and Remark 2.8. Moreover, we have
$\operatorname{H}^{t}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})=0$
for any $t\geq 0$ by the isomorphism (14). So we have the exact sequences (35) and (2.21) for sufficiently large $N^{\prime}$.
Assume that the sequences (35) and (2.21) are exact for $t>N^{\prime}(\geq N_{1})$. If the first map in the sequence (35) is injective for $t=N^{\prime}+1$, the second map in the sequence (35) is surjective for $t=N^{\prime}$. So, in order to prove that the sequence (35) is exact for $t=N^{\prime}$, it suffices to show that the first map in the sequence (35), that is, the homomorphism
$\operatorname{H}^{t+s-1}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\operatorname{H}^{t+s}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}$ (37)
is injective for $t=N^{\prime}$.
We prove the injectivity of the homomorphism (37) for $t=N^{\prime}$ by induction on $s$. By [17, p.93, III, Corollary 1.28], it suffices to prove the injectivity of the homomorphism (37) in the case where $A$ is a henselian regular local ring with $\operatorname{dim}(A)=s$.
In the case where $s=0$, $(\operatorname{Spec}(A))^{(0)}\cap\operatorname{Spec}(C)=\emptyset$. So the homomorphism (37) is injective.
In the case where $s=1$, we prove the injectivity of the homomorphism (37). Since $A$ is a henselian valuation ring, we have an isomorphism
$\operatorname{H}^{u}_{\mathrm{\acute{e}t}}\left(A,j_{!}\mathcal{F}_{\operatorname{Spec}(k(A))}\right)\simeq 0$
for any integer $u\geq 0$ by the isomorphism (14) and so we have an isomorphism
$\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(k(A),\mathcal{F}_{\operatorname{Spec}(k(A))})\xrightarrow{\simeq}\operatorname{H}^{N^{\prime}+1}_{x}(A,j_{!}\mathcal{F}_{\operatorname{Spec}(k(A))})$
for $x\in(\operatorname{Spec}(A))^{(1)}\cap\operatorname{Spec}(C)$. Since the diagram
${\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(k(A),\mathcal{F}_{\operatorname{Spec}(k(A))})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\simeq}$${\operatorname{H}^{N^{\prime}}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{N^{\prime}+1}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{\operatorname{Spec}(k(A))})}$ (38)
is anti-commutative by [25, Lemma 3.11], the homomorphism (37) is injective by Condition 2.11 and Lemma 2.17.
Next, we prove the injectivity of the homomorphism (37) in the case where $s=2$. Put
$E^{u,v}_{1}(D,\mathcal{G})=\displaystyle\bigoplus_{x\in\operatorname{Spec}(D)^{(u)}}\operatorname{H}^{u+v}_{x}(D_{\mathrm{\acute{e}t}},\mathcal{G})$
for an equidimensional ring $D$ and a sheaf $\mathcal{G}$ on $\operatorname{Spec}(D)_{\mathrm{\acute{e}t}}$. Then we have a commutative diagram
${0}$${E^{1,N^{\prime}}_{1}\left(C,\mathcal{F}_{\operatorname{Spec}(C)}\right)_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{2,N^{\prime}}_{1}\left(A,j_{!}\mathcal{F}_{U}\right)_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{2,N^{\prime}}_{1}\left(A,\mathcal{F}_{\operatorname{Spec}(A)}\right)_{T_{q}}}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{0,N^{\prime}}_{1}\left(C,\mathcal{F}_{\operatorname{Spec}(C)}\right)_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{1,N^{\prime}}_{1}\left(A,j_{!}\mathcal{F}_{U}\right)_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{1,N^{\prime}}_{1}\left(A,\mathcal{F}_{\operatorname{Spec}(A)}\right)_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{0,N^{\prime}}_{1}\left(A,j_{!}\mathcal{F}_{U}\right)_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{0,N^{\prime}}_{1}\left(A,\mathcal{F}_{\operatorname{Spec}(A)}\right)_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
where the columns are exact by Remark 2.6, the assumption and Proposition 2.9. Moreover, the middle row is exact by the above. So the homomorphism (37) is injective by the anti-commutative diagram (38) and Lemma 2.19.2.
Assume that the homomorphism (37) is injective for $s\leq u$ and $u\geq 2$. Then we prove the injectivity of the homomorphism (37) for $s=u+1$. Consider the following diagram:
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{u-2,N^{\prime}}_{1}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E^{u-1,N^{\prime}}_{1}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{u,N^{\prime}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${E_{1}^{u-2,N^{\prime}}(A,j_{!}\mathcal{F}_{U})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{u-1,N^{\prime}}(A,j_{!}\mathcal{F}_{U})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{u,N^{\prime}}(A,j_{!}\mathcal{F}_{U})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{u+1,N^{\prime}}(A,j_{!}\mathcal{F}_{U})_{T_{q}}}$${E_{1}^{u-2,N^{\prime}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{u-1,N^{\prime}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{u,N^{\prime}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}}$${0}$
where the upper row and the bottom row are exact by Condition 2.11 and the assumption. Moreover, the sequence
$E^{u-1,N^{\prime}}(A,j_{!}\mathcal{F}_{U})_{T_{q}}\to E^{u,N^{\prime}}(A,j_{!}\mathcal{F}_{U})_{T_{q}}\to E^{u+1,N^{\prime}}(A,j_{!}\mathcal{F}_{U})_{T_{q}}$
is exact by Remark 2.6 and Proposition 2.9. By the assumption, the columns except for the right column are also exact. So the homomorphism (37) is injective for $s=u+1$ by Lemma 2.19.1. Hence the sequence (35) is exact for $x\in(\operatorname{Spec}(A))^{(s)}\cap\operatorname{Spec}(C)$. Finally, we prove the exactness of the sequence (2.21). Since we have an anti-commutative diagram
${\displaystyle\bigoplus_{x\in\operatorname{Spec}(C)^{(0)}}\operatorname{H}_{x}^{N^{\prime}}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}_{x}^{N^{\prime}+1}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})}$${\operatorname{H}_{\mathrm{\acute{e}t}}^{N^{\prime}}(C,\mathcal{F}_{\operatorname{Spec}(C)})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}_{x}^{N^{\prime}}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
for $N^{\prime}\geq N$ by the anti-commutative diagram (38), the exactness of the sequence (2.21) follows from Condition 2.11 and the exact sequences (2.21) and (35). This completes the proof. ∎
We prove Theorem 1.2. The proof of the following is similar to the proof of [25, Theorem 1.3].
Theorem 2.22.
Let $Y$ be a normal crossing variety over the spectrum of a field $k$. Let $Y_{1},\cdots,Y_{a}$ be the irreducible components of $Y$.
Let $(\mathcal{F}_{X},N)$ be a pair of an étale sheaf of abelian groups on an object $X$ in $\mathcal{V}$ and a non-negative integer which satisfies Condition 2.11. Let the notation $T$ be the same as in Proposition 2.4 and $N_{1}=\operatorname{max}\{N,T\}$. Suppose that $\operatorname{cd}_{q}(k)<\infty$ for a prime number $q$.
Then we have the followings:
  • Property $\operatorname{P}_{1}(a)$: Let $s\geq 0$ be an integer and $i:Z\to Y$ a closed immersion with
    $Z=\displaystyle\bigcup^{a-1}_{m=1}Y_{m}.$
    Then the homomorphism
    $\operatorname{H}^{N^{\prime}+s}_{y}(Z_{\mathrm{\acute{e}t}},i^{*}\mathcal{F}_{Y})_{T_{q}}\xrightarrow{\sim}\operatorname{H}^{N^{\prime}+s}_{y}(Z_{\mathrm{\acute{e}t}},\mathcal{F}_{Z})_{T_{q}}$
    is an isomorphism for $N^{\prime}\geq N_{1}$ and $y\in Y^{(s)}\cap Z$.
  • Property $\operatorname{P}_{2}(a)$: Let the notations be the same as above and $j:U\hookrightarrow Y$ the immersion of the open complement of $Z$ in $Y$. Then a distinguished triangle
    $\cdots\to j_{!}\mathcal{F}_{U}\to\mathcal{F}_{X}\to i_{*}i^{*}\mathcal{F}_{X}\to\cdots.$ (39)
    induces an exact sequence
    $0\to\operatorname{H}^{N^{\prime}+s}_{y}(Y_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\operatorname{H}^{N^{\prime}+s}_{y}(Y_{\mathrm{\acute{e}t}},\mathcal{F}_{Y})_{T_{q}}\to\operatorname{H}^{N^{\prime}+s}_{y}(Z_{\mathrm{\acute{e}t}},\mathcal{F}_{Z})_{T_{q}}\to 0$ (40)
    for $y\in Y^{(s)}\cap Z$ and $N^{\prime}\geq N_{1}$.
  • Property $\operatorname{P}_{3}(a)$: Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point of $y\in Y$. Then the sequence
    $0\to\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}$
    $\to\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\to\cdots$
    is exact for $N^{\prime}\geq N_{1}$.
Proof.
We prove the statement by induction on $a$.
In the case where $a=1$, then $Z=\phi$ and so $\operatorname{P}_{1}(1)$ and $\operatorname{P}_{2}(1)$ hold. Since $Y$ is a smooth scheme over the spectrum of a field, $\operatorname{P}_{3}(1)$ holds by Condition 2.11.
Assume that $\operatorname{P}_{u}(v)$ holds for $u=1,2,3$ and $v\leq a$. Then we prove $\operatorname{P}_{u}(a+1)$ for $u=1,2,3$.
First, we prove $\operatorname{P}_{1}(a+1)$. By [17, p.93, III, Corollary 1.28], it suffices to prove $\operatorname{P}_{1}(a+1)$ in the case where $Y$ is the spectrum of a henselian local ring $A$ and $\operatorname{dim}(Y)=s$. Since $(Z^{(0)})=a$, $\operatorname{P}_{1}(a+1)$ follows from $\operatorname{P}_{3}(a)$ and Lemma 2.18.
Next we prove $\operatorname{P}_{2}(a+1)$ and $\operatorname{P}_{3}(a+1)$ by downward induction on $N^{\prime}$. Since $\operatorname{P}_{1}(a+1)$ holds, the distinguished triangle (39) induces an exact sequence
$\operatorname{H}_{y}^{N_{1}+s}(Y_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\operatorname{H}_{y}^{N_{1}+s}(Y_{\mathrm{\acute{e}t}},\mathcal{F}_{Y})_{T_{q}}\to\operatorname{H}_{y}^{N_{1}+s}(Z_{\mathrm{\acute{e}t}},\mathcal{F}_{Z})_{T_{q}}$
$\to$ $\operatorname{H}_{y}^{N_{1}+s+1}(Y_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\cdots$ (41)
for $y\in Y^{(s)}\cap Z$.
Let $x\in Y^{(s)}$ and $\mathcal{G}$ be a sheaf of abelian groups on $Y_{\mathrm{\acute{e}t}}$. Since $\operatorname{cd}_{q}(k)$ and $\operatorname{tr.deg}_{k}\mathcal{O}_{Y,y}$ are finite for any $y\in Y^{(0)}$, we have
$\operatorname{H}^{N^{\prime}+s}_{x}(Y_{\mathrm{\acute{e}t}},\mathcal{G})_{T_{q}}=0$
for sufficiently large $N^{\prime}$ by Proposition 2.7 and Remark 2.8. Moreover, we have
$\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}=0$
for sufficiently large $N^{\prime}$ by (17) and Remark 2.8. So the sequence (40) is exact and $\operatorname{P}_{3}(a+1)$ holds for sufficiently large $N^{\prime}$.
Assume that the sequence (40) is exact and $\operatorname{P}_{3}(a+1)$ holds for $N^{\prime}>N^{\prime\prime}(\geq N_{1})$. Then the sequence
$\operatorname{H}^{N^{\prime\prime}+s}_{y}(Y_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\operatorname{H}^{N^{\prime\prime}+s}_{y}(Y_{\mathrm{\acute{e}t}},\mathcal{F}_{Y})_{T_{q}}\to\operatorname{H}^{N^{\prime\prime}+s}_{y}(Z_{\mathrm{\acute{e}t}},\mathcal{F}_{Z})_{T_{q}}\to 0$
is exact for $y\in Y^{(s)}\cap Z$ by the exact sequence (2.2). In order to prove $\operatorname{P}_{2}(a+1)$, it suffices to show that the homomorphism
$\operatorname{H}^{N^{\prime\prime}+s}_{y}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U})_{T_{q}}\to\operatorname{H}^{N^{\prime\prime}+s}_{y}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}$ (42)
is injective in the case where $A$ is the henselization at a point $y$ of $Y^{(s)}\cap Z$ by [17, p.93, III, Corollary 1.28]. Since we have
$U=Y\setminus Z=Y_{a+1}\setminus Z\cap Y_{a+1},$
there is an open immersion $j^{\prime}:Y\setminus Z\to Y_{a+1}$. Let $i_{a+1}:Y_{a+1}\to Y$ be a closed immersion. Then we have
$j=i_{a+1}\circ j^{\prime}.$
Let $A_{a+1}$ be the henselization of $\mathcal{O}_{Y_{a+1},y}$ and $U^{\prime}=\operatorname{Spec}(A_{a+1})\times_{Y_{a+1}}U$. For notational simplicity, a base change of $j$ (resp. $j^{\prime}$, $i$) is also denoted by $j$ (resp. $j^{\prime}$, $i$), unless confusion arises. Then we have an isomorphism
$\operatorname{H}_{x}^{u}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})=\left\{\begin{array}[]{ll}\operatorname{H}^{u}_{x}((A_{a+1})_{\mathrm{\acute{e}t}},(j^{\prime})_{!}\mathcal{F}_{U^{\prime}})&\textrm{for}~~x\in(Y_{a+1})^{(s)}\\[5.0pt] 0&\textrm{for}~~x\in Y^{(s)}\setminus Y_{a+1}\end{array}\right$. (45)
for any $u\geq 0$ by Lemma 2.20 (or [17, p.169, V, Proposition 1.13]) and [25, Lemma 3.7]. By Condition 2.11 and $\operatorname{P}_{3}(a)$, the sequences
$0\to$ $\operatorname{H}_{\mathrm{\acute{e}t}}^{N^{\prime}}(A_{a+1},\mathcal{F}_{\operatorname{Spec}(A_{a+1})})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(A_{a+1})^{(0)}}\operatorname{H}^{N^{\prime}}_{x}((A_{a+1})_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A_{a+1})})_{T_{q}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(A_{a+1})^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}((A_{a+1})_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A_{a+1})})_{T_{q}}\to\cdots$
and
$0\to$ $\operatorname{H}_{\mathrm{\acute{e}t}}^{N^{\prime}}(C_{a+1},\mathcal{F}_{\operatorname{Spec}(C_{a+1})})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(C_{a+1})^{(0)}}\operatorname{H}^{N^{\prime}}_{x}((C_{a+1})_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C_{a+1})})_{T_{q}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(C_{a+1})^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}((C_{a+1})_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C_{a+1})})_{T_{q}}\to\cdots$
are exact for $N^{\prime}\geq N_{1}$ where $\operatorname{Spec}(C_{a+1})=\operatorname{Spec}(A_{a+1})\times_{Y}Z$. So the sequence
$0\to$ $\operatorname{H}_{\mathrm{\acute{e}t}}^{N^{\prime}}(A_{a+1},(j^{\prime})_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(A_{a+1})^{(0)}}\operatorname{H}^{N^{\prime}}_{x}((A_{a+1})_{\mathrm{\acute{e}t}},(j^{\prime})_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(A_{a+1})^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}((A_{a+1})_{\mathrm{\acute{e}t}},(j^{\prime})_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\to\cdots$
is exact for $y\in(Y_{a+1})^{(s)}$ and $N^{\prime}\geq N_{1}$ by Lemma 2.21. Hence the sequence
$0\to$ $\operatorname{H}_{\mathrm{\acute{e}t}}^{N^{\prime}}(A,j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\to\cdots$ (46)
is exact for $N^{\prime}\geq N_{1}$ by the isomorphism (45).
Then we prove that the homomorphism (42) is injective by induction on $s$. Put $\operatorname{Spec}(C)=\operatorname{Spec}(A)\times_{Y}Z$ and $U^{\prime}=\operatorname{Spec}(A)\times_{Y}U$. Remark that we have an isomorphism
$\operatorname{H}^{u}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})\xrightarrow{\sim}\operatorname{H}^{u}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})$
for $x\in U=Y\setminus Z$ and any integer $u\geq 0$.
In the case where $s=0$, then we have an isomorphism
$\operatorname{H}_{y}^{u}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})\xrightarrow{\simeq}\operatorname{H}_{y}^{u}(C_{\mathrm{\acute{e}t}},i^{*}\mathcal{F}_{\operatorname{Spec}(A)})$
for $y\in Y^{(0)}\cap Z$ and $u\geq 0$. So the homomorphism (42) is injective for $s=0$.
In the case where $s=1$, we have the commutative diagram
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{N^{\prime\prime}}_{\mathrm{\acute{e}t}}(A,j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{N^{\prime\prime}}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{N^{\prime\prime}+1}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${\operatorname{H}^{N^{\prime\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{N^{\prime\prime}}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{N^{\prime\prime}+1}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{N^{\prime\prime}}_{\mathrm{\acute{e}t}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{x\in\operatorname{Spec}(C)^{(0)}}\operatorname{H}^{N^{\prime\prime}}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}}$
where the upper and the middle rows are exact by the exact sequence (2.2). Since $Y_{1},\cdots,Y_{a}$ are the irreducible components of $Z$, the lower row is exact by $\operatorname{P}_{3}(a)$. Moreover, the left and the middle columns are exact by the distinguished triangle (39) and the assumption. Hence the morphism (42) is injective for $s=1$ by Lemma 2.19.
Let $t$ be an integer and $t\geq 1$. Assume that $\operatorname{P}_{2}(a+1)$ holds for $s\leq t$ and $\operatorname{dim}(A)=t+1$. Put
$E^{u,v}_{1}(D,\mathcal{G})=\displaystyle\bigoplus_{x\in\operatorname{Spec}(D)^{(u)}}\operatorname{H}^{u+v}_{x}(D_{\mathrm{\acute{e}t}},\mathcal{G})$
for an equidimensional ring $D$ and a sheaf $\mathcal{G}$ on $\operatorname{Spec}(D)_{\mathrm{\acute{e}t}}$. Then we have the commutative diagram
${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{t-1,N^{\prime\prime}}(A,j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{t,N^{\prime\prime}}(A,j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{t+1,N^{\prime\prime}}(A,j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${B_{t}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{t-1,N^{\prime\prime}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{t,N^{\prime\prime}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{t+1,N^{\prime\prime}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}}$${D_{t}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{t-1,N^{\prime\prime}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${E_{1}^{t,N^{\prime\prime}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}}$${0}$
where
$B_{t}=\begin{cases}\operatorname{H}^{N^{\prime\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}&(\textrm{if}~~t=1)\\ \\ E_{1}^{t-2,N^{\prime\prime}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}&(\textrm{if}~~t\geq 2)\end{cases}$
and
$D_{t}=\begin{cases}\operatorname{H}^{N^{\prime\prime}}_{\mathrm{\acute{e}t}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}&(\textrm{if}~~t=1)\\ \\ E_{1}^{t-2,N^{\prime\prime}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}&(\textrm{if}~~t\geq 2).\end{cases}$
Then the columns except for the right map are exact by Condition 2.11 and $\operatorname{P}_{2}(a+1)$ for $s\leq t$. If $\mathcal{G}=j_{!}\mathcal{F}_{U^{\prime}}$ or $\mathcal{G}=\mathcal{F}_{\operatorname{Spec}(A)}$, then the sequence
$E_{1}^{t-1,N^{\prime\prime}}(A,\mathcal{G})_{T_{q}}\to E_{1}^{t,N^{\prime\prime}}(A,\mathcal{G})_{T_{q}}\to E_{1}^{t+1,N^{\prime\prime}}(A,\mathcal{G})_{T_{q}}\to 0$
is exact by the exact sequence (2.2), Proposition 2.9 and $\operatorname{P}_{3}(a+1)$ for $N^{\prime}\geq N^{\prime\prime}+1$. Since $Y_{1},\cdots,Y_{a}$ are the irreducible components of $Z$, the lower row in the diagram (47) is exact by $\operatorname{P}_{3}(a)$. So the homomorphism (42) is injective for $s=t+1$ by Lemma 2.19. Hence $\operatorname{P}_{2}(a+1)$ holds.
Finally, we prove $\operatorname{P}_{3}(a+1)$. Let the notations $A$, $C$ and $U^{\prime}$ be the same as above. By (2.2) and $\operatorname{P}_{3}(a)$, the sequences
$0\to$ $\operatorname{H}_{\mathrm{\acute{e}t}}^{N^{\prime}}(A,j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}(A_{\mathrm{\acute{e}t}},j_{!}\mathcal{F}_{U^{\prime}})_{T_{q}}\to\cdots$
and
$0\to$ $\operatorname{H}_{\mathrm{\acute{e}t}}^{N^{\prime}}(C,\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(C)^{(0)}}\operatorname{H}^{N^{\prime}}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}$
$\to$ $\bigoplus_{x\in\operatorname{Spec}(C)^{(1)}}\operatorname{H}^{N^{\prime}+1}_{x}(C_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(C)})_{T_{q}}\to\cdots$
are exact for $N^{\prime}\geq N_{1}$. Since $A$ is henselian, we have isomorphisms
$\operatorname{H}^{u}_{\mathrm{\acute{e}t}}(A,j_{!}\mathcal{F}_{U^{\prime}})=0$
for any $u\geq 0$ and
$\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})\xrightarrow{\sim}\operatorname{H}^{N^{\prime}}_{\mathrm{\acute{e}t}}(C,\mathcal{F}_{\operatorname{Spec}(C)})$
for $N^{\prime}\geq N$ by Condition 2.11 and the isomorphism (14). Hence $\operatorname{P}_{3}(a+1)$ follows from $\operatorname{P}_{2}(a+1)$. This completes the proof. ∎
Corollary 2.23.
Let the notations be the same as in Theorem 2.22. Suppose that $N_{1}=0$. Then the sequence
$0\to\operatorname{H}^{t}_{\mathrm{\acute{e}t}}(A,\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{t}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}$
$\to\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{t+1}_{x}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})_{T_{q}}\to\cdots$
is exact for any integer $t$. Especially, we have
$\operatorname{H}_{x}^{s+u}(A_{\mathrm{\acute{e}t}},\mathcal{F})_{T_{q}}=0$
for $x\in\operatorname{Spec}(A)^{(s)}$ and $u<0$.
Proof.
Consider the spectral sequence
$E^{u,v}_{1}=\bigoplus_{x\in\operatorname{Spec}(A)^{(u)}}\operatorname{H}_{x}^{u+v}(A_{\mathrm{\acute{e}t}},\mathcal{F}_{\operatorname{Spec}(A)})\Rightarrow E^{u+v}=\operatorname{H}_{\mathrm{\acute{e}t}}^{u+v}(A,\mathcal{F}_{\operatorname{Spec}(A)}).$
By Theorem 2.22 and [17, p.93, III, Corollary 1.28], it suffices to prove that
$(E^{s,t}_{1})_{T_{q}}=0$ (48)
for $t<0$ and $\operatorname{dim}(A)=s$. We prove the statement by induction on $s$.
If $s=1$, then the isomorphism (48) for t=1t=-1 follows from Lemma 2.17. So the isomorphism (48) holds for $s=1$ and $t<0$.
Assume that we have the isomorphism (48) for $s\leq s^{\prime}$ and $s^{\prime}\geq 1$. Suppose that $\operatorname{dim}(A)=s^{\prime}+1$. Then it suffices to prove that
$(E_{2}^{s^{\prime}+1,t})_{T_{q}}=0$
for $t<0$ by the assumption.
By Lemma 2.17, the homomorphism
$E^{s^{\prime}+t+1}\to E^{0,s^{\prime}+t+1}_{1}$
is injective. Moreover, we have
$(E^{u,v}_{2})_{T_{q}}=\begin{cases}(E^{u+v})_{T_{q}}&(\textrm{if}~~u=0)\\ \\ 0&(\textrm{if}~~s^{\prime}\geq 2~~\textrm{and}~~0<u\leq s^{\prime}-1)\end{cases}$
for $u+v=s^{\prime}+t$. Hence, the statement follows from Lemma 2.2. This completes the proof.
Lemma 2.24.
Let $A$ be the strict henselian of a local ring of a normal crossing variety over the spectrum of a field of positive characteristic. Then $A$ is a direct limit of a family of the strict henselian rings $A_{i}$ ($i\in I$) of a local ring of a normal crossing variety over the spectrum of the prime field which satisfies the following:
Let $A/(a_{s})$ $(s\in S)$ be the irreducible components of $A$. Then
$\operatorname{Im}(A_{i}\to A\to A/(a_{s}))$
$(s\in S)$ are regular for any $i\in I$.
Proof.
By the definition of a normal crossing variety $Y$, there exists an étale map of local rings $\mathcal{O}_{Y,y}\to A^{\prime}$ such that $A^{\prime}$ is also a quotient of a local ring $C$ of a smooth algebra over a field at a point and $\operatorname{dim}(C)=\operatorname{dim}(A^{\prime})+1$. By Quillen’s method (cf.[20, §7, The proof of Theorem 5.11]), there exists a local ring $C^{\prime}$ of a regular algebra of finite type over a field $k^{\prime}$ such that $k^{\prime}$ is finitely generated over the prime field and $C=k\otimes_{k^{\prime}}C^{\prime}.$ Then we have isomorphism
$C=\displaystyle\lim_{\begin{subarray}{c}\to\\ i\in I\end{subarray}}k_{i}\otimes_{k^{\prime}}C^{\prime}$
where $k_{i}$ run over subfields of $k$ which contain $k^{\prime}$ and are finitely generated over the prime field. Let $\mathfrak{m}$ be the maximal ideal of $C$, $f_{i}:k_{i}\otimes_{k^{\prime}}C^{\prime}\to C$ the base change and $\mathfrak{m}_{i}=(f_{i})^{-1}(\mathfrak{m}).$ Put $C_{i}=\left(k_{i}\otimes_{k^{\prime}}C^{\prime}\right)_{\mathfrak{m}_{i}}.$ Then $C_{i}$ is regular by [16, p.182, Theorem 23.7 (i)]. Let $c^{\prime}_{s}$ ($s\in S$) be elements of $C$ such that $C/(c^{\prime}_{s})\simeq A^{\prime}/(a^{\prime}_{s})$ where $A^{\prime}/(a^{\prime}_{s})$ ($s\in S$) are the irreducible components of $A^{\prime}$. Since we are able to choose a directed subset $J$ of $I$ such that $c_{s}^{\prime}\in\mathfrak{m}_{j}$ and
$\mathfrak{m}_{j}/\mathfrak{m}_{j}^{2}\otimes_{c_{j}/\mathfrak{m}_{j}}C/\mathfrak{m}\simeq\mathfrak{m}/\mathfrak{m}^{2}$
for any $s\in S$ and $j\in J$, we may assume that $I=J$ and $\operatorname{dim}(C)=\operatorname{dim}(C_{i})$ for any $i\in I$. Put
$A_{i}^{\prime}=\operatorname{Im}\left(C_{i}\to C\to A^{\prime}\right)$
for any $i\in I$ and $A$ (resp. $A_{i}$) denotes the strict henselian of $A^{\prime}$ (resp. $A^{\prime}_{i}$). Then $A$ is also the strict henselian of $\mathcal{O}_{Y,y}$ and $A_{i}$ satisfies the property of the statement. This completes the proof. ∎
Proposition 2.25.
Let $Y$ be a normal crossing variety over the spectrum of a field of positive characteristic $p>0$ and $y\in Y^{(s)}$. Then we have
$\operatorname{H}_{y}^{t}(Y_{\mathrm{\acute{e}t}},\lambda_{r}^{n})=0$
for $t<s$.
Proof.
Let $\{Y_{i}\}_{i\in I}$ be the irreducible components of $Y$. For an integer $m$, we define
$Y^{\langle m\rangle}:=\bigsqcup_{\{i_{1},i_{2},\cdots,i_{m}\}\subset I}Y_{i_{1}}\cap Y_{i_{2}}\cap\cdots\cap Y_{i_{m}}$
where the indices are pairwise distinct for each subset $\{i_{1},\cdots,i_{m}\}\subset I$. Then there is an exact sequence on $Y_{\mathrm{\acute{e}t}}$
$0\to\lambda^{n}_{Y,r}\to a_{1*}W_{r}\Omega^{n}_{Y^{\langle 1\rangle},\log}\to a_{2*}W_{r}\Omega^{n}_{Y^{\langle 2\rangle},\log}\to\cdots\to a_{N+1-n*}W_{r}\Omega^{n}_{Y^{\langle N+1-n\rangle},\log}\to 0$
by [26, p.727, Proposition 3.2.1] and Lemma 2.24. Here $N=\operatorname{dim}(Y)$ and $a_{m}$ is the canonical finite morphism $Y^{\langle m\rangle}\to Y$. Put
$\Lambda^{n,m}_{r}=\operatorname{Im}\left(a_{m*}W_{r}\Omega^{n}_{Y^{\langle m\rangle},\log}\to a_{m+1*}W_{r}\Omega^{n}_{Y^{\langle m+1\rangle},\log}\right).$
Then there is an exact sequence
$0\to\Lambda^{n,m-1}_{r}\to a_{m*}W_{r}\Omega^{n}_{Y^{\langle m\rangle},\log}\to\Lambda^{n,m}_{r}\to 0$
by the definition. Since $Y^{\langle m\rangle}$ is a regular scheme of positive characteristic, we have isomorphisms
$\operatorname{H}^{t}_{y}(Y_{\mathrm{\acute{e}t}},a_{m*}W_{r}\Omega^{n}_{X^{\langle m\rangle},\log})\simeq\operatorname{H}^{t}_{y}((Y^{\langle m\rangle})_{\mathrm{\acute{e}t}},W_{r}\Omega^{n}_{X^{\langle m\rangle},\log})\simeq 0$
for $y\in Y^{(s)}$ and t<sm+1t<s-m+1 by [25, Lemma 3.7] and [29, p.584, Corollary 3.4]. Hence we have isomorphisms
$\operatorname{H}^{t}_{y}(Y_{\mathrm{\acute{e}t}},\lambda^{n}_{r})\simeq\operatorname{H}^{t-1}_{y}(Y_{\mathrm{\acute{e}t}},\Lambda^{n,1}_{r})\simeq\cdots\simeq\operatorname{H}^{t-N+n}_{y}(Y_{\mathrm{\acute{e}t}},\Lambda^{n,N-n}_{r})$
$\simeq$ $\operatorname{H}^{t-N+n}_{y}(Y_{\mathrm{\acute{e}t}},a_{N+1-n*}W_{r}\Omega_{X^{\langle N+1-n\rangle},\log}^{n})\simeq 0$
for $y\in Y^{(s)}$ and $t<s$. This completes the proof. ∎
Theorem 2.26.
Let $A$ be the henselization of the local ring $\mathcal{O}_{Y,y}$ of a normal crossing variety $Y$ over the spectrum of a field of positive characteristic $p>0$ at a point $y\in Y$. Then the sequence
$0\to\operatorname{H}^{s}_{\mathrm{\acute{e}t}}(A,\lambda^{n}_{r})\to\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{s}_{x}(A_{\mathrm{\acute{e}t}},\lambda_{r}^{n})\to\bigoplus_{x\in\operatorname{Spec}(A)^{(1)}}\operatorname{H}^{s+1}_{x}(A_{\mathrm{\acute{e}t}},\lambda_{r}^{n})\to\cdots$ (49)
is exact for any integers $n\geq 0$, $s$ and $r>0$.
Proof.
By [4, Part 1, (1.2)], we have a spectral sequence
$E_{1}^{u,v}=\bigoplus_{x\in\operatorname{Spec}(A)^{(u)}}\operatorname{H}^{u+v}_{x}(A_{\mathrm{\acute{e}t}},\lambda^{n}_{r})\Rightarrow E^{u+v}=\operatorname{H}^{u+v}_{\mathrm{\acute{e}t}}(A,\lambda^{n}_{r}).$
If $s<0$, the sequence (49) is exact by Proposition 2.25. If $s\geq 1$, the sequence (49) is exact by Remark 2.12 and Theorem 2.22. So it suffices to show that the sequence (49) is exact for $s=0$, that is,
$E_{2}^{0,0}=E^{0}$ and $E^{u,0}_{2}=0~~(u>0).$
By [1, Exposé X, Théorème 5.1], we have
$E^{u,0}_{\infty}=E^{u}=0$ (50)
for $u\geq 2$. By Proposition 2.25, we have
$E^{u,v}_{r}=0$ (51)
for $v<0$ and $r\geq 2$. If $u>2$, we have
$E^{u-r,r-1}_{r}=0$
for $r\geq 2$ by Remark 2.12 and Theorem 2.22. So we have
$E^{u,0}_{2}=0$
for $u>2$. By (51), the sequence
$0\to E^{1,0}_{2}\to E^{1}\to E^{0,1}_{2}\to E^{2,0}_{2}$
is exact. By Theorem 2.22, we have an isomorphism
$E^{1}\xrightarrow{\simeq}E^{0,1}_{2}$
and so we have isomorphisms
$E^{1,0}_{2}\simeq 0$ and $\operatorname{Im}(E^{0,1}_{2}\to E^{2,0}_{2})\simeq 0.$
Hence we have isomorphisms
$E^{0}\xrightarrow{\simeq}E^{0,0}_{2}$ and $E^{u,0}_{2}\simeq 0~~(u=1,2)$
by (50) and (51). This completes the proof. ∎
2.3 Mixed characteristic case
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $K$ the quotient field of $B$. Let $\mathfrak{X}$ be a semistable family over $\operatorname{Spec}(B)$, that is, a regular scheme of pure dimension which is flat of finite type over $\operatorname{Spec}(B)$, the generic fiber $\mathfrak{X}\otimes_{B}K$ is smooth over $\operatorname{Spec}(K)$ and the special fiber $Y$ of $\mathfrak{X}$ is a reduced divisor with normal crossings on $\mathfrak{X}$.
As an application of Theorem 2.26, we prove the relative version of the Gersten-type conjecture for the $p$-adic étale $\mathfrak{T}_{r}(n)$ (cf. Definition 1.4) over the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of a semistable family $\mathfrak{X}$ over the spectrum of a discrete valuation ring of mixed characteristic $(0,p)$ in the case where $B$ contains $p^{r}$-th roots of unity (cf. Theorem 2.28).
Proposition 2.27.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let $\mathfrak{X}$ be a semistable family over $\operatorname{Spec}(B)$ and $R$ the local ring of $\mathfrak{X}$ at a point $x$ of the closed fiber $Z$ of $\mathfrak{X}$. Let $j:U\to\mathfrak{X}$ be the inclusion of the generic fiber of $\mathfrak{X}$. Then the homomorphism
$\operatorname{H}^{q}_{x}(X_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})\xrightarrow{\simeq}\operatorname{H}^{q}_{x}(X_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))$
is an isomorphism for $x\in X^{(s)}\cap Z$ and $q\geq s+n+2$. Moreover, the homomorphism
$\operatorname{H}^{q}_{R/(\pi)}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})\xrightarrow{\simeq}\operatorname{H}^{q}_{R/(\pi)}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))$
is an isomorphism for $q\geq n+3$.
Proof.
Let $i:Z\to\mathfrak{X}$ be the inclusion of the closed fiber of $\mathfrak{X}$. Since we have a spectral sequence
$E_{2}^{s,t}=\operatorname{H}_{\mathrm{\acute{e}t}}^{s}(R/(\pi),\mathcal{H}^{t}(i^{*}\mathfrak{T}_{r}(n)))\Rightarrow E^{s+t}=\operatorname{H}^{s+t}_{\mathrm{\acute{e}t}}(R/(\pi),i^{*}\mathfrak{T}_{r}(n)),$
we have an isomorphism
$\operatorname{H}_{\mathrm{\acute{e}t}}^{q}(R/(\pi),i^{*}\mathfrak{T}_{r}(n))\simeq 0$ (52)
for $q\geq n+2$ by [1, Exposé X, Théorème 5.1]. By the similar argument as above, we have an isomorphism
$\operatorname{H}^{q}_{x}\left((R/(\pi))_{\mathrm{\acute{e}t}},i^{*}\mathfrak{T}_{r}(n)\right)\simeq 0$ (53)
for $x\in(\operatorname{Spec}(R))^{(s)}\cap Z$ and $q\geq n+s+1$ by [24, Proposition 2.6]. Since we have a distinguished triangle
$\cdots\to j_{!}\mu_{p^{r}}^{\otimes n}\to\mathfrak{T}_{r}(n)\to i_{*}i^{*}\mathfrak{T}_{r}(n)\to\cdots,$
the statement follows from (52) and (53). ∎
Theorem 2.28.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let $\mathfrak{X}$ be a semistable family over $\operatorname{Spec}(B)$ and $R$ the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of $\mathfrak{X}$ at a point $x$ of the closed fiber of $\mathfrak{X}$. Put $Z=\operatorname{Spec}(R/(\pi))$. Suppose that $B$ contains $p^{r}$-th roots of unity. Then the sequence
$0\to\operatorname{H}_{Z}^{q}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\to\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}_{z}^{q}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\to\displaystyle\bigoplus_{z\in Z^{(1)}}\operatorname{H}_{z}^{q+1}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\to\cdots$
is exact for $q\geq n+2$.
Proof.
Let $i:Z\to\operatorname{Spec}(R)$ be the inclusion of the closed fiber of $\operatorname{Spec}(R)$ and $j:U\to\operatorname{Spec}(R)$ the inclusion of the generic fiber of $\operatorname{Spec}(R)$. Since
$\operatorname{H}^{s}_{z}(R_{\mathrm{\acute{e}t}},i_{*}Ri^{!}\mathfrak{T}_{r}(n))=\begin{cases}\operatorname{H}^{s}_{z}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))&(z\in Z)\\ \\ 0&(z\in U),\end{cases}$
we have a spectral sequence
$E^{s,q}_{1}(n)=\displaystyle\bigoplus_{z\in\operatorname{Spec}(R)^{(s)}\cap Z}\operatorname{H}^{s+q}_{z}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\Rightarrow E^{s+q}(n)=\operatorname{H}_{Z}^{s+q}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))$ (54)
by [24, Proposition 3.8]. Moreover, we have a commutative diagram
${\operatorname{H}_{\mathrm{\acute{e}t}}^{n+1}(R,i_{*}i^{*}\mathfrak{T}_{r}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{Z}^{n+2}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{Z}^{n+2}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}_{z}^{n+1}(R_{\mathrm{\acute{e}t}},i_{*}i^{*}\mathfrak{T}_{r}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}_{z}^{n+2}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}_{z}^{n+2}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))}$
where the sequences are exact by [25, Theorem 1.3] and (52). Since we have an isomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(R,\mathfrak{T}_{r}(n))\simeq\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(R,i_{*}i^{*}\mathfrak{T}_{r}(n))$
by [7, p.777, The proof of Proposition 2.2.b)], we have an isomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(R,i_{*}i^{*}\mathfrak{T}_{r}(n))\simeq\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(Z,\lambda_{r}^{n}).$ (56)
by [24, Theorem 1.2] and [24, Theorem 1.4]. Moreover, we have an isomorphism
$\operatorname{H}^{n+1}_{z}(R_{\mathrm{\acute{e}t}},i_{*}i^{*}\mathfrak{T}_{r}(n))\simeq\operatorname{H}^{1}_{z}(Z_{\mathrm{\acute{e}t}},\lambda_{r}^{n})$
for $z\in Z^{(0)}$ by [25, Theorem 3.5] and [25, Lemma 3.7]. So the left map in the diagram (55) corresponds to the canonical map
$\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(Z,\lambda^{n}_{r})\to\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}^{1}_{z}(Z,\lambda^{n}_{r})$
and the left map in the diagram (55) is injective by Theorem 2.26. Moreover, the right map in the diagram (55) is injective by [24, Theorem 3.10]. Hence the middle map in the diagram (55) is injective. Since $B$ contains $\mu_{p^{r}}$ by the assumption, we have isomorphisms
$\operatorname{H}_{Z}^{v}(R,\mathfrak{T}_{r}(n))\simeq\operatorname{H}_{Z}^{v}(R,j_{!}\mu_{p^{r}}^{\otimes(v-3)})\xrightarrow{\simeq}\operatorname{H}_{Z}^{v}(R,\mathfrak{T}_{r}(v-3))$ (57)
for $v\geq n+3$ by Proposition 2.27. Moreover, we have isomorphisms
$\operatorname{H}_{z}^{v}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\simeq\operatorname{H}_{z}^{v}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes(v-s-3)})\xrightarrow{\simeq}\operatorname{H}_{z}^{v}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(v-s-3))$ (58)
for $z\in\operatorname{Spec}(R)^{(s+1)}\cap Z$ and $v\geq n+s+3$ by Proposition 2.27 and the assumption. So the homomorphism
$E^{v}(n)\to E_{\infty}^{1,v-1}(n)$ (59)
is injective for $v\geq n+2$. Hence we have an isomorphism
$E^{s,q}_{\infty}(n)=0$ (60)
for $s>1$ and $s+q\geq n+2$ by the spectral sequence (54). Then we prove that
$E^{s,q}_{2}(n)=\begin{cases}E^{q+1}(n)&(s=1)\\ \\ 0&(s\geq 2)\end{cases}$ (61)
for any integers $n\geq 0$ and $q\geq n+1$ by induction on $s$.
In the case where $s=1$, we have
$E_{2}^{1,n+1}(n)=E^{n+2}(n)$ (62)
by [24, Theorem 3.10]. Consider a commutative diagram
${\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}_{z}^{n+1}(R_{\mathrm{\acute{e}t}},i_{*}i^{*}\mathfrak{T}_{r}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}_{z}^{n+2}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}_{z}^{n+2}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{z\in Z^{(1)}}\operatorname{H}_{z}^{n+2}(R_{\mathrm{\acute{e}t}},i_{*}i^{*}\mathfrak{T}_{r}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{z\in Z^{(1)}}\operatorname{H}_{z}^{n+3}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\displaystyle\bigoplus_{z\in Z^{(1)}}\operatorname{H}_{z}^{n+3}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))}$
where the sequences are exact by [25, Theorem 1.3]. Then the kernel of the left map is $\operatorname{H}_{\mathrm{\acute{e}t}}^{n+1}(R,i_{*}i^{*}\mathfrak{T}_{r}(n))$ by Theorem 2.22, the isomorphism (56) and [25, Theorem 1.2]. Moreover, the kernel of the right map is $\operatorname{H}_{Z}^{n+2}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{r}(n))$ by the isomorphism (62). By the upper sequence in the diagram (55), the sequence
$0\to\operatorname{H}^{n+2}_{Z}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})\to\displaystyle\bigoplus_{z\in Z^{(0)}}\operatorname{H}^{n+2}_{z}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})\to\displaystyle\bigoplus_{z\in Z^{(1)}}\operatorname{H}^{n+3}_{z}(R_{\mathrm{\acute{e}t}},j_{!}\mu_{p^{r}}^{\otimes n})$
is exact and so we have
$E^{1,q-1}_{2}(n)=E^{q}(n)$ (63)
for $q\geq n+3$ by (57) and (58). Hence we have the equation (61) for any integer $n\geq 0$ and $s=1$. Moreover, we have
$E^{1,q-1}_{\infty}(n)=E^{1,q-1}_{2}(n)$
for $q\geq n+3$ by the injectivity of the homomorphism (59) and the isomorphism (63). So we have
$\operatorname{Im}\left(d^{1,u+n-1}_{u-1}:E^{1,u+n-1}_{u-1}(n)\to E^{u,n+1}_{u-1}(n)\right)=0$ (64)
for $u\geq 3$.
Suppose that the equation (61) holds for any integers $n\geq 0$ and $s\leq t$. By the assumption and (64), we have
$\operatorname{Im}\left(E^{t+1-r,n+r}_{r}(n)\to E^{t+1,n+1}_{r}(n)\right)=0$
for $r\geq 2$. Moreover, we have
$E^{t+r+1,n+2-r}_{r}(n)=0$
for $r\geq 2$ by [27, p.540, Theorem 4.4.7] and [24, Theorem 2.7]. So we have
$E_{2}^{t+1,n+1}(n)=E_{\infty}^{t+1,n+1}(n)=0$ (65)
by (60). By [25, Theorem 1.3] and Proposition 2.27, the sequence
$0\to\displaystyle\bigoplus_{z\in\operatorname{Spec}(R)^{(u)}\cap Z}\operatorname{H}_{z}^{u}(Z_{\mathrm{\acute{e}t}},\lambda_{r}^{n+1})\to E_{1}^{u,n+2}(n)\to E_{1}^{u,(n+1)+1}(n+1)\to 0$
is exact for any integer $u\geq 1$. So we have
$E^{t+1,n+2}_{2}(n)=0$
by Theorem 2.26 and (65). Hence the equation (61) holds for any integer $n\geq 0$ and $s=t+1$. This completes the proof. ∎
3 A relation between the Gersten-type conjecture and motivic cohomology groups
Let $\mathbb{Z}(n)$ be Bloch’s cycle complex for Zariski topology and $\mathbb{Z}/m(n):=\mathbb{Z}(n)\otimes\mathbb{Z}/m\mathbb{Z}$ for any positive integer $m$. As an application of the Beilinson-Lichtenbaum conjecture (cf. [7, p.774, Theorem 1.2.2] and [30]), we obtain the following:
Proposition 3.1.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $R$ a local ring (resp. a henselian local ring) of a regular scheme of finite type over $\operatorname{Spec}(B)$. Let $U$ be the generic fiber of $\operatorname{Spec}(R)$ and $n$ a non-negative integer. Suppose that $B$ contains $p$-th roots of unity.
Then the sequence
$0\to$ $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(U,\mu_{p}^{\otimes n})\to\displaystyle\bigoplus_{x\in U^{(0)}}\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{p}^{\otimes n})\to\displaystyle\bigoplus_{x\in U^{(1)}}\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{p}^{\otimes(n-1)})$
$\to$ $\cdots$ (66)
is exact if and only if we have isomorphisms
$\operatorname{H}^{n+1}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n))\simeq 0$
and
$\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n))\simeq\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n+1))$
for any integer $s\geq 1$. Here $\mu_{p}$ is the sheaf of $p$-th roots of unity.
Proof.
The sequence (3.1) is exact if and only if
$\operatorname{H}^{s}_{\mathrm{Zar}}(U,R^{n+1}\epsilon_{*}\mu_{p}^{\otimes n})=\begin{cases}\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(U,\mu_{p}^{\otimes n})&(\textrm{if}~~s=0)\\ \\ 0&(\textrm{if}~~s\geq 1)\end{cases}$
(cf. The proof of [25, Lemma 5.3 (i)]). Here $\epsilon:U_{\mathrm{\acute{e}t}}\to U_{\mathrm{Zar}}$ is the canonical map of sites and $\epsilon_{*}$ is the forgetful functor. Moreover, we have a distinguished triangle
$\cdots\to\tau_{\leq n}R\epsilon_{*}\mu_{p}^{\otimes n}\to\tau_{\leq(n+1)}R\epsilon_{*}\mu_{p}^{\otimes n}\to R^{n+1}\epsilon_{*}\mu_{p}^{\otimes n}[-(n+1)]\to\cdots$
and a quasi-isomorphism
$\mathbb{Z}/p(N)_{\mathrm{Zar}}^{U}\simeq\tau_{\leq N}R\epsilon_{*}\mu_{p}^{\otimes N}$
for $N=n,n+1$ by [7, p.774, Theorem 1.2] and [30]. Since $B$ contains $p$-th roots of unity, we have an exact sequence
$\cdots\to$ $\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n))\to\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n+1))\to\operatorname{H}^{s}_{\mathrm{Zar}}(U,R^{n+1}\epsilon_{*}\mu_{p}^{\otimes n})$
$\to$ $\operatorname{H}^{n+2+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n))\to\cdots$
for any integer $s\geq 0$ by the above. So the statement follows. ∎
Remark 3.2.
Let $R$ be a local ring of a semistable family over the spectrum of a discrete valuation ring and $U$ the generic fiber of $\operatorname{Spec}(R)$. Let $m$ be a positive integer. Then we have an isomorphism
$\operatorname{H}^{q}_{\mathrm{Zar}}(R,\mathbb{Z}/m(n))\xrightarrow{\simeq}\operatorname{H}^{q}_{\mathrm{Zar}}(U,\mathbb{Z}/m(n))$
for $q\geq n+2$ by the localization theorem ([7, p.779, Theorem 3.2]) and [24, Proposition 2.1]. So we have $\operatorname{H}^{q}_{\mathrm{Zar}}(R,\mathbb{Z}/m(n))=0$ for $q\geq n+2$ if and only if $\operatorname{H}^{q}_{\mathrm{Zar}}(U,\mathbb{Z}/m(n))=0$ for $q\geq n+2$.
The following is an affirmative answer to [27, p.524, Conjecture 1.4.1] in a special case.
Corollary 3.3.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$, $k$ the residue field of $B$ and $k:k^{p}]\leq p^{t}$. Let $\mathfrak{X}$ be a semistable family over $\operatorname{Spec}(B)$ and $n=\operatorname{dim}(\mathfrak{X})+t$. Suppose that $B$ contains $p$-th roots of unity. Then we have a quasi-isomorphism
$\mathbb{Z}/p^{r}(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\simeq\mathfrak{T}_{r}(n)$
for any integer $r\geq 1$.
Proof.
Let $R$ be the strict Henselization $\mathcal{O}_{\mathfrak{X},x}^{sh}$ of $\mathfrak{X}$ at a point $x$ of the closed fiber of $\mathfrak{X}$. By [28, p.209, Remark 7.2], we have a quasi-isomorphism
$\tau_{\leq n}\mathbb{Z}/p^{r}(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\simeq\mathfrak{T}_{r}(n)$
for any positive integer $r$. Moreover, the sequence
$\cdots\to\operatorname{H}^{u}_{\mathrm{Zar}}(R,\mathbb{Z}/p^{r}(n))\to\operatorname{H}^{u}_{\mathrm{Zar}}(R,\mathbb{Z}/p^{r+r^{\prime}}(n))\to\operatorname{H}^{u}_{\mathrm{Zar}}(R,\mathbb{Z}/p^{r^{\prime}}(n))\to\cdots$
is exact for positive integers $r$ and $r^{\prime}$. So it suffices to show that we have an isomorphism
$\operatorname{H}^{u}_{\mathrm{Zar}}(R,\mathbb{Z}/p(n))\simeq 0$
for $u\geq n+1$. Let $U$ be the generic fiber of $\operatorname{Spec}(R)$. Then it suffices to show that we have an isomorphism
$\operatorname{H}^{u}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n))\simeq 0$
for $u\geq n+2$ by [24, Proposition 1.5], [24, Proposition 2.1] and the localization theorem ([7, p.779, Theorem 3.2]).
Let $x^{\prime}$ be an inverse image of a point $x$ under the morphism $\mathfrak{X}\times B^{sh}\to\mathfrak{X}.$ Then we have an isomorphism
$\mathcal{O}^{sh}_{\mathfrak{X},x}\simeq\mathcal{O}^{sh}_{\mathfrak{X}\times_{B}B^{sh},x^{\prime}}$
and an equation
$k_{s}:k^{p}_{s}]=[k:k^{p}]\leq p^{t}$
where $k_{s}$ is a separable closed field of $k$. So we are able to assume that $k$ is a separable closed field.
We consider the cohomological $p$-dimension $\operatorname{cd}_{p}(k(\mathcal{O}_{\mathfrak{X},x}^{sh}))$ of $k(\mathcal{O}_{\mathfrak{X},x}^{sh})$. We have an equation
$\operatorname{cd}_{p}(k(B))=1+t$
by [6, p.48, Théorème 1.2]. Since $k(\mathcal{O}_{\mathfrak{X},x}^{sh})$ is an algebraic extension of $k(\mathcal{O}_{\mathfrak{X},x})$, we have
$\operatorname{tr.deg}_{k(B)}k(\mathcal{O}^{sh}_{\mathfrak{X},x})=\operatorname{dim}(\mathfrak{X})-1$
by [16, pp.118–119, Theorem 15.5] and [16, p.119, Theorem 15.6]. Here $\operatorname{tr.deg}_{k(B)}k(\mathcal{O}^{sh}_{\mathfrak{X},x})$ is the transcendence degree of $k(\mathcal{O}^{sh}_{\mathfrak{X},x})$ over $k(B)$. So we have
$\operatorname{cd}_{p}(k(\mathcal{O}^{sh}_{\mathfrak{X},x}))\leq\operatorname{dim}(\mathfrak{X})+t=n$
by [6, p.57, Corollaire 2.3.2].
Since the homomorphism
$\operatorname{H}^{u}_{\mathrm{\acute{e}t}}(U,\mu_{p}^{\otimes n})\to\operatorname{H}^{u}_{\mathrm{\acute{e}t}}(k(R),\mu_{p}^{\otimes n})$
is injective for $u\geq n+1$ by [24, Proposition 4.2] and [24, Theorem 3.10], we have an isomorphism
$\operatorname{H}^{u}_{\mathrm{\acute{e}t}}(U,\mu_{p}^{\otimes n})\simeq 0$
for $u\geq n+1$. Here $\mu_{p}$ is the sheaf of $p$-th roots of unity. Moreover, we have
$\operatorname{cd}_{p}(\kappa(x))\leq n-\operatorname{dim}(\mathcal{O}_{\mathfrak{X},x})$
for $x\in U$ by [6, p.57, Corollaire 2.3.2] and [16, pp.118–119, Theorem 15.5]. Hence we have isomorphisms
$\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n))\simeq\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n+1))\simeq\cdots\simeq\operatorname{H}^{n+1+s}_{\mathrm{Zar}}(U,\mathbb{Z}/p(n+s))\simeq 0$
for any integer $s\geq 1$ by Proposition 3.1. This completes the proof. ∎
Corollary 3.4.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $R$ the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of a semistable family $\mathfrak{X}$ over $\operatorname{Spec}(B)$ at a point $x\in\mathfrak{X}$. Suppose that $\operatorname{dim}(R)=2$ and $B$ contains $p$-th roots of unity. Then the sequence
$0\to$ $\operatorname{H}^{s}_{\mathrm{\acute{e}t}}(R,\mathfrak{T}_{1}(n))\to\bigoplus_{x\in\operatorname{Spec}(R)^{(0)}}\operatorname{H}^{s}_{x}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to$
$\bigoplus_{x\in\operatorname{Spec}(R)^{(1)}}\operatorname{H}^{s+1}_{x}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to\bigoplus_{x\in\operatorname{Spec}(R)^{(2)}}\operatorname{H}^{s+2}_{x}(R_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to 0$
is exact for any integer $s\geq 0$.
Proof.
Let $U$ be the generic fiber of $\operatorname{Spec}(R)$ and $N$ an integer larger than or equal to $n$. Then we have an isomorphism
$\operatorname{H}^{N+1}_{\mathrm{Zar}}(U,\mathbb{Z}/p(N))\simeq 0$
by [24, Proposition 4.5], [24, Proposition 2.1] and the localization theorem ([7, p.779, Theorem 3.2]). Moreover, we have an isomorphism
$\operatorname{H}^{N+t}_{\mathrm{Zar}}(U,\mathbb{Z}/p(N))\simeq 0$
for any integer $t\geq 2$ by the proof of [24, Corollary 4.6].
In the case where $s\leq n$, the statement follows from [27, p.540, Theorem 4.4.7], [24, Theorem 1.1] and Proposition 3.1. In the case where $s\geq n+1$, the statement follows from Theorem 2.28 and Proposition 3.1. This completes the proof. ∎
Corollary 3.5.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $R$ the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of a semistable family $\mathfrak{X}$ over $\operatorname{Spec}(B)$ at a point $x\in\mathfrak{X}$. Suppose that $B$ contains $p$-th roots of unity. Then the sequence
$0$ $\to\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(R,\mathbb{Z}/p\mathbb{Z})\to\bigoplus_{x\in\operatorname{Spec}(R)^{(0)}}\operatorname{H}^{1}_{x}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/p\mathbb{Z})\to\bigoplus_{x\in\operatorname{Spec}(R)^{(1)}}\operatorname{H}^{2}_{x}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/p\mathbb{Z})$
$\to\cdots$
is exact.
Proof.
By the definition, we have a quasi-isomorphism
$\mathbb{Z}(0)_{\mathrm{Zar}}^{\operatorname{Spec}(R)}\simeq\mathbb{Z}.$
Moreover, we have a quasi-isomorphism
$\mathbb{Z}(1)^{\operatorname{Spec}(R)}_{\mathrm{Zar}}\simeq\mathbb{G}_{m}[-1]$
by [15, Lemma 11.2]. So we have an isomorphism
$\operatorname{H}_{\mathrm{Zar}}^{N+t}(U,\mathbb{Z}/p(N))\simeq 0$
for $N=0,1$ and an integer $t\geq 1$ by [24, Proposition 2.1] and the localization theorem ([7, p.779, Theorem 3.2]). Here $U$ is the generic fiber of $\operatorname{Spec}(R)$. Hence the statement follows from Theorem 2.28 and Proposition 3.1. ∎
Remark 3.6.
If $\mathfrak{X}$ is a smooth scheme over the spectrum of a discrete valuation ring, then we have $\mathcal{H}^{s}(\mathbb{Z}(n)^{\mathfrak{X}})=0$ for $s>n$ by [7, p.786, Corollary 4.4]. Hence [25, Theorem 1.4] follows from Theorem 2.28 and Proposition 3.1.
4 Computations of motivic cohomology groups on global cases
4.1 Equi-characteristic cases
We compute motivic cohomology groups in equi-characteristic cases and prepare to compute motivic cohomology groups in mixed characteristic cases.
Proposition 4.1.
Let $k$ be a field and
$D=k[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a})$
for $0\leq a\leq N$. Then we have
$\operatorname{H}^{q}_{\mathrm{Zar}}(D,\mathbb{Z}(n))=0$
for any integer $n\geq 0$ and $q\geq n+1$.
Proof.
By [7, p.781, Corollary 3.5], it suffices to show the statement in the case where $a=N$. Then $\operatorname{Spec}(D)$ is isomorphic to
$\operatorname{Spec}(k[T_{0},\cdots,T_{a}]/(T_{0}\cdots T_{a-1}))\cup\operatorname{Spec}(k[T_{0},\cdots,T_{a}]/(T_{a})).$
Let
$Z=\operatorname{Spec}(k[T_{0},\cdots,T_{a}]/(T_{0}\cdots T_{a-1},T_{a}))\simeq\operatorname{Spec}(k[T_{0},\cdots,T_{a-1}]/(T_{0}\cdots T_{a-1}))$
and
$U=\operatorname{Spec}(D)\setminus Z.$
Then we have
$U\simeq\operatorname{Spec}\left((k[T_{0},\cdots,T_{a}]/(T_{0}\cdots T_{a-1}))_{T_{a}}\right)\oplus\operatorname{Spec}\left((k[T_{0},\cdots,T_{a-1}])_{T_{0}\cdots T_{a-1}}\right).$
Moreover, the homomorphism
$\operatorname{H}^{n}_{\mathrm{Zar}}\left((k[T_{0},\cdots,T_{a-1}])_{T_{0}\cdots T_{a-1}},\mathbb{Z}(n)\right)\to\operatorname{H}^{n-1}_{\mathrm{\acute{e}t}}(Z,\mathbb{Z}(n-1))$
is surjective by the localization theorem ([7, p.779, Theorem 3.2]) and [7, p.781, Corollary 3.5]. Since we have a commutative diagram
${\operatorname{H}^{n}_{\mathrm{Zar}}\left((k[T_{0},\cdots,T_{a-1}])_{T_{0}\cdots T_{a-1}},\mathbb{Z}(n)\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n-1}_{\mathrm{Zar}}(Z,\mathbb{Z}(n-1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n}_{\mathrm{Zar}}(U,\mathbb{Z}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n-1}_{\mathrm{Zar}}(Z,\mathbb{Z}(n-1)),}$
the homomorphism
$\operatorname{H}^{n}_{\mathrm{Zar}}(U,\mathbb{Z}(n))\to\operatorname{H}^{n-1}_{\mathrm{Zar}}(Z,\mathbb{Z}(n-1))$
is surjective. So there exists an exact sequence
$0\to$ $\operatorname{H}^{n+1}_{\mathrm{Zar}}(D,\mathbb{Z}(n))\to\operatorname{H}^{n+1}_{\mathrm{Zar}}(U,\mathbb{Z}(n))\to\operatorname{H}^{n}_{\mathrm{Zar}}(Z,\mathbb{Z}(n-1))$
$\to$ $\operatorname{H}^{n+2}_{\mathrm{Zar}}(D,\mathbb{Z}(n))\to\operatorname{H}^{n+2}_{\mathrm{Zar}}(U,\mathbb{Z}(n))\to\cdots$
by the localization theorem ([7, p.779, Theorem 3.2]). Moreover, we have exact sequences
$\operatorname{H}^{q}_{\mathrm{Zar}}\left(k[T_{0},\cdots,T_{a}]/(T_{0}\cdots T_{a-1}),\mathbb{Z}(n)\right)\to\operatorname{H}^{q}_{\mathrm{Zar}}\left((k[T_{0},\cdots,T_{a}]/(T_{0}\cdots T_{a-1}))_{T_{a}},\mathbb{Z}(n)\right)$
$\to\operatorname{H}^{q-1}_{\mathrm{Zar}}\left(k[T_{0},\cdots,T_{a-1}]/(T_{0}\cdots T_{a-1}),\mathbb{Z}(n-1)\right)$
and
$\operatorname{H}^{q}_{\mathrm{Zar}}\left(k[T_{0},\cdots,T_{a-1}],\mathbb{Z}(n)\right)\to\operatorname{H}^{q}_{\mathrm{Zar}}\left((k[T_{0},\cdots,T_{a-1}])_{T_{0}\cdots T_{a-1}},\mathbb{Z}(n)\right)$
$\to\operatorname{H}^{q-1}_{\mathrm{Zar}}\left(k[T_{0},\cdots,T_{a-1}]/(T_{0}\cdots T_{a-1}),\mathbb{Z}(n-1)\right)$
for any integer $q$. Hence, the statement follows by induction on $a$. ∎
We use the following results later.
Proposition 4.2.
Let $Y$ be a scheme which is essentially of finite type over the spectrum of a field $k$. Then we have a quasi-isomorphism
$\tau_{\leq n+1}\left(\mathbb{Z}(n)^{Y}_{\mathrm{Zar}}\right)\xrightarrow{\simeq}\tau_{\leq n+1}\left(R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{Y}\right)$ (67)
for any integer $n$. Here $\epsilon:Y_{\mathrm{\acute{e}t}}\to Y_{\mathrm{Zar}}$ is the canonical map of sites.
Proof.
Let $A$ be a local ring $\mathcal{O}_{Y,y}$ of $Y$ at a point $y\in Y$. It suffices to prove the quasi-isomorphism (67) for $\operatorname{Spec}(A)$. We prove the statement by induction on $\operatorname{dim}(A)$.
Assume that $\operatorname{dim}(A)=0,$ that is, $A$ is a field. Then we have the quasi-isomorphism (67) by [7, p.774, Theorem 1.2.2] and [30].
Assume that the quasi-isomorphism (67) holds for $\operatorname{Spec}(A^{\prime})$ in the case where $A^{\prime}$ is a local ring of a scheme of finite type over the spectrum of a field at a point and $\operatorname{dim}(A^{\prime})<\operatorname{dim}(A).$
Let $j:U\to\operatorname{Spec}(A)$ be an open immersion of $\operatorname{Spec}(A)$ such that $U$ is a regular scheme. Let $\iota:Z\to\operatorname{Spec}(A)$ be a closed subscheme of codimension $c\geq 1$ with open complement $j:U\to\operatorname{Spec}(A)$. Since $A$ is an excellent ring, we are able to choose such a non-empty open subset $U$ of $\operatorname{Spec}(A)$.
Let $A^{\prime\prime}$ be a regular local ring which is essentially finite type over a field $k$. By Quillen’s method (cf. [20, §7, The proof of Theorem 5.11]),
$A^{\prime\prime}=\lim_{\begin{subarray}{c}\to\\ s\end{subarray}}A^{\prime\prime}_{s}$
where $A^{\prime\prime}_{s}$ is a local ring of a smooth algebra over the prime field and the maps $A^{\prime\prime}_{s}\to A^{\prime\prime}_{t}$ are flat. So we have a quasi-isomorphism
$\mathbb{Z}(n)^{U}_{\mathrm{Zar}}\xrightarrow{\simeq}\tau_{\leq n+1}R\epsilon_{*}\mathbb{Z}(n)^{U}_{\mathrm{\acute{e}t}}$ (68)
by [7, p.774, Theorem 1.2.2] and [30].
In the case where $U=\operatorname{Spec}(A)$, then the quasi-isomorphism (67) holds. So it suffices to prove the quasi-isomorphism (67) in the case where $U\neq\operatorname{Spec}(A)$.
Consider a map of distinguished triangles in the derived category of étale sheaves on $\operatorname{Spec}(A)$
${\iota_{*}\mathbb{Z}(n-c)[-2c]_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\epsilon^{*}Rj_{*}\mathbb{Z}(n)_{\mathrm{Zar}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\iota_{*}R\iota^{!}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Rj_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}.}$ (69)
Then the right map agrees with the composite
$\epsilon^{*}Rj_{*}\mathbb{Z}(n)_{\mathrm{Zar}}\xrightarrow[\simeq]{(*)}\epsilon^{*}Rj_{*}\left(\tau_{\leq n+1}R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$
$\to$ $\epsilon^{*}Rj_{*}\left(R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)\xleftarrow[\simeq]{(**)}Rj_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$
where the maps ()(*) and ()(**) are quasi-isomorphisms by the quasi-isomorphism (68) and [7, p.776, Proposition 2.2. a)]. Since the homomorphism
$\epsilon^{*}R^{q}j_{*}\left(\tau_{\leq n+1}R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)\to\epsilon^{*}R^{q}j_{*}\left(R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$
is an isomorphism for $q\leq n+1$ and injective for $q=n+2$, the homomorphism
$\epsilon^{*}R^{q}j_{*}\mathbb{Z}(n)_{\mathrm{Zar}}\to R^{q}j_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$
is an isomorphism for $q\leq n+1$ and injective for $q=n+2$. So we have a quasi-isomorphism
$\tau_{\leq n+2}\left(\iota_{*}\mathbb{Z}(n-c)_{\mathrm{\acute{e}t}}[-2c]\right)\xrightarrow{\simeq}\tau_{\leq n+2}\left(\iota_{*}R\iota^{!}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$ (70)
by the five lemma. Moreover, we consider a map of distinguished triangles in the derived category of Zariski sheaves on $\operatorname{Spec}(A)$
${\iota_{*}\mathbb{Z}(n-c)[-2c]_{\mathrm{Zar}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\mathbb{Z}(n)_{\mathrm{Zar}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${Rj_{*}\mathbb{Z}(n)_{\mathrm{Zar}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R\epsilon_{*}\iota_{*}R\iota^{!}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R\epsilon_{*}Rj_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}.}$
Then the left map is the composition of the map
$\iota_{*}\mathbb{Z}(n-c)[-2c]_{\mathrm{Zar}}\to\iota_{*}R\epsilon_{*}\mathbb{Z}(n-c)[-2c]_{\mathrm{\acute{e}t}}$
and the map
$R\epsilon_{*}\iota_{*}\mathbb{Z}(n-c)[-2c]_{\mathrm{\acute{e}t}}\to R\epsilon_{*}\iota_{*}R\iota^{!}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$
which is obtained by applying $R\epsilon_{*}$ to the left map in the commutative diagram (69). So we have a quasi-isomorphism
$\tau_{\leq n+2}\left(\iota_{*}\mathbb{Z}(n-c)[-2c]_{\mathrm{Zar}}\right)\xrightarrow{\simeq}\tau_{\leq n+2}\left(R\epsilon_{*}\iota_{*}R\iota^{!}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$
by the assumption and the quasi-isomorphism (70). Hence, the quasi-isomorphism (67) holds for $\operatorname{Spec}(R)$ by the five lemma. This completes the proof. ∎
Proposition 4.3.
Let $A$ be a local ring of a normal crossing variety over the spectrum of a field $k$. Let $n\geq 0$, $l>0$ be integers with $(l,\operatorname{char}(k))=1$. Then we have isomorphisms
$\mathcal{H}^{q}(\mathbb{Z}/l(n)_{\mathrm{Zar}}^{\operatorname{Spec}(A)})=0$ (71)
for $q<0$ and
$\operatorname{H}^{0}_{\mathrm{Zar}}(A,\mathbb{Z}/l(n))\xrightarrow{\simeq}\bigoplus_{x\in\operatorname{Spec}(A)^{(0)}}\operatorname{H}^{0}_{\mathrm{Zar}}(\kappa(x),\mathbb{Z}/l(n))\simeq\underbrace{\mu_{l}^{\otimes n}\oplus\cdots\oplus\mu_{l}^{\otimes n}}_{\#\left((\operatorname{Spec}(A))^{(0)}\right)~\textrm{times}}.$ (72)
Moreover, we have an isomorphism
$\mathcal{H}^{q}(\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(A)})=0$ (73)
for $q\geq\#\left((\operatorname{Spec}(A))^{(0)}\right)$.
Proof.
First we prove the isomorphism (71). By [7, p.779, Theorem 3.2.b)], we have isomorphisms
$\operatorname{H}^{q}_{\mathrm{Zar}}(A_{x},\mathbb{Z}/l(n))\simeq\Gamma(\operatorname{Spec}(A_{x}),\mathcal{H}^{q}(\mathbb{Z}/l(n)_{\mathrm{Zar}}^{\operatorname{Spec}(A)}))\simeq\mathcal{H}^{q}(\mathbb{Z}/l(n)^{\operatorname{Spec}(A)}_{\mathrm{Zar}})_{x}$
for $x\in\operatorname{Spec}(A)$ and any integer $q$ where $A_{x}$ is the local ring of $A$ at a point $x\in\operatorname{Spec}(A)$.
Let $Z_{x}$ be an irreducible component of $\operatorname{Spec}(A_{x})$ and $U_{x}=\operatorname{Spec}(A_{x})\setminus Z_{x}$. Then there exists an exact sequence
$\operatorname{H}^{q}_{\mathrm{Zar}}(Z_{x},\mathbb{Z}/l(n))\to\operatorname{H}^{q}_{\mathrm{Zar}}(A_{x},\mathbb{Z}/l(n))\to\operatorname{H}^{q}_{\mathrm{Zar}}(U_{x},\mathbb{Z}/l(n))$
by the localization theorem ([7, p.779, Theorem 3.2]). Since we have isomorphisms
$\operatorname{H}^{q}_{\mathrm{Zar}}(Z_{x},\mathbb{Z}/l(n))\simeq\operatorname{H}^{q}_{\mathrm{\acute{e}t}}(Z_{x},\mu_{l}^{\otimes n})=0$
for $q<0$ by [7, p.774, Theorem 1.2] and [30], it suffices to prove that
$\operatorname{H}^{q}_{\mathrm{Zar}}(U_{x},\mathbb{Z}/l(n))=0$
for $q<0$.
Let $Z^{\prime}_{x}$ be the union of the irreducible components of $\operatorname{Spec}(A_{x})$ excluding $Z_{x}$. Then $Z^{\prime}_{x}$ and $Z^{\prime}_{x}\cap Z_{x}$ are the spectrum of a local ring of a normal crossing variety over $\operatorname{Spec}(k)$ and there exists an exact sequence
$\operatorname{H}^{q}_{\mathrm{Zar}}(Z_{x}^{\prime},\mathbb{Z}/l(n))\to\operatorname{H}^{q}_{\mathrm{Zar}}(U_{x},\mathbb{Z}/l(n))\to\operatorname{H}^{q-1}_{\mathrm{Zar}}(Z_{x}\cap Z_{x}^{\prime},\mathbb{Z}/l(n-1))$
by the localization theorem ([7, p.779, Theorem 3.2]). Since we have
$\#\left((Z^{\prime}_{x})^{(0)}\right)=\#\left((Z^{\prime}_{x}\cap Z_{x})^{(0)}\right)=\#\left((\operatorname{Spec}(A_{x}))^{(0)}\right)-1,$
the isomorphism (71) follows by induction on $\left((\operatorname{Spec}(A_{x}))^{(0)}\right)$.
Next we prove the isomorphism (72). We also prove the isomorphism (72) by induction on $\left((\operatorname{Spec}(A))^{(0)}\right)$.
In the case where $\left((\operatorname{Spec}(A))^{(0)}\right)=1$, the isomorphism (72) follows from [7, p.774, Theorem 1.2.4] and [30].
Assume that the isomorphism (72) holds in the case where $\left((\operatorname{Spec}(A))^{(0)}\right)\leq s$. Suppose that $\left((\operatorname{Spec}(A))^{(0)}\right)=s+1$. Let $\bar{\{y\}}$ be an irreducible component of $\operatorname{Spec}(A)$ and $U_{0}=\operatorname{Spec}(A)\setminus\bar{\{y\}}$. Then there exists an exact sequence
$0\to\operatorname{H}^{0}_{\mathrm{Zar}}(\bar{\{y\}},\mathbb{Z}/l(n))\to\operatorname{H}^{0}_{\mathrm{Zar}}(\operatorname{Spec}(A),\mathbb{Z}/l(n))\to\operatorname{H}^{0}_{\mathrm{Zar}}(U_{0},\mathbb{Z}/l(n))$
by the localization theorem ([7, p.779, Theorem 3.2]) and the isomorphism (71). Let $Z_{1}$ be the union of the irreducible components of $\operatorname{Spec}(A)$ excluding $\bar{\{y\}}$. Since $Z_{1}\cap\bar{\{y\}}$ is the spectrum of a local ring of a normal crossing variety over $\operatorname{Spec}(k)$, we have
$\operatorname{H}^{q}_{\mathrm{Zar}}(Z_{1}\cap\bar{\{y\}},\mathbb{Z}/l(n-1))=0$
for $q<0$ by the isomorphism (71). Moreover, we have
$U_{0}=Z_{1}\setminus\left(Z_{1}\cap\bar{\{y\}}\right)$
and $((Z_{1})^{(0)})=s$. So we have isomorphisms
$\operatorname{H}^{0}_{\mathrm{Zar}}(Z_{1},\mathbb{Z}/l(n))\xrightarrow{\simeq}\operatorname{H}^{0}_{\mathrm{Zar}}(U_{0},\mathbb{Z}/l(n))\xrightarrow{\simeq}\bigoplus_{z\in(Z_{1})^{(0)}}\operatorname{H}^{0}_{\mathrm{Zar}}(\kappa(z),\mathbb{Z}/l(n))$
by the localization theorem ([7, p.779, Theorem 3.2]) and the assumption. Moreover, the composition
$\operatorname{H}^{0}_{\mathrm{Zar}}(\bar{\{z\}},\mathbb{Z}/l(n))\to\operatorname{H}^{0}_{\mathrm{Zar}}(A,\mathbb{Z}/l(n))\to\operatorname{H}^{0}_{\mathrm{Zar}}(U_{0},\mathbb{Z}/l(n))\to\operatorname{H}^{0}_{\mathrm{Zar}}(\kappa(z),\mathbb{Z}/l(n))$
is an isomorphism for $z\in(Z_{1})^{(0)}$. So we have the isomorphism (72) in the case where $((\operatorname{Spec}(A))^{(0)})=s+1$. Hence, the isomorphism (72) holds.
Finally, we prove the isomorphism (73). By [24, Proposition 2.1], it suffices to prove the isomorphism (73) in the case where $((\operatorname{Spec}(A))^{(0)})<n+1$. Let $A_{\bar{x}}$ be the strict henselian of $A$ at a point $x\in\operatorname{Spec}(A)$. If $l^{\prime}$ is a prime number with $(l^{\prime},\operatorname{char}(k))=1$, we have $\operatorname{cd}_{l^{\prime}}(A_{\bar{x}})=0$. So we have isomorphisms
$\operatorname{H}^{q}_{\mathrm{\acute{e}t}}(A_{\bar{x}},\mathbb{Z}/l(n))\simeq\Gamma\left(A_{\bar{x}},\mathcal{H}^{q}(\mathbb{Z}/l(n))_{\mathrm{\acute{e}t}}\right)\simeq\mathcal{H}^{q}\left(\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}\right)_{\bar{x}}$
for $q>0$. By [7, p.774, Theorem 1.2.2] and [30], it suffices to prove that
$\operatorname{H}^{q}_{\mathrm{Zar}}(A_{\bar{x}},\mathbb{Z}/l(n))=0$
for $q\geq\#((\operatorname{Spec}(A_{\bar{x}}))^{(0)})=\#((\operatorname{Spec}(A_{x}))^{(0)}).$ By a similar argument as in the proof of the isomorphism (71), the isomorphism (73) holds as follows.
Let $Z_{\bar{x}}$ be an irreducible component of $\operatorname{Spec}(A_{\bar{x}})$ and $Z_{\bar{x}}^{\prime}$ the union of the irreducible components of $\operatorname{Spec}(A_{\bar{x}})$ excluding $Z_{\bar{x}}$. Then we have
$U_{\bar{x}}:=\operatorname{Spec}(A_{\bar{x}})\setminus Z_{\bar{x}}=Z^{\prime}_{\bar{x}}\setminus Z^{\prime}_{\bar{x}}\cap Z_{\bar{x}}$
and isomorphisms
$\operatorname{H}^{q}_{\mathrm{Zar}}(Z_{\bar{x}},\mathbb{Z}/l(n))\simeq\operatorname{H}^{q}_{\mathrm{\acute{e}t}}(Z_{\bar{x}},\mathbb{Z}/l(n))\simeq\operatorname{H}^{q}_{\mathrm{\acute{e}t}}(Z_{\bar{x}},\mu_{l}^{\otimes n})\simeq 0$
for $q>0$ by [7, p.774, Theorem 1.2.2 and Theorem 1.2.4], [30] and [7, p.786, Corollary 4.4]. So we have an isomorphism
$\operatorname{H}^{q}_{\mathrm{Zar}}(A_{\bar{x}},\mathbb{Z}/l(n))\simeq\operatorname{H}^{q}_{\mathrm{Zar}}(U_{\bar{x}},\mathbb{Z}/l(n))$
for $q>0$. Moreover, we have an exact sequence
$\operatorname{H}^{q}_{\mathrm{Zar}}(Z^{\prime}_{\bar{x}},\mathbb{Z}/l(n))\to\operatorname{H}^{q}_{\mathrm{Zar}}(U_{\bar{x}},\mathbb{Z}/l(n))\to\operatorname{H}^{q-1}_{\mathrm{Zar}}(Z^{\prime}_{\bar{x}}\cap Z_{\bar{x}},\mathbb{Z}/l(n-1)$
for $q>0$ by the localization theorem ([7, p.779, Theorem 3.2]). Since we have
$\#((Z^{\prime}_{\bar{x}})^{(0)})=\#((Z^{\prime}_{\bar{x}}\cap Z_{\bar{x}})^{(0)})=\#((\operatorname{Spec}(A_{\bar{x}}))^{(0)})-1,$
the isomorphism (73) follows by induction on $((\operatorname{Spec}(A_{\bar{x}}))^{(0)})$. This completes the proof. ∎
Remark 4.4.
Let $A_{\bar{x}}$ be the strict henselization of a normal crossing variety $Y$ over the spectrum of a field $k$ at a point $x\in Y$ and $l$ an integer which is prime to $\operatorname{char}(k)$.
If $((\operatorname{Spec}(A_{\bar{x}}))^{(0)})<n+1$ and $s=\#((\operatorname{Spec}(A_{\bar{x}}))^{(0)})-1$, then
$\mathcal{H}^{s}(\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}^{Y})_{\bar{x}}\simeq\operatorname{H}^{s}_{\mathrm{\acute{e}t}}(A_{\bar{x}},\mathbb{Z}/l(n))\neq 0.$
In fact, we have isomorphisms
$\operatorname{H}^{s}_{\mathrm{\acute{e}t}}(A_{\bar{x}},\mathbb{Z}/l(n))\simeq\operatorname{H}^{s-1}_{\mathrm{\acute{e}t}}(Z_{\bar{x}}\cap Z_{\bar{x}}^{\prime},\mathbb{Z}/l(n-1))\simeq\cdots\simeq\mu_{l}^{\otimes(n-s)}$
by the same argument as in the proof of the isomorphism (73).
4.2 Mixed characteristic cases
We compute the motivic cohomology in mixed characteristic.
Proposition 4.5.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let
$C=B[T_{0},T_{1},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi)$
for $0\leq a\leq N$. Then we have
$\operatorname{H}^{q}_{\mathrm{Zar}}(C,\mathbb{Z}(n))=0$ (74)
for $q>n+1$. Moreover, we have
$\operatorname{H}^{q}_{\mathrm{Zar}}(C,\mathbb{Z}/m(n))=0$ (75)
for any integers $m\geq 0$ and $q>n+1$.
Proof.
Since we have an exact sequence
$\cdots\to\operatorname{H}^{s}_{\mathrm{Zar}}(C,\mathbb{Z}(n))\xrightarrow{\times m}\operatorname{H}^{s}_{\mathrm{Zar}}(C,\mathbb{Z}(n))\to\operatorname{H}^{s}_{\mathrm{Zar}}(C,\mathbb{Z}/m(n))\to\cdots$
for any integer $s$, the equation (75) follows from the equation (74). So it suffices to prove the equation (74).
If $a<N$, we have an isomorphism
$\operatorname{H}^{q}_{\mathrm{Zar}}(C,\mathbb{Z}(n))\simeq\operatorname{H}^{q}_{\mathrm{Zar}}(B[T_{0},\cdots,T_{a}]/(T_{0}\cdots T_{a}-\pi),\mathbb{Z}(n))$
by [7, p.781, Corollary 3.5]. So it suffices to show (74) in the case where $a=N$. We prove the statement by induction on $a=N$.
Let $a=N=0$. Then we have an isomorphism $C\simeq B$. So we have the equation (74) by [7, p.786, Corollary 4.4].
Assume that the equation (74) holds in the case where $a=N\leq b$. We have a homomorphism of polynomial rings over $B$
$B[T_{0}^{\prime},\cdots,T_{b}^{\prime},T_{b+1}^{\prime}]\to B[T_{0},\cdots,T_{b},T_{b+1}]$
which sends $T_{b}^{\prime}$ to $T_{b}T_{b+1}$ and sends $T_{r}^{\prime}$ to $T_{r}$ for $r\neq b$. Then this homomorphism induces an isomorphism
$\left(B[T_{0}^{\prime},\cdots,T_{b}^{\prime},T_{b+1}^{\prime}]/(T_{0}^{\prime}\cdots T_{b}^{\prime}-\pi)\right)_{T_{b+1}^{\prime}}$
$\xrightarrow{\simeq}$ $\left(B[T_{0},\cdots,T_{b},T_{b+1}]/(T_{0}\cdots T_{b}\cdot T_{b+1}-\pi)\right)_{T_{b+1}}.$ (76)
Moreover, we have an isomorphism
$\left(B[T_{0}^{\prime},\cdots,T_{b-1}^{\prime},T_{b}^{\prime},T_{b+1}^{\prime}]/(T_{0}^{\prime}\cdots T_{b-1}^{\prime}\cdot T_{b}^{\prime}-\pi)\right)_{T_{b+1}^{\prime}}$
$\simeq$ $\bigl(\left(B[T_{0}^{\prime},\cdots,T_{b-1}^{\prime},T_{b}^{\prime}]/(T_{0}^{\prime}\cdots T_{b-1}^{\prime}\cdot T_{b}^{\prime}-\pi)\right)\otimes_{B}B[T_{b+1}^{\prime}]\bigr)_{T_{b+1}^{\prime}}.$
Since there exists an exact sequence
$\operatorname{H}^{q}_{\mathrm{Zar}}\left((B[T_{0}^{\prime},\cdots,T_{b}^{\prime}]/(T_{0}^{\prime}\cdots T_{b}^{\prime}-\pi))\otimes_{B}B[T_{b+1}^{\prime}],\mathbb{Z}(n)\right)$
$\to$ $\operatorname{H}^{q}_{\mathrm{Zar}}(C_{T_{b+1}},\mathbb{Z}(n))\to\operatorname{H}^{q-1}_{\mathrm{Zar}}(B[T_{0}^{\prime},\cdots,T_{b}^{\prime}]/(T_{0}^{\prime}\cdots T_{b}^{\prime}-\pi),\mathbb{Z}(n-1))$
for $N=a=b+1$ by the localization theorem ([7, p.779, Theorem 3.2]), we have
$\operatorname{H}^{q}_{\mathrm{Zar}}(C_{T_{b+1}},\mathbb{Z}(n))=0$ (77)
for $N=a=b+1$ and $q>n+1$ by [7, p.781, Corollary 3.5] and the assumption. Moreover, we have isomorphisms
$\operatorname{H}_{\mathrm{Zar}}^{q}(C/(T_{b+1}),\mathbb{Z}(n-1))\simeq\operatorname{H}^{q}_{\mathrm{Zar}}((B/(\pi))[T_{0},\cdots,T_{b}],\mathbb{Z}(n-1))\simeq 0$ (78)
for $N=a=b+1$ and $q\geq n$ by [7, p.781, Corollary 3.5]. Hence we have the equation (74) for $q>n+1$ by (77), (78) and the localization theorem ([7, p.779, Theorem 3.2]). This completes the proof. ∎
Proposition 4.6.
Let $\mathfrak{X}$ be a regular scheme which is essentially of finite type over the spectrum of a discrete valuation ring in mixed characteristic $(0,p)$. Then we have a quasi-isomorphism
$\tau_{\leq n+1}\left(\mathbb{Z}(n)_{\mathrm{Zar}}^{\mathfrak{X}}\right)\xrightarrow{\simeq}\tau_{\leq n+1}\left(R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\right)$ (79)
where $\epsilon:\mathfrak{X}_{\mathrm{\acute{e}t}}\to\mathfrak{X}_{\mathrm{Zar}}$ is the canonical map of sites and $\epsilon_{*}$ is the forgetful functor. Moreover, we have a quasi-isomorphism
$\tau_{\leq n+1}\left(\mathbb{Z}(n)_{\mathrm{Nis}}^{\mathfrak{X}}\right)\xrightarrow{\simeq}\tau_{\leq n+1}\left(R\alpha_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\right)$ (80)
where $\alpha:\mathfrak{X}_{\mathrm{\acute{e}t}}\to\mathfrak{X}_{\mathrm{Nis}}$ is the canonical map of sites.
Proof.
Let $i:Y\to\mathfrak{X}$ be the inclusion of the closed fiber of $\mathfrak{X}$. Then we have a quasi-isomorphism
$\tau_{\leq n+2}\left(R\epsilon_{*}\mathbb{Z}(n-1)_{\mathrm{\acute{e}t}}^{Y}[-2]\right)\xrightarrow{\simeq}\tau_{\leq n+2}i_{*}Ri^{!}\left(R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\right)$
by [23, p.33, Proposition 2.1]. Moreover, we have a quasi-isomorphism
$\mathbb{Z}(n-1)_{\mathrm{Zar}}^{Y}[-2]\xrightarrow{\simeq}i_{*}Ri^{!}\mathbb{Z}(n)_{\mathrm{Zar}}^{\mathfrak{X}}$
(cf. [7, p.780, (6)]). Let $Z$ be a scheme which is finite type over the spectrum of a field. Then we have a quasi-isomorphism
$\tau_{\leq n+1}\mathbb{Z}(n)_{\mathrm{Zar}}^{Z}\xrightarrow{\simeq}\tau_{\leq n+1}R\epsilon_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{Z}$
by Proposition 4.2. So the quasi-isomorphism (79) follows from the five lemma. Let $\beta:\mathfrak{X}_{\mathrm{Nis}}\to\mathfrak{X}_{\mathrm{Zar}}$ be the canonical map of sites. By applying $\beta^{*}$ to the quasi-isomorphism (79), we obtain the quasi-isomorphism (80). This completes the proof. ∎
Proposition 4.7.
Let $X$ be a regular scheme which is essentially finite type over the spectrum of a discrete valuation ring of mixed characteristic $(0,p)$, $Z$ an irreducible component of the closed fiber of $X$ and $i:Z\to X$ the corresponding closed immersion. Let $n$ be a non-negative integer. Then we have a quasi-isomorphism
$\tau_{\leq n+2}\left(\mathbb{Z}(n-1)_{\mathrm{\acute{e}t}}[-2]\right)\xrightarrow{\simeq}\tau_{\leq n+2}Ri^{!}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$ (81)
and a quasi-isomorphism
$\tau_{\leq n+1}\left(\mathbb{Z}/m(n-1)_{\mathrm{\acute{e}t}}[-2]\right)\xrightarrow{\simeq}\tau_{\leq n+1}Ri^{!}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}$
for any positive integer $m$.
Proof.
It suffices to show the quasi-isomorphism (81). Let $\mathcal{F}^{\bullet}$ be a bounded below complex on $X_{\mathrm{\acute{e}t}}$. Then we have an isomorphism
$\operatorname{H}^{q}\left((\mathcal{O}_{X,\bar{x}})_{\mathrm{\acute{e}t}},\mathcal{F}^{\bullet}\right)\xrightarrow{\simeq}\Gamma\left((\mathcal{O}_{X,\bar{x}})_{\mathrm{\acute{e}t}},\mathcal{H}^{q}(\mathcal{F}^{\bullet})\right)$
for any integer $q\geq 0$ where $\mathcal{O}_{X,\bar{x}}$ is the strict henselian of $X$ at a point $x\in X$. So it suffices to prove that the homomorphism
$\operatorname{H}^{q-2}(Z_{\mathrm{\acute{e}t}},\mathbb{Z}(n-1))\xrightarrow{\simeq}\operatorname{H}^{q}_{Z}(X_{\mathrm{\acute{e}t}},\mathbb{Z}(n))$
is an isomorphism for $q\leq n+2$.
Let $Y$ be the closed fiber of $X$ and $U=X\setminus Z$. Let $j:W\to Y$ be an open immersion with $W=Y\cap U$. Then we have quasi-isomorphisms
$\tau_{\leq n}\epsilon^{*}Rj_{*}\mathbb{Z}(n-1)_{\mathrm{Zar}}\simeq\tau_{\leq n}\epsilon^{*}Rj_{*}\left(\tau_{\leq n}R\epsilon_{*}\mathbb{Z}(n-1)_{\mathrm{\acute{e}t}}\right)\simeq\tau_{\leq n}Rj_{*}\mathbb{Z}(n-1)_{\mathrm{\acute{e}t}}$
by Proposition 4.2. Here $\epsilon:W_{\mathrm{\acute{e}t}}\to W_{\mathrm{Zar}}$ is the canonical map of sites. So we have a commutative diagram
$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${\operatorname{H}^{q-2}_{\mathrm{\acute{e}t}}(Z,\mathbb{Z}(n-1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q-2}_{\mathrm{\acute{e}t}}(Y,\mathbb{Z}(n-1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q-2}_{\mathrm{\acute{e}t}}(W,\mathbb{Z}(n-1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q}_{Z}(X_{\mathrm{\acute{e}t}},\mathbb{Z}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q}_{Y}(X_{\mathrm{\acute{e}t}},\mathbb{Z}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q}_{W}(U_{\mathrm{\acute{e}t}},\mathbb{Z}(n))}$
for $q\leq n+2$ by [7, p.776, Proposition 2.2.a)] and Lemma 2.13. Here the sequences are exact. Moreover, we have isomorphisms
$\operatorname{H}_{\mathrm{\acute{e}t}}^{q-2}(Y,\mathbb{Z}(n-1))\xrightarrow{\simeq}\operatorname{H}^{q}_{Y}(X_{\mathrm{\acute{e}t}},\mathbb{Z}(n))$
and
$\operatorname{H}_{\mathrm{\acute{e}t}}^{q-2}(W,\mathbb{Z}(n-1))\xrightarrow{\simeq}\operatorname{H}^{q}_{W}(U_{\mathrm{\acute{e}t}},\mathbb{Z}(n))$
for $q\leq n+2$ by [23, p.33, Proposition 2.1]. Therefore the statement follows from the five lemma. ∎
Corollary 4.8.
Let $R$ be a local ring of a regular scheme which is finite type over the spectrum of a discrete valuation ring of mixed characteristic $(0,p)$. Let $\mathfrak{m}$ be the maximal ideal of $\operatorname{Spec}(R)$, $i:\operatorname{Spec}(R/\mathfrak{m})\to\operatorname{Spec}(R)$ the corresponding closed immersion and $c=\operatorname{dim}(R)$. Let $n$ be a non-negative integer. Then we have a quasi-isomorphism
$\tau_{\leq n+2}\left(\mathbb{Z}(n-c)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(R/\mathfrak{m})}[-2c]\right)\xrightarrow{\sim}\tau_{\leq n+2}Ri^{!}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(R)}$ (82)
and a quasi-isomorphism
$\tau_{\leq n+1}\left(\mathbb{Z}/m(n-c)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(R/\mathfrak{m})}[-2c]\right)\xrightarrow{\sim}\tau_{\leq n+1}Ri^{!}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(R)}$
for any positive integer $m$.
Proof.
It suffices to show the quasi-isomorphism (82). Let $Z$ be an irreducible component of the closed fiber of $X$. Put $U=\operatorname{Spec}(R)\setminus\operatorname{Spec}(R/\mathfrak{m})$ and $W=Z\setminus\operatorname{Spec}(R/\mathfrak{m})$. Since we have a commutative diagram
$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${\operatorname{H}^{q-2c}_{\mathrm{\acute{e}t}}(\kappa(\mathfrak{m}),\mathbb{Z}(n-c))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q-2}_{\mathrm{\acute{e}t}}(Z,\mathbb{Z}(n-1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q-2}_{\mathrm{\acute{e}t}}(W,\mathbb{Z}(n-1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q}_{\mathfrak{m}}(R_{\mathrm{\acute{e}t}},\mathbb{Z}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q}_{Z}(R_{\mathrm{\acute{e}t}},\mathbb{Z}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{q}_{W}(U_{\mathrm{\acute{e}t}},\mathbb{Z}(n))}$
for $q\leq n+2$, the statement follows from Proposition 4.7. ∎
Corollary 4.9.
Let $X$ be a regular scheme which is essentially finite type over the spectrum of a discrete valuation ring of mixed characteristic $(0,p)$, $Z$ an irreducible closed subscheme of $X$ and $i:Z\to X$ the corresponding closed immersion of codimension $c$. Let $n$ be a non-negative integer. Then we have a quasi-isomorphism
$\tau_{\leq n+1}\left(\mathbb{Z}/p^{r}(n-c)^{Z}_{\mathrm{\acute{e}t}}[-2c]\right)\simeq\tau_{\leq n+1}\left(Ri^{!}\mathbb{Z}/p^{r}(n)^{X}_{\mathrm{\acute{e}t}}\right)$
for any positive integer $r$.
Proof.
It suffices to prove the statement in the case where $X$ is the spectrum of a local ring $R$ of a regular scheme which is finite type over the spectrum of a discrete valuation ring of mixed characteristic $(0,p)$. We prove the statement by induction on $\operatorname{dim}(R)$.
Assume that $\operatorname{dim}(R)\leq 2$. Then the statement follows from [23, p.33, Proposition 2.1].
Assume that the statement holds in the case where $\operatorname{dim}(R)=d\geq 2$. Suppose that $\operatorname{dim}(R)=d+1$. Put $U=\operatorname{Spec}(R)\setminus\operatorname{Spec}(R/\mathfrak{m})$ and $W=Z\setminus\operatorname{Spec}(R/\mathfrak{m})$. Since $\operatorname{char}(R/\mathfrak{m})=p>0$ and
(n2d)(nd1)=1d<0,(n-2d)-(n-d-1)=1-d<0,
we have
$\operatorname{H}^{n-2d}_{\mathrm{\acute{e}t}}\left(R/\mathfrak{m},\mathbb{Z}/p^{r}(n-d-1)\right)=0$
by [9, Theorem 8.5]. So we have a commutative diagram
$\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces$${\operatorname{H}^{n-2d-1}_{\mathrm{\acute{e}t}}(R/\mathfrak{m},\mathbb{Z}/p^{r}(n-d-1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n-2c+1}_{\mathrm{\acute{e}t}}(Z,\mathbb{Z}/p^{r}(n-c))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n-2c+1}_{\mathrm{\acute{e}t}}(W,\mathbb{Z}/p^{r}(n-c))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0}\textstyle{\cdots\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{\mathfrak{m}}^{n+1}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/p^{r}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{Z}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/p^{r}(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{W}(U_{\mathrm{\acute{e}t}},\mathbb{Z}/p^{r}(n)).}$
By using the five lemma, the statement in the case where $\operatorname{Spec}(R)=d+1$ follows from Corollary 4.8 and the assumption of induction. This completes the proof. ∎
Proposition 4.10.
Let the notations be the same as in Proposition 4.5. Then we have
$\operatorname{H}^{n+1}_{\mathrm{Zar}}(C,\mathbb{Z}/m(n))=0$
for any positive number $m$.
Proof.
Since we have an exact sequence
$\operatorname{H}^{n+1}_{\mathrm{Zar}}(C,\mathbb{Z}/m(n))\to\operatorname{H}^{n+1}_{\mathrm{Zar}}(C,\mathbb{Z}/(m+m^{\prime})(n))\to\operatorname{H}^{n+1}_{\mathrm{Zar}}(C,\mathbb{Z}/m^{\prime}(n))$
for any positive numbers $m$ and $m^{\prime}$, it suffices to prove the statement in the case where $m$ is a prime number. By [7, p.781, Corollary 3.5], it suffices to prove the statement in the case where
$C=B[T_{0},\cdots,T_{a}]/(T_{0}\cdots T_{a}-\pi)$
for any positive integer $a$. Moreover, we have exact sequences
$\operatorname{H}^{n+1}_{\mathrm{Zar}}(B[T_{0}^{\prime},\cdots,T_{a-1}^{\prime}]/(T_{0}^{\prime}\cdots T_{a-1}^{\prime}-\pi)\otimes_{B}B[T_{a}^{\prime}],\mathbb{Z}/m(n))$
$\to$ $\operatorname{H}^{n+1}_{\mathrm{Zar}}(C_{T_{a}},\mathbb{Z}/m(n))\to\operatorname{H}^{n}_{\mathrm{Zar}}(B[T_{0}^{\prime},\cdots,T_{a-1}^{\prime}]/(T_{0}^{\prime}\cdots T_{a-1}^{\prime}-\pi),\mathbb{Z}/m(n-1))$
and
$\operatorname{H}^{n}_{\mathrm{Zar}}(C_{T_{a}},\mathbb{Z}/m(n))\to\operatorname{H}^{n-1}_{\mathrm{Zar}}\left((B/(\pi))[T_{0},\cdots,T_{a-1}],\mathbb{Z}/m(n-1)\right)$
$\to$ $\operatorname{H}^{n+1}_{\mathrm{Zar}}(C,\mathbb{Z}/m(n))\to\operatorname{H}^{n+1}_{\mathrm{Zar}}(C_{T_{a}},\mathbb{Z}/m(n))$
by the isomorphism (4.2) and the localization theorem ([7, p.779, Theorem 3.2]). So, by induction on $a$, it suffices to prove that the homomorphism
$\operatorname{H}^{n}_{\mathrm{Zar}}(C_{T_{a}},\mathbb{Z}/m(n))\to\operatorname{H}^{n-1}_{\mathrm{Zar}}\left((B/(\pi))[T_{0},\cdots,T_{a-1}],\mathbb{Z}/m(n-1)\right)$
is surjective for any positive integer $a$ and any prime number $m$. Moreover, it suffices to prove that the homomorphism
$\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(C_{T_{a}},\mathbb{Z}/m(n))\to\operatorname{H}^{n-1}_{\mathrm{\acute{e}t}}\left((B/(\pi))[T_{0},\cdots,T_{a-1}],\mathbb{Z}/m(n-1)\right)$ (83)
is surjective for any positive integer $a$ and any prime number $m$ by Proposition 4.6 and Proposition 4.7.
Let $\mathfrak{X}$ be a semistable family over the spectrum of a discrete valuation ring $A$ of mixed-characteristic $(0,p)$. Then we have quasi-isomorphisms
$\tau_{\leq n}\left(\mathbb{Z}/p(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\right)\simeq\mathfrak{T}_{1}(n)^{\mathfrak{X}}$
and
$\tau_{\leq n}\left(\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\right)\simeq\mu_{m}^{\otimes n}$
in the case where $m$ is a prime number which is prime to $\operatorname{char}(B/(\pi))$ by [28, p.209, Remark 7.2] and [24, Remark 4.7]. So we have a commutative diagram
${\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(\mathfrak{X},\mathbb{Z}/m(1))\otimes\cdots\otimes\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(\mathfrak{X},\mathbb{Z}/m(1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cup}$${\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(\mathfrak{X},\mathbb{Z}/m(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(k(\mathfrak{X}),\mathbb{Z}/m(1))\otimes\cdots\otimes\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(k(\mathfrak{X}),\mathbb{Z}/m(1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}\scriptstyle{\cup}$${\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(k(\mathfrak{X}),\mathbb{Z}/m(n))}$
by [27, p.538, Proposition 4.2.6]. Moreover, we have a commutative diagram
${\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(C_{T_{a}},\mathbb{Z}/m(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n-1}_{\mathrm{\acute{e}t}}\left((B/(\pi))[T_{0},\cdots,T_{a-1}],\mathbb{Z}/m(n-1)\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(k(C_{T_{a}}),\mathbb{Z}/m(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\delta}$${\operatorname{H}^{n-1}_{\mathrm{\acute{e}t}}(k((B/(\pi))[T_{0},\cdots,T_{a-1}]),\mathbb{Z}/m(n-1))}$
where the right map is injective by [7, p.781, Corollary 3.5] and [25, Corollary 6.9]. Then an element
$\{a_{0},\cdots,a_{n-1},T_{a}\}$
$(a_{0},\cdots,a_{n-1}\in B^{*})$ maps to
$\{\bar{a_{0}},\cdots,\bar{a_{n-1}}\}$
via $\delta$. So the homomorphism (83) is surjective by the Bloch-Kato conjecture ([30]) and [3, p.113, Theorem (2.1)]. This completes the proof. ∎
Corollary 4.11.
Let $B$ be a discrete valuation ring and $\pi$ a prime element of $B$. Let
$C=B[T_{0},T_{1}]/(T_{0}T_{1}-\pi)$
and $l$ be a positive integer which is invertible in $A$. Suppose that $A$ contains $l$-th roots of unity. Then there exists an exact sequence
$0\to$ $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(C,\mu_{l}^{\otimes n})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(C)^{(0)}}\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{l}^{\otimes n})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(C)^{(1)}}\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{l}^{\otimes(n-1)})$
$\to$ $\displaystyle\bigoplus_{x\in\operatorname{Spec}(C)^{(2)}}\operatorname{H}^{n-1}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{l}^{\otimes(n-2)})\to 0$
for any positive integer $n$.
Proof.
Let $R$ be the henselization of a regular local ring which is essentially of finite type over $\operatorname{Spec}(B)$. Let $\mathfrak{m}$ be the maximal ideal of $R$ and $g\in\mathfrak{m}\setminus\mathfrak{m}^{2}$. Suppose that $\operatorname{dim}(R)=2$. Then we have
$\operatorname{H}^{q-2}_{\mathrm{Zar}}(R/(g),\mathbb{Z}/l(n-1))=\operatorname{H}^{q}_{\mathrm{Zar}}(R_{g},\mathbb{Z}/l(n))=0$
for $q\geq n+2$ by [7, p.786, Corollary 4.4] and the vanishing theorem. So we have
$\operatorname{H}^{q}_{\mathrm{Zar}}(R,\mathbb{Z}/l(n))=0$
for $q\geq n+1$ by the localization theorem ([7, p.779, Theorem 3.2]) and [22, p.37, Corollary 7]. Moreover, we have a quasi-isomorphism
$\tau_{\leq n}\left(\mathbb{Z}/l(n)_{\mathrm{Nis}}^{\operatorname{Spec}(C)}\right)\simeq\tau_{\leq n}\left(R\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(C)}\right)$
by Proposition 4.6. Hence we have a quasi-isomorphism
$\mathbb{Z}/l(n)_{\mathrm{Nis}}^{\operatorname{Spec}(C)}\simeq\tau_{\leq n}R\alpha_{*}\mu_{l}^{\otimes n}$
by [24, Remark 4.7]. Since $B$ contains $l$-th roots of unity, we have quasi-isomorphisms
$\tau_{\leq N}R\alpha_{*}\mu_{l}^{\otimes n}\simeq\tau_{\leq N}R\alpha_{*}\mu_{l}^{\otimes N}\simeq\mathbb{Z}/l(N)_{\mathrm{Nis}}^{\operatorname{Spec}(C)}$
for any integer $N\geq 0$ by the above. So we have a distinguished triangle
$\cdots\to\mathbb{Z}/l(n)_{\mathrm{Nis}}^{\operatorname{Spec}(C)}\to\mathbb{Z}/l(n+1)_{\mathrm{Nis}}^{\operatorname{Spec}(C)}\to R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}[-(n+1)]\to\cdots.$
Hence we have isomorphisms
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(C,\mu_{l}^{\otimes n})\simeq\operatorname{H}^{0}_{\mathrm{Nis}}(C,R\alpha_{*}^{n+1}\mu_{l}^{\otimes n})$ and $\operatorname{H}_{\mathrm{Nis}}^{j}(C,R\alpha_{*}^{n+1}\mu_{l}^{\otimes n})\simeq 0$
for $j>0$ by Proposition 4.5 and Proposition 4.10. Moreover, we have a flabby resolution
$0\to R\alpha_{*}^{n+1}\mu_{l}^{\otimes n}\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(C)^{(0)}}(i_{x})_{*}R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(C)^{(1)}}(i_{x})_{*}R^{n}\alpha_{*}\mu_{l}^{\otimes(n-1)}$
$\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(C)^{(2)}}(i_{x})_{*}R^{n-1}\alpha_{*}\mu_{l}^{\otimes(n-2)}\to 0$
by [22, p.38, Theorem 9]. Here we write $i_{x}$ for the canonical map $\operatorname{Spec}(\kappa(x))\to\operatorname{Spec}(C)$. Therefore the statement follows. ∎
4.3 Application
Let $B$ be a henselian valuation ring of mixed characeristic $(0,p)$, $\pi$ a prime element of $B$ and $C=B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi)$ for $a\leq N$. Then we observe a relation between $\Gamma\left(C,R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$ and $\Gamma\left(C/(\pi),R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$ for an integer $l$ which is prime to $p$.
Lemma 4.12.
Let $B$ be a discrete valuation ring and $\pi$ a prime element of $B$. Let $n\geq 0$ and $m>0$ be integers. Then we have
$\operatorname{H}^{q}_{\mathrm{Zar}}\left((B[T_{0},\cdots,T_{N}]/(T_{b}\cdots T_{a}-\pi))_{T_{0}\cdots T_{b}},\mathbb{Z}/m(n)\right)=0$
for integers $a\leq N$, $b<a$ and $q\geq n+1$.
Proof.
We have a homomorphism of polynomial rings over $B$
$B[T_{0}^{(1)},\cdots,T^{(1)}_{N}]\to B[T_{0},\cdots,T_{N}]$
which sends $T^{(1)}_{b+1}$ to $T_{b}T_{b+1}$ and sends $T^{(1)}_{r}$ to $T_{r}$ for $r\neq b+1$. Then this homomorphism induces an isomorphism
$\left(B[T^{(1)}_{0},\cdots,T^{(1)}_{N}]/(T^{(1)}_{b+1}\cdots T^{(1)}_{a}-\pi)\right)_{T^{(1)}_{0}\cdots T^{(1)}_{b}}$
$\xrightarrow{\simeq}$ $\left(B[T_{0},\cdots,T_{N}]/(T_{b}\cdots T_{a}-\pi)\right)_{T_{0}\cdots T_{b}}.$ (84)
Moreover, we have a homomorphism of polynomial rings over $B$
$B[T^{(1)}_{0},\cdots T^{(1)}_{N}]\to B[T^{(2)}_{0},\cdots T^{(2)}_{N}$]
which sends $T^{(1)}_{r}$ to $T^{(2)}_{r}$ for $0\leq r\leq b-1$, sends $T_{b}^{(1)}$ to $T_{N}^{(2)}$ and sends $T^{(1)}_{r}$ to $T^{(2)}_{r-1}$ for $b+1\leq r\leq N$. Then this homomorphism induces
$\left(B[T^{(1)}_{0},\cdots,T^{(1)}_{N}]/(T^{(1)}_{b+1}\cdots T^{(1)}_{a}-\pi)\right)_{T^{(1)}_{0}\cdots T^{(1)}_{b}}$
$\xrightarrow{\simeq}$ $\left(B[T^{(2)}_{0},\cdots,T^{(2)}_{N}]/(T^{(2)}_{b}\cdots T^{(2)}_{a-1}-\pi)\right)_{(T^{(2)}_{0}\cdots T^{(2)}_{b-1})\cdot T^{(2)}_{N}}$.
So we have an exact sequence
$\operatorname{H}^{q+1}_{\mathrm{Zar}}\left(\left(B[T^{(2)}_{0},\cdots,T^{(2)}_{N}]/(T^{(2)}_{b}\cdots T^{(2)}_{a-1}-\pi)\right)_{T^{(2)}_{0}\cdots T^{(2)}_{b-1}},\mathbb{Z}/m(n)\right)$
$\to$ $\operatorname{H}^{q+1}_{\mathrm{Zar}}\left(\left(B[T_{0},\cdots,T_{N}]/(T_{b}\cdots T_{a}-\pi)\right)_{T_{0}\cdots T_{b}},\mathbb{Z}/m(n)\right)$
$\to$ $\operatorname{H}^{q}_{\mathrm{Zar}}\left(\left(B[T^{(2)}_{0},\cdots,T^{(2)}_{N-1}]/(T^{(2)}_{b}\cdots T^{(2)}_{a-1}-\pi)\right)_{T^{(2)}_{0}\cdots T^{(2)}_{b-1}},\mathbb{Z}/m(n-1)\right)$ (85)
for any integer $q$ by the localization theorem ([7, p.779, Theorem 3.2]). By induction on $a$, the statement follows from the isomorphism (4.3) and the exact sequence (4.3). ∎
Let $B$ be a discrete valuation of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let
$C=B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi)$
for $0\leq a\leq N$ and $i:\operatorname{Spec}(C/(\pi))\to\operatorname{Spec}(C)$ be the closed fiber of $\operatorname{Spec}(C)$ and $l$ an integer which is prime to $p$. Then we have the composition
$\tau_{\leq n+1}(\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(C)})\simeq\mu_{l}^{\otimes n}\to i_{*}\mu_{l}^{\otimes n}$
$\to$ $i_{*}\left(\tau_{\leq 0}(\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(C/(\pi))})\right)\to i_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(C/(\pi))}$ (86)
by [24, Remark 4.7] and Proposition 4.3.
Proposition 4.13.
Let the notations be the same as above. Let $n$ be a non-negative integer. Suppose that $B$ is a henselian discrete valuation ring. Then the homomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi),\mathbb{Z}/l(n))$
$\to$ $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left((B/(\pi))[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}),\mathbb{Z}/l(n)\right)$ (87)
is injective where the homomorphism (4.13) is induced by the composition (4.3).
Proof.
By the localization theorem ([7, p.779, Theorem 3.2]), Lemma 4.12 and the isomorphism (4.3), the homomorphism
$\operatorname{H}^{n}_{\mathrm{Zar}}\left((B[S_{0},\cdots,S_{N}]/(S_{b}\cdots S_{a}-\pi))_{S_{0}\cdots S_{b-1}\cdot S_{b}},\mathbb{Z}/l(n)\right)$
$\to$ $\operatorname{H}^{n-1}_{\mathrm{Zar}}\left((B[S_{0},\cdots,S_{N}]/(S_{b},\pi))_{S_{0}\cdots S_{b-1}},\mathbb{Z}/l(n-1)\right)$
is surjective. So the homomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left((B[S_{0},\cdots,S_{N}]/(S_{b}\cdots S_{a}-\pi))_{S_{0}\cdots S_{b-1}},\mathbb{Z}/l(n)\right)$
$\to$ $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left((B[S_{0},\cdots,S_{N}]/(S_{b}\cdots S_{a}-\pi))_{S_{0}\cdots S_{b-1}\cdot S_{b}},\mathbb{Z}/l(n)\right)$
is injective by Proposition 4.6 and Proposition 4.7. Moreover, we have a homomorphism of polynomial ring over $B$
$B[S^{\prime}_{0},\cdots,S^{\prime}_{N}]\to B[S_{0},\cdots,S_{N}]$
which $S_{r}^{\prime}$ sends to $S_{r}$ for $r\neq b+1$ and $S_{b+1}^{\prime}$ sends to $S_{b}S_{b+1}$. Then this homomorphism induces an isomorphism
$\left(B[S^{\prime}_{0},\cdots,S^{\prime}_{N}]/(S^{\prime}_{b+1}\cdots S^{\prime}_{a}-\pi)\right)_{S^{\prime}_{0}\cdots S^{\prime}_{b}}\xrightarrow{\simeq}\left(B[S_{0},\cdots,S_{N}]/(S_{b}\cdots S_{a}-\pi)\right)_{S_{0}\cdots S_{b}}$
and so the homomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left(\left(B[S_{0},\cdots,S_{N}]/(S_{b}\cdots S_{a}-\pi)\right)_{S_{0}\cdots S_{b-1}},\mathbb{Z}/l(n)\right)$
$\to$ $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left(\left(B[S^{\prime}_{0},\cdots,S^{\prime}_{N}]/(S^{\prime}_{b+1}\cdots S^{\prime}_{a}-\pi)\right)_{S^{\prime}_{0}\cdots S^{\prime}_{b}},\mathbb{Z}/l(n)\right)$
is injective. Moreover, we have a homomorphism of polynomial rings over $B$
$B[S_{0},\cdots,S_{N}]\to B[S^{\prime}_{0},\cdots,S^{\prime}_{N}]$
which $S_{r}$ sends to $S^{\prime}_{r}$ for $0\leq r\leq a-1$, $S_{a}$ sends to $S^{\prime}_{N}+\pi$ and $S_{r}$ sends to $S^{\prime}_{r-1}$ for $a+1\leq r\leq N$. Then this homomorphism induces an isomorphism
$\left(B[S_{0}\cdots S_{N}]/(S_{a}-\pi)\right)_{S_{0}\cdots S_{a-1}}\xrightarrow{\simeq}\left(B[S^{\prime}_{0}\cdots S^{\prime}_{N}]/(S^{\prime}_{N})\right)_{S^{\prime}_{0}\cdots S^{\prime}_{a-1}}.$
So we have an injective homomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left(B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi),\mathbb{Z}/l(n)\right)$
$\to$ $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left(B[R_{0},\cdots,R_{N-1}]_{R_{0}\cdots R_{a-1}},\mathbb{Z}/l(n)\right)$
where $R_{0},\cdots,R_{N-1}$ are indeterminates over $B$. Since we have a commutative diagram
${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi),\mathbb{Z}/l(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(B[R_{0},\cdots,R_{N-1}]_{R_{0}\cdots R_{a-1}},\mathbb{Z}/l(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left((B/(\pi))[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}),\mathbb{Z}/l(n)\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left((B/(\pi))[R_{0},\cdots,R_{N-1}]_{R_{0}\cdots R_{a-1}},\mathbb{Z}/l(n)\right)}$
and the upper map is injective, it suffices to prove that the right map
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left(B[R_{0},\cdots,R_{N-1}]_{R_{0}\cdots R_{a-1}},\mathbb{Z}/l(n)\right)$
$\to$ $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left((B/(\pi))[R_{0},\cdots,R_{N-1}]_{R_{0}\cdots R_{a-1}},\mathbb{Z}/l(n)\right)$ (88)
is injective.
By the localization theorem ([7, p.779, Theorem 3.2]), [7, p.781, Corollary 3.5] and [7, p.786, Corollary 4.4], we have
$\operatorname{H}^{n+1}_{\mathrm{Zar}}\left((B/(\pi))[R_{0},\cdots R_{N-1}]_{R_{0}\cdots R_{b}},\mathbb{Z}/l(n)\right)=0$
for $0\leq b\leq a-1$ and so the homomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left((B/(\pi))[R_{0},\cdots,R_{N}]_{R_{0}\cdots R_{b}},\mathbb{Z}/l(n)\right)\to\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}\left((B/(\pi))[R_{0},\cdots,R_{N}]_{R_{0}\cdots R_{b+1}},\mathbb{Z}/l(n)\right)$
is injective for $0\leq b\leq a-1$ by the same argument as above. Let $E^{M}_{b}$ be a ring which is isomorphic to $B[R_{0},\cdots,R_{M}]_{R_{0}\cdots R_{b}}$. Then we have a commutative diagram
${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(E^{N-1}_{b-1},\mathbb{Z}/l(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(E^{N-1}_{b},\mathbb{Z}/l(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(E^{N-2}_{b-1},\mathbb{Z}/l(n-1))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${0\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(E^{N-1}_{b-1}/(\pi),\mathbb{Z}/l(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(E^{N-1}_{b}/(\pi),\mathbb{Z}/l(n))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(E^{N-2}_{b-1}/(\pi),\mathbb{Z}/l(n-1))}$
by [7, p.774, Theorem 1.2.4] and the absolute purity conjecture ([5]) where the sequences are exact. Since $A$ is a henselian discrete valuation ring, the homomorphism
$\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(B[R_{0},\cdots,R_{M}],\mathbb{Z}/l(n-1))\to\operatorname{H}^{n}_{\mathrm{\acute{e}t}}\left((B/(\pi))[R_{0},\cdots,R_{M}],\mathbb{Z}/l(n-1)\right)$
is an isomorphism by [7, p.774, Theorem 1.2.3], [17, p.240, VI, Corollary 4.20] and the isomorphism (14). Hence the homomorphism (4.3) is injective by induction on $b$. This completes the proof. ∎
Lemma 4.14.
Let $\mathfrak{X}$ be a semistable family over the spectrum of a Dedekind domain and $\alpha:\mathfrak{X}_{\mathrm{\acute{e}t}}\to\mathfrak{X}_{\mathrm{Nis}}$ the canonical map of sites. Let $m$ be any positive integer. Then the homomorphism
$\operatorname{H}^{n+2}_{\mathrm{Nis}}\left(\mathfrak{X},\tau_{\leq n}(R\alpha_{*}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}})\right)\to\operatorname{H}^{n+2}_{\mathrm{Nis}}(\mathfrak{X},\mathbb{Z}/m(n))$ (89)
is injective. Here the homomorphism (89) is induced by the composite
$\tau_{\leq n}(R\alpha_{*}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}})\xleftarrow{\simeq}\tau_{\leq n}(\mathbb{Z}/m(n)_{\mathrm{Nis}})\to\mathbb{Z}/m(n)_{\mathrm{Nis}}$
where the first map is a quasi-isomorphism by Proposition 4.6.
Proof.
By [22, p.37, Corollary 7] and [24, Proposition 4.5], we have a quasi-isomorphism
$\tau_{\leq n+1}\left(\mathbb{Z}/m(n)_{\mathrm{Nis}}^{\mathfrak{X}}\right)\xrightarrow{\simeq}\tau_{\leq n}\left(R\alpha_{*}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\right).$
Hence the statement follows. ∎
Corollary 4.15.
Let $B$ be a henselian discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let $n$ be a non-negative integer and $l$ an integer which is prime to $p$. Let
$C=B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi)$
for $0\leq a\leq N$ and $\alpha:\operatorname{Spec}(C)_{\mathrm{\acute{e}t}}\to\operatorname{Spec}(C)_{\mathrm{Nis}}$ the canonical map of sites. Then the homomorphism
$\Gamma\left(\operatorname{Spec}(C),R^{n+1}\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}\right)\to\Gamma\left(\operatorname{Spec}(C/(\pi)),R^{n+1}\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}\right)$
is injective where the homomorphism is induced by (4.3).
Proof.
By Proposition 4.13, it suffices to show that the homomorphisms
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(C,\mathbb{Z}/l(n))\xrightarrow{\simeq}\Gamma\left(\operatorname{Spec}(C),R^{n+1}\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}\right)$ (90)
and
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(C/(\pi),\mathbb{Z}/l(n))\xrightarrow{\simeq}\Gamma\left(\operatorname{Spec}(C/(\pi)),R^{n+1}\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}\right)$ (91)
are isomorphisms. Since the sequence
$\operatorname{H}^{n+1}_{\mathrm{Nis}}\left(C,\tau_{\leq n}(R\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}})\right)\to\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(C,\mathbb{Z}/l(n))$
$\to$ $\Gamma\left(\operatorname{Spec}(C),R^{n+1}\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}\right)\to\operatorname{H}^{n+2}_{\mathrm{Nis}}\left(C,\tau_{\leq n}(R\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}})\right)$
is exact, the isomorphism (90) follows from Proposition 4.10, Lemma 4.14 and Proposition 4.5. Since we have a quasi-isomorphism
$\tau_{\leq n}\left(R\alpha_{*}\mathbb{Z}/l(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(C/(\pi))}\right)\simeq\mathbb{Z}/l(n)_{\mathrm{Nis}}^{\operatorname{Spec}(C/(\pi))}$
by Proposition 4.2 and [24, Proposition 2.1], the isomorphism (91) follows from Proposition 4.1. This completes the proof. ∎
Lemma 4.16.
Let $A$ be a henselian local ring with $\operatorname{dim}(A)=1$ and $l$ an integer which is prime to $\operatorname{char}(A)$. Suppose that the maximal ideal of $A$ is principal. Then the homomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(A,\mu_{l}^{\otimes n})\to\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{l}^{\otimes n})$
is injective for any integer $n\geq 0$ and $x\in\operatorname{Spec}(A)^{(0)}$.
Proof.
Consider a commutative diagram
${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(A,\mu_{l}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{l}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(A/x,\mu_{l}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{l}^{\otimes n})}$
where $x\in\operatorname{Spec}(A)^{(0)}$. Since $A/x$ is a discrete valuation ring by [16, p.78, Theorem 11.1], the bottom map is injective by [4, Theorem B.2.1 and Examples B.1.1.(2)]. Moreover, the left map is an isomorphism by [7, p.777, The proof of Proposition 2.2.b)]. So the upper map is injective. This completes the proof. ∎
Lemma 4.17.
Let $B$ be a discrete valuation ring and $\pi$ a prime element of $B$. Let $R$ be the henselization of the local ring of $C=B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi T_{N}^{b})$ at a point $x$ of $\operatorname{Spec}(C)$ and $l$ an integer which is prime to $\operatorname{char}(B)$. Then the homomorphism
$\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(R,\mu_{l}^{\otimes n})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(R)^{(0)}}\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{l}^{\otimes n})$ (92)
is injective for any integer $n\geq 0$.
Proof.
Let $\mathfrak{p}\in\operatorname{Spec}(B[T_{0},\cdots,T_{N}])$ be the image of $x\in\operatorname{Spec}(C)$ under $\operatorname{Spec}(C)\to\operatorname{Spec}(B[T_{0},\cdots,T_{N}])$. Then $T_{0}\cdots T_{a}-\pi T_{N}^{b}\in\mathfrak{p}.$
First we assume that $T_{N}\notin\mathfrak{p}$. We have a homomorphism of polynomial rings over $B$
$B[T_{0},\cdots,T_{N}]\to B[T_{0}^{\prime},\cdots,T_{N}^{\prime}]$
which sends $T_{i}$ to $T_{i}^{\prime}$ for $i\neq 0$ and sends $T_{0}$ to $T_{0}^{\prime}(T_{N}^{\prime})^{b}$. Then this homomorphism induces an isomorphism
$\left(B[T_{0},\cdots,T_{N}]/(T_{0}\cdots T_{a}-\pi T_{N}^{b})\right)_{T_{N}}\simeq\left(B[T_{0}^{\prime},\cdots,T_{N}^{\prime}]/(T^{\prime}_{0}\cdots T_{a}^{\prime}-\pi)\right)_{T_{N}^{\prime}}$
and so $R$ is the henselization of the local ring of a semistable family over $\operatorname{Spec}(B)$. By [22, p.35, Proposition 5], the homomorphism (92) is injective in the case where $T_{N}\notin\mathfrak{p}$.
Next we assume that $T_{N}\in\mathfrak{p}$. Then $T_{i}\in\mathfrak{p}$ for some $i$ ($0\leq i\leq a$) and we may assume that $T_{0}\in\mathfrak{p}$. So we have $(T_{0},T_{N})\subset\mathfrak{p}$. Since $T_{1},\cdots,T_{a}$ are invertible elements of
$(B[T_{0},\cdots,T_{N}]_{\mathfrak{p}})_{(T_{0},T_{N})}=B[T_{0},\cdots,T_{N}]_{(T_{0},T_{N})},$
so $T_{1},\cdots,T_{a}$ are also invertible elements of
$(B[T_{0},\cdots,T_{N}]_{\mathfrak{p}}^{h})_{(T_{0},T_{N})}^{h}$
where $D^{h}_{\mathfrak{q}}$ is the henselization of the local ring of a ring $D$ at a point $\mathfrak{q}\in\operatorname{Spec}(D)$. Hence
$\left(T_{N},T_{0}\cdots T_{a}-\pi T_{N}^{b}\right)=\left(T_{N},T_{0}\cdots T_{a}\right)=(T_{0},T_{N})$
in $(B[T_{0},\cdots,T_{N}]_{\mathfrak{p}}^{h})_{(T_{0},T_{N})}^{h}$ and the maximal ideal of
$E=\left(B[T_{0},\cdots,T_{N}]_{\mathfrak{p}}^{h}/(T_{0}\cdots T_{a}-\pi T_{N}^{b})\right)_{(T_{0},T_{N})}^{h}$
is principal. Since the residue field $\kappa((T_{0},T_{N}))$ of $(T_{0},T_{N})\in\operatorname{Spec}(R)$ agrees with $E/(T_{0},T_{N})$, we have a commutative diagram
${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(R,\mu_{l}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(E,\mu_{l}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(R^{\prime},\mu_{l}^{\otimes n})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\kappa((T_{0},T_{N})),\mu_{l}^{\otimes n})}$
where $R^{\prime}=R/(T_{0},T_{N})$ and so $R^{\prime}$ is the henselization of the local ring of $B[T_{1},\cdots,T_{N-1}]$. Since the left map is an isomorphism by the isomorphism (14) and the bottom map is injective by [22, p.35, Proposition 5], the upper map is injective. Moreover, the homomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(E,\mu_{l}^{\otimes n})\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(E)^{(0)}}\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\kappa(x),\mu_{l}^{\otimes n})$
is injective by Lemma 4.16. Hence, the homomorphism (92) is injective in the case where $T_{N}\in\mathfrak{p}$. This completes the proof. ∎
Corollary 4.18.
Let $B$ be a henselian discrete valuation ring of mixed characteristic $(0,p)$ and $\pi$ a prime element of $B$. Let
$\mathfrak{X}=\operatorname{Proj}\left(B[T_{0},\cdots,T_{N+1}]/(T_{0}\cdots T_{a}-\pi T_{N+1}^{a+1})\right)$
for $0\leq a\leq N$ and $Y$ the closed fiber of $\mathfrak{X}$. Then the homomorphism
$\Gamma\left(\mathfrak{X},R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)\to\Gamma\left(Y,R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$ (93)
is injective for any integers $n\geq 0$ and $l>0$ with $(l,p)=1$. Here $\alpha:\mathfrak{X}_{\mathrm{\acute{e}t}}\to\mathfrak{X}_{\mathrm{Nis}}$ is the canonical map of sites.
Proof.
Let $j:D(T_{N+1})\to\mathfrak{X}$ be an open immersion of $\mathfrak{X}$ such that
$D(T_{N+1}):=\{\mathfrak{p}\in\mathfrak{X}|T_{N+1}\notin\mathfrak{p}\}=\operatorname{Spec}(B[S_{0},\cdots,S_{N}]/(S_{0}\cdots S_{a}-\pi)).$
Then the homomorphism
$R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\to j_{*}j^{*}R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}$
is injective by Lemma 4.17. So the homomorphism
$\Gamma\left(\mathfrak{X},R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)\to\Gamma\left(\operatorname{Spec}(C),R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)$ (94)
is injective where
$C=B[S_{0},\cdots,S_{N}]/(S_{0}\cdots S_{a}-\pi).$
Then the composite of the homomorphism (94) and the homomorphism
$\Gamma\left(\operatorname{Spec}(C),R^{n+1}\alpha_{*}\mathbb{Z}/l(n)\right)\to\Gamma\left(\operatorname{Spec}(C/(\pi)),R^{n+1}\alpha_{*}\mathbb{Z}/l(n)\right)$
agrees with the composite of the homomorphism (93) and the composition
$\Gamma\left(Y,R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)\to\Gamma\left(\operatorname{Spec}(C/(\pi)),R^{n+1}\alpha_{*}\mu_{l}^{\otimes n}\right)\to\Gamma\left(\operatorname{Spec}(C/(\pi)),R^{n+1}\alpha_{*}\mathbb{Z}/l(n)\right)$
(cf. Proposition 4.3). Hence the statement follows from Corollary 4.15. ∎
Remark 4.19.
Let the notations be the same as above and $i:Y\to\mathfrak{X}$ the inclusion of the closed fiber. Then we have an isomorphism
$\Gamma\left(Y,i^{*}R^{n+1}\alpha_{*}\mu^{\otimes n}_{l}\right)\xrightarrow{\simeq}\Gamma\left(Y,R^{n+1}\alpha_{*}\mu^{\otimes n}_{l}\right)$
by [7, p.776, Proposition 2.2.b)].
5 Proper base change theorem
5.1 A generalization of Artin’s theorem
In this subsection, we prove a generalization of Artin’s theorem ([11, p.98, Théorème (3.1)]). For a scheme $X$, $\alpha:X_{\mathrm{\acute{e}t}}\to X_{\mathrm{Nis}}$ denotes the canonical map of sites.
Lemma 5.1.
Let $Y$ be a one dimensional scheme. Then we have an isomorphism
$\operatorname{H}^{2}_{\mathrm{\acute{e}t}}(Y,\mathbb{G}_{m})\xrightarrow{\simeq}\Gamma\left(Y,R^{2}\alpha_{*}\mathbb{G}_{m}\right).$
Proof.
By [17, p.124, III, Proposition 4.9], we have an isomorphism
$R^{1}\alpha_{*}\mathbb{G}_{m}\simeq 0$
and so we have a distinguished triangle
$\cdots\to\tau_{\leq 0}R\alpha_{*}\mathbb{G}_{m}^{Y}\to R\alpha_{*}\mathbb{G}_{m}^{Y}\to\tau_{\geq 2}R\alpha_{*}\mathbb{G}_{m}^{Y}\to\cdots.$
By [19, pp.279–280, 1.32. Theorem], we have an isomorphism
$\operatorname{H}^{q}_{\mathrm{Nis}}(Y,\tau_{\leq 0}R\alpha_{*}\mathbb{G}_{m})=0$
for $q\geq 2$. Hence the statement follows. ∎
Theorem 5.2.
(cf. Artin’s theorem [11, p.98, Théorème (3.1)]) Let $B$ be a henselian excellent discrete valuation ring, $\mathfrak{X}$ a two dimensional regular scheme which is proper and flat over $\operatorname{Spec}(B)$ and $Y$ the closed fiber of $\mathfrak{X}$.
Then the canonical map
$\Gamma\left(\mathfrak{X},R^{2}\alpha_{*}\mathbb{G}_{m}\right)\to\Gamma\left(Y,R^{2}\alpha_{*}\mathbb{G}_{m}\right)$
is an isomorphism.
Proof.
By the same argument as in the proof of Lemma 5.1, we have a distinguished triangle
$\cdots\to\tau_{\leq 0}R\alpha_{*}\mathbb{G}_{m}^{\mathfrak{X}}\to R\alpha_{*}\mathbb{G}_{m}^{\mathfrak{X}}\to\tau_{\geq 2}R\alpha_{*}\mathbb{G}_{m}^{\mathfrak{X}}\to\cdots.$
Moreover, we have a commutative diagram
${\operatorname{H}^{2}_{\mathrm{\acute{e}t}}(\mathfrak{X},\mathbb{G}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Gamma\left(\mathfrak{X},R^{2}\alpha_{*}\mathbb{G}_{m}\right)\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}^{2}_{\mathrm{\acute{e}t}}(k(\mathfrak{X}),\mathbb{G}_{m})\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\Gamma\left(k(\mathfrak{X}),R^{2}\alpha_{*}\mathbb{G}_{m}\right).}$
Then the left map is injective by [17, p.145, IV, Corollary 2.6]. Moreover, the bottom map is an isomorphism. So the homomorphism
$\operatorname{H}^{2}_{\mathrm{\acute{e}t}}(\mathfrak{X},\mathbb{G}_{m})\to\Gamma\left(\mathfrak{X},R^{2}\alpha_{*}\mathbb{G}_{m}\right)$ (95)
is injective. Moreover, we have
$\operatorname{H}^{3}_{\mathrm{Nis}}(\mathfrak{X},\tau_{\leq 0}R\alpha_{*}\mathbb{G}_{m})=0$
by [19, pp.279–280, 1.32. Theorem]. So the homomorphism (95) is an isomorphism. Hence the statement follows from [11, p.98, Théorème (3.1)] and Lemma 5.1. ∎
Lemma 5.3.
Let $B$ be a local ring and $X$ a scheme over $\operatorname{Spec}(B)$. Let $i:Y\to X$ be the inclusion of the closed fiber of $X$ with open complement $j:U\to X$ and $\mathcal{F}^{\bullet}$ a bounded below complex of sheaves on $X_{\mathrm{\acute{e}t}}$. Then we have a spectral sequence
$E^{s,t}_{2}=\operatorname{H}^{s}_{\mathrm{Nis}}(X,j_{!}j^{*}R^{t}\alpha_{*}\mathcal{F}^{\bullet})\Rightarrow E^{s+t}=\operatorname{H}^{s+t}_{\mathrm{\acute{e}t}}(X,j_{!}j^{*}\mathcal{F}^{\bullet}).$ (96)
Suppose that $B$ is a henselian local ring, $X$ is proper over $\operatorname{Spec}(B)$ and $\mathcal{H}^{t}(\mathcal{F}^{\bullet})$ are torsion sheaves for any integer $t$. Then we have
$E^{n}=\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(X,j_{!}j^{*}\mathcal{F}^{\bullet})=0$ (97)
Proof.
By [7, p.776, Proposition 2.2], we have a quasi-isomorphism
$i_{*}i^{*}R\alpha_{*}\mathcal{F}^{\bullet}\simeq R\alpha_{*}i_{*}i^{*}\mathcal{F}^{\bullet}.$
Since $_{*}, ii^{*}$, $j_{!}$ and $j^{*}$ are exact (cf.[17, p.76, II, Proposition 3.14 (b)]), we have a distinguished triangle
$\cdots\to j_{!}j^{*}\mathcal{G}^{\bullet}\to\mathcal{G}^{\bullet}\to i_{*}i^{*}\mathcal{G}^{\bullet}\to\cdots$
for a bounded below complex $\mathcal{G}^{\bullet}$ of sheaves on $X_{\mathrm{\acute{e}t}}$. So we have a quasi-isomorphism
$j_{!}j^{*}R\alpha_{*}\mathcal{F}^{\bullet}\simeq R\alpha_{*}j_{!}j^{*}\mathcal{F}^{\bullet}$
and the spectral sequence (96).
Suppose that $B$ is a henselian local ring, $X$ is proper over $\operatorname{Spec}(B)$ and $\mathcal{H}^{t}(\mathcal{F}^{\bullet})$ are torsion sheaves for any integer $t$. Then we have the equation (97) by a spectral sequence
$E^{s,t}_{2}=\operatorname{H}^{s}_{\mathrm{\acute{e}t}}(X,j_{!}j^{*}\mathcal{H}^{t}(\mathcal{F}^{\bullet}))\Rightarrow E^{s+t}=\operatorname{H}^{s+t}_{\mathrm{\acute{e}t}}(X,j_{!}j^{*}\mathcal{F}^{\bullet})$
and [17, p.224, VI, Corollary 2.7]. ∎
Lemma 5.4.
Let $B$ be a discrete valuation ring of mixed characteristic $(0,p)$, $X$ a semistable family over $\operatorname{Spec}(B)$ and $Y$ the closed fiber of $X$. Suppose that $\operatorname{dim}(X)=2$ and $B$ contains $p$-th roots of unity. Then we have
$\operatorname{H}^{s}_{\mathrm{Nis}}(X,R^{t}\alpha_{*}\mathfrak{T}_{1}(n))=\dfrac{\operatorname{Ker}\left(\displaystyle\bigoplus_{x\in X^{(s)}}\operatorname{H}^{s+t}_{x}(X,\mathfrak{T}_{1}(n))\to\bigoplus_{x\in X^{(s+1)}}\operatorname{H}^{s+t+1}_{x}(X,\mathfrak{T}_{1}(n))\right)}{\operatorname{Im}\left(\displaystyle\bigoplus_{x\in X^{(s-1)}}\operatorname{H}^{s+t-1}_{x}(X,\mathfrak{T}_{1}(n))\to\bigoplus_{x\in X^{(s)}}\operatorname{H}^{s+t}_{x}(X,\mathfrak{T}_{1}(n))\right)}$
and
$\operatorname{H}^{s}_{\mathrm{Nis}}(Y,R^{t}\alpha_{*}\lambda_{1}^{n})=\dfrac{\operatorname{Ker}\left(\displaystyle\bigoplus_{y\in Y^{(s)}}\operatorname{H}^{s+t}_{y}(Y,\lambda_{1}^{n})\to\bigoplus_{y\in Y^{(s+1)}}\operatorname{H}^{s+t+1}_{y}(Y,\lambda_{1}^{n})\right)}{\operatorname{Im}\left(\displaystyle\bigoplus_{y\in Y^{(s-1)}}\operatorname{H}^{s+t-1}_{y}(Y,\lambda_{1}^{n})\to\bigoplus_{y\in Y^{(s)}}\operatorname{H}^{s+t}_{y}(Y,\lambda_{1}^{n})\right)}$
for any non-negative integers $s$, $t$ and $n$.
Proof.
By the same argument as in the proof of [25, Lemma 5.3], the statement follows from Corollary 3.4 and Theorem 2.26. ∎
Theorem 5.5.
Let $B$ be a henselian excellent discrete valuation ring of mixed characteristic $(0,p)$ and $\mathfrak{X}$ a semistable family and proper over $\operatorname{Spec}(B)$. Let $i:Y\to\mathfrak{X}$ be the inclusion of the closed fiber of $\mathfrak{X}$. Suppose that $\operatorname{dim}(\mathfrak{X})=2$ and $B$ contains $p$-th roots of unity. Then the homomorphism
$\operatorname{H}^{s}_{\mathrm{Nis}}(\mathfrak{X},R^{t}\alpha_{*}\mathfrak{T}_{1}(n))\xrightarrow{\simeq}\operatorname{H}^{s}_{\mathrm{Nis}}(Y,i^{*}R^{t}\alpha_{*}\mathfrak{T}_{1}(n))$ (98)
is an isomorphism for integers $s\geq 0$ and $t\geq 2$. Moreover, we have an isomorphism
$\operatorname{H}^{s}_{\mathrm{Nis}}(\mathfrak{X},R^{n+1}\alpha_{*}\mathfrak{T}_{1}(n))\xrightarrow{\simeq}\operatorname{H}^{s}_{\mathrm{Nis}}(Y,R^{1}\alpha_{*}\lambda_{1}^{n})$ (99)
for integers $s\geq 0$ and $n\geq 1$. Thus, the sequence
$0\to\displaystyle\bigoplus_{x\in\mathfrak{X}^{(0)}}\operatorname{H}_{x}^{n+r}(\mathfrak{X}_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to\displaystyle\bigoplus_{x\in\mathfrak{X}^{(1)}}\operatorname{H}_{x}^{n+r+1}(\mathfrak{X}_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))$
$\to\displaystyle\bigoplus_{x\in\mathfrak{X}^{(2)}}\operatorname{H}_{x}^{n+r+2}(\mathfrak{X}_{\mathrm{\acute{e}t}},\mathfrak{T}_{1}(n))\to 0$ (100)
is exact for integers $n\geq 1$ and $r\geq 2$.
Proof.
Assume that the isomorphism (98) holds. By [24, Theorem 1.2 and Theorem 1.4], we have an isomorphism
$i^{*}R^{n+1}\alpha_{*}\mathfrak{T}_{1}(n)\simeq R^{1}\alpha_{*}\lambda_{1}^{n}$ (101)
for any integer $n\geq 0$. So the isomorphism (99) follows from the isomorphism (98). Moreover, we have an isomorphism
$i^{*}R^{t}\alpha_{*}\mathfrak{T}_{1}(n)\simeq 0$
for $t\geq n+2$ by the analog results of [17, p.69, II, Theorem 3.2] and [17, p.88, III, Theorem 1.15] for the Nisnevich sheaves and [1, Exposé X, Théorème 5.1]. So we have an isomorphism
$\operatorname{H}^{s}_{\mathrm{Nis}}(\mathfrak{X},R^{t}\alpha_{*}\mathfrak{T}_{1}(n))\simeq 0$
for integers $s\geq 0$ and $t\geq n+2$ by the isomorphism (98). So the sequence (5.5) is exact by Lemma 5.4. Hence it suffices to prove the isomorphism (98).
Let $j:U\to\mathfrak{X}$ be the inclusion of the generic fiber of $\mathfrak{X}$. Since $j^{*}$ is exact by [17, p.76, II, Proposition 3.14 (b)], we have isomorphisms
$j^{*}R^{t}\alpha_{*}\mathfrak{T}_{1}(n)\simeq R^{t}(j^{*}\alpha_{*})\mathfrak{T}_{1}(n)\simeq R^{t}(\alpha_{*}j^{*})\mathfrak{T}_{1}(n)\simeq R^{t}\alpha_{*}\mu_{p}^{\otimes n}$
by [7, p.776, Proposition 2.2 a)]. So the sequence
$0\to j_{!}R^{t}\alpha_{*}\mu_{p}^{\otimes n}\to R^{t}\alpha_{*}\mathfrak{T}_{1}(n)\to i_{*}i^{*}R^{t}\alpha_{*}\mathfrak{T}_{1}(n)\to 0$
is exact. Hence the isomorphism (98) follows from the isomorphism
$\operatorname{H}^{s}_{\mathrm{Nis}}(\mathfrak{X},j_{!}R^{t}\alpha_{*}\mu_{p})\simeq 0$ (102)
for any integers $s\geq 0$ and $t\geq 2$. Hence it suffices to prove the isomorphism (102).
By [19, p.279–280, 1.32. Theorem], we have an isomorphism (102) for $s\geq 3$.
By [23, p.31, Theorem 1.2] (or [24, Theorem 1.4]), we have an exact sequence
$0\to\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(\kappa(x),\lambda_{1}^{n})\to\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(k(\mathcal{O}_{\mathfrak{X},x}^{h}),\mathfrak{T}_{1}(n))\to\operatorname{H}^{n+2}_{x}(\mathfrak{X},\mathfrak{T}_{1}(n))$
for $x\in\mathfrak{X}^{(1)}\cap Y$. Here $\mathcal{O}_{\mathfrak{X},x}^{h}$ is the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of $\mathfrak{X}$ at $x\in\mathfrak{X}^{(1)}\cap Y$ and $k(\mathcal{O}^{h}_{\mathfrak{X},x})$ is the fraction field of $\mathcal{O}^{h}_{\mathfrak{X},x}$. So the homomorphism
$\operatorname{H}^{0}_{\mathrm{Nis}}(\mathfrak{X},R^{n+1}\alpha_{*}\mathfrak{T}_{1}(n))\to\operatorname{H}^{0}_{\mathrm{Nis}}(Y,R^{1}\alpha_{*}\lambda_{1}^{n})$
is injective for $n\geq 1$ by Lemma 5.4 and [24, Theorem 1.7]. Hence we have an isomorphism (102) for $s=0$ and $t\geq 2$ by the isomorphism (101).
Moreover, we have an isomorphism
$\operatorname{H}_{\mathrm{\acute{e}t}}^{s}(\mathfrak{X},j_{!}j^{*}\mathfrak{T}_{1}(n))\simeq 0$ (103)
for any integer $s$ by Lemma 5.3. Since we have an isomorphism (102) for $s\neq 1,2$ and $t\geq 2$ by the above, we have an isomorphism (102) for $s=1,2$ and $t\geq 2$ by the spectral sequence (96) and the isomorphism (103). This completes the proof.
Remark 5.6.
Let the notations be the same as in Theorem 5.5.
  1. 1.
    Let $l$ be an integer which is prime to $p$. Then the natural map
    $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(k(\mathfrak{X}),\mu_{l}^{\otimes n})\to\displaystyle\bigoplus_{\mathfrak{p}\in\mathfrak{X}^{(1)}\setminus Y^{(0)}}\operatorname{H}^{n}_{\mathrm{\acute{e}t}}(\kappa(\mathfrak{p}),\mu_{l}^{\otimes(n-1)})\oplus\displaystyle\bigoplus_{\mathfrak{p}\in Y^{(0)}}\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(k(\tilde{\mathcal{O}_{\mathfrak{X},\mathfrak{p}}}),\mu_{l}^{\otimes n})$
    is injective (cf. [24, The remark below Theorem 1.7]). Here $\tilde{\mathcal{O}_{\mathfrak{X},\mathfrak{p}}}$ is the henselization of a local ring $\mathcal{O}_{\mathfrak{X},\mathfrak{p}}$ of $\mathfrak{p}$ in $\mathfrak{X}$. So we have an isomorphism
    $\operatorname{H}^{s}_{\mathrm{Nis}}(\mathfrak{X},R^{t}\alpha_{*}\mu_{l}^{\otimes n})\xrightarrow{\simeq}\operatorname{H}^{s}_{\mathrm{Nis}}(Y,i^{*}R^{t}\alpha_{*}\mu_{l}^{\otimes n})$
    for integers $s\geq 0$ and $t\geq 2$ by the same argument as in the proof of Theorem 5.5. Since we have quasi-isomorphisms
    $\tau_{\leq n+1}\left(\mathbb{Z}/p^{r}(n)^{\mathfrak{X}}_{\mathrm{\acute{e}t}}\right)\simeq\mathfrak{T}_{r}(n)$
    and
    $\tau_{\leq n+1}\left(\mathbb{Z}/l(n)^{\mathfrak{X}}_{\mathrm{\acute{e}t}}\right)\simeq\mu_{l}^{\otimes n}$
    for an integer with $(l,p)=1$ by ([28, p.209, Remark 7.2], [24, Proposition 4.5]) and [24, Remark 4.7], we have an isomorphism
    $\operatorname{H}^{s}_{\mathrm{Nis}}(\mathfrak{X},R^{n+1}\alpha_{*}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}})\xrightarrow{\simeq}\operatorname{H}^{s}_{\mathrm{Nis}}(Y,i^{*}R^{n+1}\alpha_{*}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}})$ (104)
    for integers $n\geq 1$, $s\geq 0$ and a prime number $m$ in the case where $B$ contains $m$-th roots of unity.
  2. 2.
    By [24, Proposition 4.2] and [22, p.35, Proposition 5], there exists an exact sequence
    $0\to R^{n+1}\alpha_{*}\mathbb{Z}/m(n)^{\mathfrak{X}}_{\mathrm{\acute{e}t}}\to R^{n+2}\alpha_{*}\mathbb{Z}(n)^{\mathfrak{X}}_{\mathrm{\acute{e}t}}\xrightarrow{\times m}R^{n+2}\alpha_{*}\mathbb{Z}(n)^{\mathfrak{X}}_{\mathrm{\acute{e}t}}$
    for any integers $n\geq 0$ and $m>0$. Moreover, we have a quasi-isomorphism
    $\mathbb{Z}(1)_{\mathrm{\acute{e}t}}^{\mathfrak{X}}\simeq\mathbb{G}_{m}[-1]$
    by [15, Lemma 11.2]. Thus, we have an isomorphism
    $i^{*}R^{2}\alpha_{*}\mathbb{G}_{m}^{\mathfrak{X}}\simeq R^{2}\alpha_{*}\mathbb{G}_{m}^{Y}$
    by [7, p.776, Proposition 2.2.b)] and [17, p.148, IV, Corollary 2.13]. So we can regard the isomorphism (104) for $s=0$ and $n=1$ as a generalization of Theorem 5.2.
5.2 Questions
In this subsection, we raise questions (see Question 5.10 and Question 5.18) which relate to Kato conjecture (cf. [13], [14]).
Lemma 5.7.
Let $X$ be an integral scheme and $\mathcal{F}$ a constant sheaf on $X_{\mathrm{Nis}}$. Then we have
$\operatorname{H}^{q}_{\mathrm{Nis}}(X,\mathcal{F})=0$
for $q>0$.
Proof.
Since $\mathcal{F}$ is a constant sheaf, we have an isomorphism
$\mathcal{F}\simeq(i_{\eta})_{*}(i_{\eta})^{*}\mathcal{F}$
where $i_{\eta}:\operatorname{Spec}(\kappa(\eta))\to X$ is the generic point. Since $(i_{\eta})^{*}\mathcal{F}$ is a flabby sheaf, the statement follows from [17, p.89, III, Lemma 1.19]. ∎
Lemma 5.8.
Let $X$ be a smooth scheme over the spectrum of a regular ring $B$ with $\operatorname{dim}(B)\leq 1$. Then we have
$\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(X,\mathbb{Q}/\mathbb{Z})=\Gamma\left(X,R^{1}\alpha_{*}\mathbb{Q}/\mathbb{Z}\right)$ (105)
where $\alpha:X_{\mathrm{\acute{e}t}}\to X_{\mathrm{Nis}}$ is the canonical map of sites. Moreover, the sequence
$0\to\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(X,\mathbb{Q}/\mathbb{Z})\to\bigoplus_{x\in X^{(0)}}\operatorname{H}^{1}_{x}(X_{\mathrm{\acute{e}t}},\mathbb{Q}/\mathbb{Z})\to\bigoplus_{x\in X^{(1)}}\operatorname{H}^{2}_{x}(X_{\mathrm{\acute{e}t}},\mathbb{Q}/\mathbb{Z})$ (106)
is exact.
Proof.
Since we have a distinguished triangle
$\cdots\to\mathbb{Q}/\mathbb{Z}\to R\alpha_{*}\mathbb{Q}/\mathbb{Z}\to\tau_{\geq 1}\alpha_{*}\mathbb{Q}/\mathbb{Z}\to\cdots,$
we have the isomorphism (105) by Lemma 5.7. So the sequence (106) is exact by [23, p.30, Theorem 1.1]. This completes the proof. ∎
Proposition 5.9.
Let $\mathfrak{X}$ be a proper and smooth scheme over the spectrum of a henselian discrete valuation ring and $i:Y\to\mathfrak{X}$ the inclusion of the closed fiber. Then the homomorphism
$\Gamma\left(\mathfrak{X},R^{1}\alpha_{*}\mathbb{Q}/\mathbb{Z}\right)\to\Gamma\left(Y,i^{*}R^{1}\alpha_{*}\mathbb{Q}/\mathbb{Z}\right)$
is an isomorphism.
Proof.
The statement follows from the proper base change theorem (cf. [17, p.224, VI, Corollary 2.7]) and Lemma 5.8. ∎
Question 5.10.
Let $\mathfrak{X}$ be a regular scheme which is proper and flat over the spectrum of a henselian discrete valuation ring $B$ and $i:Y\to\mathfrak{X}$ the closed fiber of $\mathfrak{X}$. Then
(a) When is the homomorphism
$\Gamma\left(\mathfrak{X},R^{n+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)\to\Gamma\left(Y,i^{*}R^{n+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$ (107)
an isomorphism ?
(b) If $\mathfrak{X}$ is smooth over $\operatorname{Spec}(B)$, then the homomorphism (107) is an isomorphism ?
Remark 5.11.
Let $\mathcal{O}_{K}$ be the ring of integers of a $p$-adic field $K$, $\mathfrak{X}$ a $d$-dimensional regular scheme which is proper and smooth over $\operatorname{Spec}(\mathcal{O}_{K})$ and $i:Y\to\mathfrak{X}$ the inclusion of the closed fiber of $\mathfrak{X}$. Then
$i^{*}R^{d+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(d)^{\mathfrak{X}}_{\mathrm{\acute{e}t}}=R^{d+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(d)^{Y}_{\mathrm{\acute{e}t}}=0$
and so we have
$\Gamma\left(Y,i^{*}R^{d+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(d)_{\mathrm{\acute{e}t}}\right)=0$
where $\alpha:\mathfrak{X}_{\mathrm{\acute{e}t}}\to\mathfrak{X}_{\mathrm{Nis}}$ is the canonical map of sites and $\alpha_{*}$ is the forgetful functor. On the other hand, if Kato conjecture holds (cf.[14, p.125, Conjecture 0.2], [14, p.125, Theorem 0.4]), then
$\Gamma\left(\mathfrak{X},R^{d+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(d)_{\mathrm{\acute{e}t}}\right)=0$
by [23, p.51, Theorem 4.6].
For a regular scheme of finite type over the ring of integers of a number field, we have the following conjectures:
Conjecture 5.12.
([21, Conjecture 2.1]) For a regular scheme $X$ of finite type over $\mathbb{F}_{p}$ or $\mathbb{Z}$, $\operatorname{H}_{\mathrm{Zar}}^{q}(X,\mathbb{Z}(r))$ is finitely generated.
Conjecture 5.13.
(Lichtenbaum) (cf. [8, p.1625, Conjecture 4.12]) If $X$ is regular and proper over the ring of integers of a number field, then the groups $\operatorname{H}^{i}_{\mathrm{\acute{e}t}}(X,\mathbb{Z}(n))$ are finitely generated for $i\leq 2n$, finite for $i=2n+1$, and of cofinite type for $i\geq 2n+2$.
A consequence of Conjectures 5.12 and 5.13 will be observed later.
Lemma 5.14.
Let $X$ be a regular scheme which is an essentially of finite type over the spectrum of a Dedekind domain. If $\mathcal{H}^{n+1}(\mathbb{Z}/m(n)_{\mathrm{Nis}}^{X})=0$ for any positive integer $m$, then the sequence
$0\to R^{n+1}\alpha_{*}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}^{X}\to R^{n+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{X}\xrightarrow{\times m}R^{n+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}^{X}$
is exact.
Proof.
Since
$\mathbb{Q}/\mathbb{Z}=\displaystyle\lim_{\xrightarrow[m]{}}\frac{1}{m}\mathbb{Z}/\mathbb{Z}=\displaystyle\lim_{\xrightarrow[m]{}}\mathbb{Z}/m\mathbb{Z}$
where $m$ runs over all positive integers, we have an exact sequence
$0\to\mathbb{Z}/m\mathbb{Z}\to\mathbb{Q}/\mathbb{Z}\xrightarrow{\times m}\mathbb{Q}/\mathbb{Z}\to 0$
and a distinguished triangle
$\cdots\to\mathbb{Z}/m(n)^{X}_{\mathrm{Nis}}\to\mathbb{Q}/\mathbb{Z}(n)^{X}_{\mathrm{Nis}}\xrightarrow{\times m}\mathbb{Q}/\mathbb{Z}(n)^{X}_{\mathrm{Nis}}\to\cdots$
for a positive integer $m$. So we have an exact sequence
$\mathcal{H}^{n}(\mathbb{Z}/m(n)_{\mathrm{Nis}}^{X})\to\mathcal{H}^{n}(\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{Nis}}^{X})\xrightarrow{\times m}\mathcal{H}^{n}(\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{Nis}}^{X})\to 0$
by the assumption. Moreover, we have an isomorphism
$\mathcal{H}^{n}(\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{Nis}}^{X})\simeq R^{n}\alpha_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{Nis}}^{X}$
by Proposition 4.6. Hence the statement follows. ∎
Remark 5.15.
In the case where $X$ is a semistable family over the spectrum of a Dedekind domain, we have
$\mathcal{H}^{n+1}(\mathbb{Z}/m(n)_{\mathrm{Nis}}^{X})=0$
for any positive integer $m$ (cf. [24, Proposition 4.5], [24, Remark 4.7]). If $X$ is smooth over the spectrum of a Dedekind domain, we have
$\mathcal{H}^{s}(\mathbb{Z}(n)^{X}_{\mathrm{Nis}})=\mathcal{H}^{s}(\mathbb{Z}/m(n)^{X}_{\mathrm{Nis}})=0$
for $s\geq n+1$ and any positive integer $m$ by [7, p.786, Corollary 4.4].
We have the following as an application of Conjectures 5.12 and 5.13.
Proposition 5.16.
Let $\mathfrak{X}$ be a semistable family over $\operatorname{Spec}(\mathbb{Z})$ and $n$ a positive integer. Assume that Conjecture 5.12 and Conjecture 5.13 hold. Moreover, assume that
$\mathcal{H}^{n+2}(\mathbb{Z}(n)_{\mathrm{Nis}}^{\mathfrak{X}})=\mathcal{H}^{n+2}(\mathbb{Z}/m(n)_{\mathrm{Nis}}^{\mathfrak{X}})=0$
for any positive integer $m$. Then
$\Gamma\left(\mathfrak{X},R\alpha_{*}^{n+1}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)=\Gamma\left(\mathfrak{X},R\alpha_{*}^{n+2}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$ (108)
is finite and
$\Gamma\left(\mathfrak{X},R\alpha_{*}^{n+1}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)=\operatorname{Ker}\left(\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(k(\mathfrak{X}),\mathbb{Q}/\mathbb{Z}(n))\to\bigoplus_{x\in\mathfrak{X}^{(1)}}\operatorname{H}^{n+2}_{x}(\mathfrak{X}_{\mathrm{\acute{e}t}},\mathbb{Q}/\mathbb{Z}(n))\right).$ (109)
Proof.
First we prove the isomorphism (108). Let $i_{\eta}:\operatorname{Spec}(\kappa(\eta))\to\mathfrak{X}$ be the generic point. Consider a commutative diagram
${R^{n+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{n+2}\alpha_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{n+1}\alpha_{*}R(i_{\eta})_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${R^{n+2}\alpha_{*}R(i_{\eta})_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}.}$ (110)
Then the stark of $R^{n+1}\alpha_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$ at $x\in\mathfrak{X}$ is $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(\mathcal{O}_{\mathfrak{X},x}^{h},\mathbb{Q}/\mathbb{Z}(n))$ and the stark of $R^{n+1}\alpha_{*}R(i_{\eta})_{*}\mathbb{Q}/\mathbb{Z}(n)_{\mathrm{\acute{e}t}}$ at $x\in\mathfrak{X}$ is $\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(k(\mathcal{O}^{h}_{\mathfrak{X},x}),\mathbb{Z}(n)).$ Here $\mathcal{O}_{\mathfrak{X},x}^{h}$ is the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of $\mathfrak{X}$ at $x\in\mathfrak{X}$ and $k(\mathcal{O}^{h}_{\mathfrak{X},x})$ is the fraction field of $\mathcal{O}_{\mathfrak{X},x}^{h}$. So the left map is injective by [24, Proposition 4.2] and [22, p.35, Proposition 5]. Since
$\operatorname{H}^{s}_{\mathrm{\acute{e}t}}(k(\mathcal{O}^{h}_{\mathfrak{X},x}),\mathbb{Q}(n))=\operatorname{H}^{s}_{\mathrm{Zar}}(k(\mathcal{O}^{h}_{\mathfrak{X},x}),\mathbb{Q}(n))=0$
for $s\geq n+1$ by [7, p.781, Proposition 3.6], the bottom map is an isomorphism and so the upper map is injective. Since we have
$\operatorname{H}^{s}_{\mathrm{Zar}}(\mathcal{O}^{h}_{\mathfrak{X},x},\mathbb{Q}/\mathbb{Z}(n))=\lim_{\begin{subarray}{c}\to\\ m\end{subarray}}\operatorname{H}^{s}_{\mathrm{Zar}}(\mathcal{O}^{h}_{\mathfrak{X},x},\mathbb{Z}/m(n))$
and the sequence
$\operatorname{H}^{s}_{\mathrm{Zar}}(\mathcal{O}^{h}_{\mathfrak{X},x},\mathbb{Z}(n))\to\operatorname{H}^{s}_{\mathrm{Zar}}(\mathcal{O}^{h}_{\mathfrak{X},x},\mathbb{Q}(n))\to\operatorname{H}^{s}_{\mathrm{Zar}}(\mathcal{O}^{h}_{\mathfrak{X},x},\mathbb{Q}/\mathbb{Z}(n))$
is exact for any integer $s$, we have
$R^{n+2}\alpha_{*}\mathbb{Q}(n)_{\mathrm{\acute{e}t}}=\mathcal{H}^{n+2}(\mathbb{Q}(n)_{\mathrm{Nis}})=0$
by the assumption. Hence the upper map in the commutative diagram (110) is surjective and so is an isomorphism. Therefore we have the isomorphism (108).
Next we prove that $\Gamma\left(\mathfrak{X},R^{n+2}\alpha_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$ is finite. Since we have a quasi-isomorphism
$\tau_{\leq n+1}R\alpha_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\simeq\tau_{\leq n+2}\mathbb{Z}(n)_{\mathrm{Nis}}$ (111)
by Proposition 4.6 and the assumption, the homomorphism
$\operatorname{H}^{n+3}_{\mathrm{Nis}}(\mathfrak{X},\tau_{\leq n+1}R\alpha_{*}\mathbb{Z}(n))\to\operatorname{H}^{n+3}_{\mathrm{Nis}}(\mathfrak{X},\mathbb{Z}(n))$
is injective and we have
$\operatorname{H}^{n+3}_{\mathrm{Zar}}(\mathfrak{X},\mathbb{Z}(n))\simeq\operatorname{H}^{n+3}_{\mathrm{Nis}}(\mathfrak{X},\mathbb{Z}(n))$
by [7, p.781, Proposition 3.6]. Moreover, the sequence
$\operatorname{H}^{n+2}_{\mathrm{\acute{e}t}}(\mathfrak{X},\mathbb{Z}(n))\to\Gamma\left(\mathfrak{X},R\alpha_{*}^{n+2}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)\to\operatorname{H}^{n+3}_{\mathrm{Nis}}(\mathfrak{X},\tau_{\leq n+1}R\alpha_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}})$
is exact. So $\Gamma\left(\mathfrak{X},R^{n+2}\alpha_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$ is finitely generated over $\mathbb{Z}$ by Conjecture 5.12 and Conjecture 5.13. Moreover, $\Gamma\left(\mathfrak{X},R^{n+2}\alpha_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$ is a torsion group by the isomorphism (108). Hence $\Gamma\left(\mathfrak{X},R^{n+2}\alpha_{*}\mathbb{Z}(n)_{\mathrm{\acute{e}t}}\right)$ is finite.
Finally, we prove the isomorphism (109). By Lemma 5.14 and Remark 5.15, it suffices to show that the sequence
$0$ $\to\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(R,\mathbb{Z}/m(n))\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(R)^{(0)}}\operatorname{H}^{n+1}_{x}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))$
$\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(R)^{(1)}}\operatorname{H}^{n+2}_{x}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))$ (112)
is exact for a prime power $m$ where $R$ is the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of $\mathfrak{X}$ at a point $x\in\mathfrak{X}$. Moreover, it suffices to show the exactness of (5.2) in the case where $\operatorname{char}(R)=(0,p)$. Note that we have a quasi-isomorphism
$\tau_{\leq n+2}\left(\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}^{\operatorname{Spec}(R)}\right)\simeq\begin{cases}\mathfrak{T}_{r}(n)&\text{for $m=p^{r}$}\\ \mu_{m}^{\otimes n}&\text{for $(m,p)=1$}\end{cases}$
by [28, p.209, Remark 7.2], [24, Remark 4.7] and the assumption.
We prove the exactness of (5.2) by induction on $\operatorname{dim}(R)$.
Suppose that $\operatorname{dim}(R)=1$. Then the sequence (5.2) is exact by [23, p.30, Theorem 1.1].
Assume that the sequence (5.2) is exact in the case where $\operatorname{dim}(R)\leq s$. Suppose that $\operatorname{dim}(R)=s+1$. Let $Z$ be an irreducible component of the closed fiber of $\operatorname{Spec}(R)$. Then we have a spectral sequence
$E^{u,v}_{1}=\displaystyle\bigoplus_{x\in\operatorname{Spec}(R)^{(u)}\cap Z}\operatorname{H}^{u+v}_{x}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))\Rightarrow E^{u+v}=\operatorname{H}^{u+v}_{Z}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))$
(cf. [24, The proof of Proposition 3.8]). Since we have
$\operatorname{H}^{n+2}_{x}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))\simeq\operatorname{H}^{n+2-2u}_{\mathrm{\acute{e}t}}(\kappa(x),\mathbb{Z}/m(n-u))$
for $x\in\operatorname{Spec}(R)^{(u)}\cap Z$ by [27, p.540, Theorem 4.4.7] and [5], we have
$E^{u,v}_{\infty}=E^{u,v}_{2}=0$
for $u+v=n+2$ and $u\geq 2$ by [29, p.600, Theorem 4.1] and [4, Corollary 2.2.2]. So the homomorphism
$\operatorname{H}^{n+2}_{Z}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(R)^{(1)}\cap Z}\operatorname{H}^{n+2}_{x}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))$
is injective. Hence the sequence
$0\to\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(R,\mathbb{Z}/m(n))\to\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(U,\mathbb{Z}/m(n))\to\displaystyle\bigoplus_{x\in\operatorname{Spec}(R)^{(1)}\cap Z}\operatorname{H}^{n+2}_{x}(R_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))$ (113)
is exact by [24, Proposition 4.2] and [22, p.35, Proposition 5] where $U=\operatorname{Spec}(R)\setminus Z$. By [24, Proposition 4.5], [22, p.37, Corollary 7] and the assumption, we have an isomorphism
$\operatorname{H}^{s}_{\mathrm{Zar}}(R,\mathbb{Z}/m(n))=0$
for $n+1\leq s\leq n+2$. Since $Z$ is the spectrum of a regular local ring of positive characteristic, we have an isomorphism
$\operatorname{H}^{s}_{\mathrm{Zar}}(Z,\mathbb{Z}/m(n))=0$
for $s\geq n+1$. So we have an isomorphism
$\operatorname{H}^{s}_{\mathrm{Zar}}(U,\mathbb{Z}/m(n))=0$
for $n+1\leq s\leq n+2$. Hence we have an isomorphism
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(U,\mathbb{Z}/m(n))=\Gamma\left(U,R^{n+1}\alpha_{*}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}\right)$ (114)
by the quasi-isomorphism (111). Since $\operatorname{dim}(U)=s$, we have
$\Gamma\left(U,R^{n+1}\alpha_{*}\mathbb{Z}/m(n)_{\mathrm{\acute{e}t}}\right)$
$=$ $\operatorname{Ker}\left(\displaystyle\bigoplus_{x\in U^{(0)}}\operatorname{H}^{n+1}_{x}(U_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))\to\displaystyle\bigoplus_{x\in U^{(1)}}\operatorname{H}^{n+2}_{x}(U_{\mathrm{\acute{e}t}},\mathbb{Z}/m(n))\right)$ (115)
by the assumption of induction. Hence the sequence (5.2) is exact by (113), (114) and (5.2). This completes the proof.
Proposition 5.17.
Let $X$ be a smooth scheme over $\operatorname{Spec}(\mathbb{Q})$. Then the canonical map
$\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(k(X),\mathbb{Q}/\mathbb{Z})\to\prod_{x\in X^{(1)}}\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(k(\mathcal{O}_{X,x}^{h}),\mathbb{Q}/\mathbb{Z})$ (116)
is injective where $\mathcal{O}_{X,x}^{h}$ is the henselization of a local ring $\mathcal{O}_{X,x}$.
Proof.
Since we have an isomorphism
$\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(\mathcal{O}^{h}_{X,x},\mathbb{Q}/\mathbb{Z})\simeq\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(\kappa(x),\mathbb{Q}/\mathbb{Z}),$
the kernel of the canonical map (116) agrees with the kernel of
$\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(X,\mathbb{Q}/\mathbb{Z})\to\prod_{x\in X^{(1)}}\operatorname{H}^{1}_{\mathrm{\acute{e}t}}(\kappa(x),\mathbb{Q}/\mathbb{Z})$
by Lemma 5.8. Since $\mathbb{Q}$ is Hilbertian, the statement follows from [32, p.53, Remark 3]. ∎
By Proposition 5.16 and Proposition 5.17, we are able to raise the following questions:
Question 5.18.
Let $\mathfrak{X}$ be a regular scheme which is proper and flat over $\operatorname{Spec}(\mathbb{Z})$. Then
(a) Is the kernel of the canonical map
$\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(k(\mathfrak{X}),\mathbb{Q}/\mathbb{Z}(n))\to\prod_{x\in\mathfrak{X}^{(1)}}\operatorname{H}^{n+1}_{\mathrm{\acute{e}t}}(k(\mathcal{O}_{\mathfrak{X},x}^{h}),\mathbb{Q}/\mathbb{Z}(n))$ (117)
finite ?
(b) When is the kernel of the canonical map (117) trivial ?
Question 5.18 relates to Kato conjecture as follows:
Remark 5.19.
Let $\mathfrak{X}$ be a $d$-dimensional regular scheme which is proper and flat over the spectrum of the ring of integers of a number field $K$. Assume that $K$ is totally imaginary. Let $\mathcal{O}_{\mathfrak{X},x}^{h}$ be the henselization of the local ring $\mathcal{O}_{\mathfrak{X},x}$ of $\mathfrak{X}$ at a point $x\in\mathfrak{X}^{(1)}$. Then we have isomorphisms
$\operatorname{H}^{d+1}_{\mathrm{\acute{e}t}}(k(\mathcal{O}_{\mathfrak{X},x}^{h}),\mathbb{Q}/\mathbb{Z}(d))\simeq\operatorname{H}^{d+2}_{x}((\mathcal{O}_{\mathfrak{X},x})_{\mathrm{\acute{e}t}},\mathbb{Q}/\mathbb{Z}(d))\simeq\operatorname{H}^{d}_{\mathrm{\acute{e}t}}(\kappa(x),\mathbb{Q}/\mathbb{Z}(d-1))$
(cf.[23, p.31, Theorem 1.2], [25, Proposition 4.8]). So, if Kato conjecture (cf.[14, p.125, Conjecture 0.3]) holds, then the canonical map
$\operatorname{H}^{d+1}_{\mathrm{\acute{e}t}}(k(\mathfrak{X}),\mathbb{Q}/\mathbb{Z}(d))\to\displaystyle\bigoplus_{x\in\mathfrak{X}^{(1)}}\operatorname{H}^{d+1}_{\mathrm{\acute{e}t}}(k(\mathcal{O}_{\mathfrak{X},x}^{h}),\mathbb{Q}/\mathbb{Z}(d))$
is injective.